### v1

# Quasi-randomization tests for network interference

Abstract
Many classical inferential approaches fail to hold when interference exists among the population units. This amounts to the treatment status of one unit affecting the potential outcome of other units in the population. Testing for such spillover effects in this setting makes the null hypothesis non-sharp. An interesting approach to tackling the non-sharp nature of the null hypothesis in this setup is constructing conditional randomization tests such that the null is sharp on the restricted population. In randomized experiments, conditional randomized tests hold finite sample validity. Such approaches can pose computational challenges as finding these appropriate sub-populations based on experimental design can involve solving an NP-hard problem. In this paper, we view the network amongst the population as a random variable instead of being fixed. We propose a new approach that builds a conditional quasi-randomization test. Our main idea is to build the (non-sharp) null distribution of no spillover effects using random graph null models. We show that our method is exactly valid in finite-samples under mild assumptions. Our method displays enhanced power over other methods, with substantial improvement in complex experimental designs. We highlight that the method reduces to a simple permutation test, making it easy to implement in practice. We conduct a simulation study to verify the finite-sample validity of our approach and illustrate our methodology to test for interference in a weather insurance adoption experiment run in rural China.
Keywords: Interference, Networks, Causal Inference, Quasi-Randomization Tests, Peer Effects, Random Graphs
1   Introduction
A common assumption in causal inference is that of the Stable Unit Treatment Value Assumption (SUTVA). This assumes away any dependence of one unit’s treatment status on the outcome of another unit. Whilst this allows for unbiased treatment effect estimation and valid inferential strategies, there are many experimental and observational settings where interference between units takes place (Cox (1958)). There are multiple pathways via which interference can take place in the population. It can be a result of the spread of treatment amongst the population units, the respective outcomes of treated units affecting the outcomes of others, or a combination of both. For example, the spread of a contagious disease is contingent upon the number of infected people in the population (Halloran and Struchiner (1995)). Thus, the estimation of any preventive measures taken will be challenged because of a likely SUTVA violation. Consider a population-level policy evaluation program, like an insurance product campaign (Cai et al. (2015)). Due to social network ties in the population, information dissemination is a plausible interference mechanism that can occur and affect evaluated estimates. Similar phenomena can be seen amongst experiments run on social media networks, like Linkedin, or online marketplaces like eBay, owing to the interconnected nature of these platforms (Pouget-Abadie et al. (2019); Blake and Coey (2014)). The total treatment effect, in such cases, can be decomposed into two effects, one of the direct effect and the other of spillover effects (Sobel (2006)).
The classical method for testing treatment effects is the “Fisher Randomization Test” or FRT (Fisher (1936)). It is used to test for a sharp null hypothesis of no treatment effect for all units in the population. The test exploits the imputability of counterfactuals under the null. This enables the determination of the randomization null distribution of a test statistic of choice. If one were to test for spillover effects, it would render the null non-sharp, leading to imputability issues for the test statistic. There is a growing body of literature that addresses this issue by considering a carefully determined subset of units in the population and treatment assignment vector, such that on these selected units and treatment vectors, the null is sharp (Aronow and Samii (2017); Athey et al. (2018); (Basse et al., 2019b)). Such conditional randomization tests also lead to valid exact p-values, as is the case with the Fisher Randomization Test. Basse et al. (2019b) build a permutation test under the special case of clustered interference. Whilst a permutation test is computationally cheaper, the method lacks generalizability because of the structured nature of interference considered. Aronow and Samii (2017) and Athey et al. (2018) generalize the above for a given network and consider a range of non-sharp null hypotheses for testing. Puelz et al. (2022) presents a graph-theoretic approach to constructing the conditioning mechanism of a sub-population of units and the corresponding treatment assignment vectors for which the null is sharp. However, the proposed methodology is NP-hard in implementation and, thus, in practice, does not involve observing the treatment assignment to construct the conditioning mechanism. Whilst the above considers the problem of testing for interference in an experimental setting, the literature on testing for spillover effects in an observational setting is still in its infancy.
In this paper, we consider the network to be coming from a distribution itself. This is in contrast with the setup considered in the current methods that assume the network to be fixed. We draw a parallel here to estimating the probability of unit treatment assignment via propensity scores in the case of observational studies. The propensity of two nodes in a network being connected can be looked at as being a function of their network characteristics, covariates, or both. This leads us to think of these linkages as coming from a generating process with an associated probability distribution. The current methods are based on strict randomization inference with no distributional assumptions for either the network or the potential outcomes, achieving a generalization of the Fisher Randomization Test. Whilst this allows for widespread validity of the methods proposed, it comes at a cost of loss of power. We emphasize that leveraging appropriate stochastic assumptions can significantly boost power and lead to methodological advancements. In this paper, we restrict the network data-generating process to be characterized by any arbitrary degree distribution. First proposed in Newman et al. (2001), random graphs with arbitrary degree distribution represent real-world social networks closely owing to their ability to capture heterogeneity in degree. Given the mild nature of the distributional assumption, we obtain a large class of network data-generating processes. This assists us in building a quasi-randomization test based on the network distribution. We show that conditional on the observed treatment, the (quasi) randomization distribution of a test statistic leads to valid p-values. The challenge of the imputability of a test statistic of choice under the null, being the major methodological contribution of the current methods in the literature, is naturally overcome in our method.
Model-based approaches with distribution assumptions on outcome model function have been previously considered in the literature for identifying and estimating peer effects (Manski (1993); Manski (2013); Bowers et al. (2013); Toulis and Kao (2013); Blume et al. (2015)). More recently, Li and Wager (2022) and Leung (2020) proposed estimation strategies for spillover effects under a distributional assumption on the network formation model. The inferential methods for the proposed estimation procedures involve either non-standard procedures or asymptotic approximations (Imbens and Manski (2004)). With a careful identification of causal estimates, such large sample normal approximations of confidence interval conform to a Neymanian perspective. Our first contribution is to extend the random graph interference model into the potential outcome framework. We generalize and define sharp/non-sharp null hypotheses for random graph models. Building upon this, we develop a randomization-based test for a given random graph model. This leads to finite sample validity of the p-value obtained for the oracle method, in line with Fisherian procedures.
Our method addresses several technical challenges encountered in the current methods. We show that randomization tests based on network-generating processes lead to imputable test statistics under the null, conditional on the treatment assignment. This overcomes a major challenge faced in randomization methods based on random treatment assignment. The current methodology involves restricting attention to a carefully chosen subset of focal units and claims null of no spillover effects for only these units. In contrast, our method tests for a stricter class of nulls with no spillover effects on all the units in line with the Fisher Randomization Test. Another hurdle faced by the current methodologies is the extension of the method to complex experimental designs. Determination of conditional randomization distribution given a conditioning mechanism is not straightforward. The restriction of fixed focal units in the randomization null reduces the sample space of possible treatment assignment vectors in the null. It can lead to an enormous loss of power or render the null degenerate completely. There is a burgeoning set of literature focussing on improving the current methods by incorporating design-based techniques (Puelz et al. (2022); (Basse et al., 2019a)). We emphasize that our method is easily generalized to various experimental designs, and we show this for two experimental designs, namely, completely randomized experiments and cluster randomized experiments.
Our second major contribution is the computational efficiency of the proposed method. We propose that the random graph model be characterized by any arbitrary degree distribution. We build a permutation test by generating the null over the set of graphs with the same degree sequence as the observed graph. Random graph generation of a prescribed degree sequence admits fast and efficient polynomial time algorithms for large graphs (Viger and Latapy (2005); Fosdick et al. (2018)). Thus, via permutation of edges of the graph in a manner that preserves degree sequence, we obtain valid randomization tests. This is closely related to literature in random graph null model used in network analysis (Van Koevering et al. (2021)) with applications in protein structure discovery and community detection among many (Milenković et al. (2009); Sah et al. (2014)). To the best of our knowledge, we are the first to formally introduce random graph null models as an extension of the Fisher Randomization Test. To illustrate our method, we re-analyze a completely randomized experiment in rural China where rice crop farmers are informed about a weather insurance product (Cai et al. (2015)). The hypothesis is that farmers with a higher proportion of informed neighbors display an enhanced adoption rate. We display our methodology on a subset of the original experiment performed and obtain significant information spillover among the farmers. The setting displays the importance of such spillover effects in evidence-based policymaking under resource-constrained settings.
The remaining paper is structured as follows. In Section 2, we set up notation and formalize the null hypothesis of interest, incorporating network random variable. We also describe the present state-of-the-art methodology of using a conditioning mechanism for testing non-sharp null. Section 3 presents a randomization test based on a given network-generating process and its application to different experimental designs. Section 4 extends this further to random graph models and presents our main methodology, which can be implemented as a permutation test. In Section 5, we present extensive numerical simulations to evaluate and compare the power of the proposed test with the current methods in the literature. Section 6 and 7 illustrate our methodology as an empirical application and conclude the paper. All theorem proofs, algorithms, and additional numerical simulation results are presented in the Appendix.
2   Setup
2.1   Notation
Consider a population =$\mathbb{P}=\{1,2,...,N\}$ where $N$ is the total number of units in the population. Consider a vector $Z\in{\mathbb{Z}}^{N}$ where each element in the vector indicates treatment assigned to the corresponding indexed unit in the population. Here, $\mathbb{Z}$ represents the set of possible treatments to which the population units can be exposed. For the purposes of this article, we take =$\mathbb{Z}=\{0,1\}$. Hence, any treatment assignment vector is an N-dimensional binary vector. Consider the treatment vector to be drawn from a probability distribution $P_{Z}:\{0,1\}^{N}\rightarrow[0,1]\subset\mathbb{R}$. Let $Z_{obs}$ be the realized outcome drawn from the treatment assignment distribution $P_{Z}$.
We consider the population $\mathbb{P}$ to be connected via an undirected graph $G=(\mathbb{P},E)$ where $E\subseteq\mathbb{P}\times\mathbb{P}$. For distinct units $i$ and $j$ such that $(i,j)\in E$ represents a connection between the two units. With some slight abuse of notation, we also represent the adjacency matrix of the graph by $G$. Thus, $G$ is an $N\times N$ binary matrix. Here, for some $i$ and $j$, $G_{ii}=0$ and $G_{ij}=G_{ji}$ owing to $G$ being a undirected graph with no self-loops. If $G_{ij}=1$, then we say unit $i$ and $j$ are neighbours. This can also be denoted by $i\in\mathcal{N}_{G}(j)$ or vice-versa. Here, $\mathcal{N}_{G}(i)$ represents all the neighbors of unit $i$ in graph $G$. We may simply write $\mathcal{N}(i)$ if the graph is clear in context. We define the total number of neighbors of unit $i$ as the degree of $i$. This is denoted by $deg(i)$ and is equal to $|N(i)|$. We denote the distance between two units $i$ and $j$ as $dist(i,j)$. It is defined as the shortest path length between the two units. By convention, we take $dist(i,i)=0$ and $dist(i,j)=\infty$ if no path exists between the two distinct units. Consider the graph $G$ to be coming from the probability distribution $P_{G}:\{0,1\}^{N\times N}\rightarrow[0,1]\subset\mathbb{R}$. We denote the observed sample from the distribution $P_{G}$ to be $G_{obs}$. Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$ where $X_{i}$ is a m-dimensional vector. $X_{i}$ is fixed and influences the outcome via the potential outcome function. Let $X$ denote the $N\times M$ matrix where each row corresponds to the respective unit’s covariate vector. We define the potential outcome function $Y:\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow\mathbb{R}^{N}$ as a real-valued function that takes treatment assignment and network adjacency matrix as arguments. We will consider $Y$ to be fixed, and each element $i$ of an image in $Y(Z,G)$, $Y_{i}(Z,G)$ denotes the outcome of $i^{th}$ unit in the population under the treatment assignment $Z$ and network assignment $G$. We denote the corresponding potential outcome as $Y_{obs}$ that is, $Y_{obs}=Y(Z_{obs},G_{obs})$. We note that the potential outcome function implicitly contains information in $X$ and can be written as $Y_{X}(Z,G)$. For the remainder of the article, we will use the notation $Y(Z,G)$ for readability.
2.2   Definitions
The assignment mechanism defines the joint probability distribution of treatment assignment vectors given covariates and potential outcomes, denoted by $P(Z|X,Y(0),Y(1))$. The following are three key assumptions that are considered in the causal inference literature for assignment mechanism: (i) individualistic assignment, which restricts the treatment assignment of a unit to its own covariates and outcomes (ii) probabilistic assignment, which assumes non-integer probabilities of treatment assignment to units (iii) unconfoundedness assumption, which restricts the dependence of the probability assignment of units to only covariates, and not the potential outcome (Imbens and Rubin (2015)). Individualistic assignment mechanism assumption, given a well-defined potential outcome function, constitutes Stable Unit Treatment Value Assumption (SUTVA). We argue that the underlying network structure among the population is an implicit form of treatment. Hence, we generalize the treatment to a bivariate treatment assignment vector $(Z,G)$. Whilst the usual assumptions described above are plausible for the marginal distribution of direct treatment $Z$, the individualistic assumption is not valid for the marginal distribution of network treatment $G$ given the nature of the variable. We look at the distribution of joint treatment assignment vector $(Z,G)$ in this setup.
$P(Z,G|X)=P(G|Z,X)\cdot P(Z|X),$
assuming joint treatment assignment is unconfounded, that is,
$P(Z,G|X,Y(0),Y(1))=P(Z,G|X).$
We remark that we took the assumption of the unconfoundedness of both treatment assignments into consideration. We formally define the standard sharp null hypothesis of no (direct) treatment effect when the population is not interconnected.
Hypothesis 1.
Sharp null hypothesis of no treatment effect:
$\forall i\in[N]\quad Y_{i}(Z)=Y_{i}(Z^{\prime})\qquad\forall\quad Z,Z^{\prime}\in\{0,1\}^{N}.$
The above stated null hypothesis imposes specified forms of restrictions on the potential outcome function. Extending this to our setup, we get the following definition of the sharp null hypothesis.
Hypothesis 2.
Sharp null hypothesis of no (joint) treatment effect:
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}.$
We now give a definition of a test statistic that will be used for testing such null hypotheses. Following our setup, we will use the joint treatment assignment, $(Z,G)$, for the rest of the article.
Definition 1.
A test statistic $T$ is a real-valued function of the treatment assignment and $Y_{obs}$ given the covariates $X$, $T_{X}(Z,G,Y_{obs})$.
Under sharp null, given an observed outcome, $Y_{obs}$, we can impute the outcome function $Y$ for all treatment assignments. This is not possible when the null hypothesis is non-sharp. We formally define the non-sharp null hypothesis below.
Hypothesis 3.
Non-sharp null hypothesis:
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\textbf{Z}_{i}\times\textbf{G}_{i}\subset\{0,1\}^{N}\times\{0,1\}^{N\times N}.$
Here, the subset restriction, denoted by $\textbf{Z}\times\textbf{G}:=\textbf{Z}_{i}\times\textbf{G}_{i}$, on treatment assignment is chosen in accordance with the null hypothesis of interest. We elaborate further using the example of testing null of no spillover effects below. Consider the null hypothesis of no spillover effects if the distance between two units is at least k.
Example 1.
(Null Hypothesis: no spillover effects at distance k and beyond) Let the $k-$distance function $D_{G}(i)$ for a unit $i$ in graph $G$ be defined as below:
[…]
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
[…]
Note the non-imputability of potential outcome function under the null considered above for different treatment assignments. Specifically, given $(Z_{obs},G_{obs})$, we can compute potential outcomes for treatment assignments in the restriction set defined above. Since the restriction set is a strict subset of the treatment assignment space, the null is not sharp.
2.3   Null hypothesis of no spillover and conditional randomization test
In this paper, we are concerned with testing first-order spillover effects. We take $k=1$ in Example 1 and state the null hypothesis of no spillover effects below.
Hypothesis 4.
(Null hypothesis of no spillover effects)
$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
[…]
We highlight that the treatment $G$ is not restricted in the above hypothesis. We now describe the spillover mechanism as a component in the potential outcome function $Y(Z,G)$. In the following example, $Y(Z,G)$ is linearly separable in two components, one of direct effect and another of spillover effect, as stated below.
We re-state the null hypothesis of no spillover effects, as described in Hypothesis 4.
Example 2.
(Null hypothesis of no additive spillover effects)
$Y_{i}(Z,G)$ $=Y^{1}(Z_{i})+Y^{2}(Z\cdot G_{i}),$
$=Y^{1}(Z_{i})+Y^{2}(Z_{\mathcal{N}_{G}(i)}).$
Using the above assumption of linearly separable joint treatment effect, we get
$\forall i\in[N]\quad Y^{2}(Z_{\mathcal{N}_{G}(i)})=Y^{2}(Z^{\prime}_{\mathcal{N}_{G^{\prime}}(i)})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
[…]
A similar approach to defining non-sharp null of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al. (2019b); Puelz et al. (2022)). The exposure functions are functions of the treatment assignment vector, $Z$, such that restriction sets can be created on $Z$ by partitioning $Z$ on the range of the exposure function. In the above Example 1, we can define an exposure function as below.
$f_{G}(Z)=\frac{Z\cdot D_{G}^{(k-1)}(i)}{|D_{G}^{(k-1)}(i)|}.$
We obtain the setup described in the literature, given a fixed $G$ using the above-described exposure function as a restriction set. The setup assumes a spillover mechanism via an exposure function. Since we consider the graph underlying the population to be random, we define restrictions on the joint treatment assignment in Hypothesis 3. A spillover mechanism, if existing, will be captured in the potential outcome function, and hence, we do not separately define a joint exposure mapping function by generalizing the exposure mappings. Careful identification of a spillover mechanism will be required for the inversion of the hypothesis test to obtain the confidence interval. We describe the currently proposed method in the literature of a conditional-randomization test procedure to obtain valid p-values (Puelz et al. (2022)).
Procedure: Consider $C$ to be a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|C)$ is imputable under the null.
  1. 1.
    Draw $Z_{obs}\sim P(Z)$ and obtain $Y_{obs}$.
  2. 2.
    Draw $C\sim P(C|Z_{obs})$ and compute $T(Z_{obs}|C)$.
  3. 3.
    Compute p-value as $P_{Z|C}\left(T(Z|C)>T(Z_{obs}|C)\right)$.
The above-stated procedure holds finite-sample validity. We note that the conditioning mechanism $C$ is drawn from $P(C|Z_{obs})$, which should be determined by the analyst. One must ensure the imputability of the test statistic $T(Z|C)$ when constructing a mechanism for $C$. We require the test statistic to be imputable to obtain its sampling distribution under the null, even though we observe only one treatment assignment vector. Below, we define the imputability of a test statistic for the joint treatment assignment.
Definition 2.
Consider a test statistic $T(Z,G,Y_{obs})$. A test statistic is called imputable under $\mathcal{H}_{0}$ if
$T(Z,G,Y(Z,G))$ $=T(Z,G,Y(Z^{\prime},G^{\prime}))$
$\quad P(Z,G|(Z,G)\in\textbf{Z}\times\textbf{G}\text{ },\mathcal{H}_{0}),P(Z^{\prime},G^{\prime}|(Z^{\prime},G^{\prime})\in\textbf{Z}\times\textbf{G}\text{ },\mathcal{H}_{0})>0.$
The choice of a (imputable) test statistic is not unique, though, and different test statistics can lead to different powers of the test. This is determined by the test statistic’s responsiveness to the null hypothesis.
3   Quasi-randomization test based on network distribution
We present the quasi-randomization test for the sharp null of no spillover effects. Our main idea is to generate the null sampling distribution of a test statistic by randomizing over the network-generating process. Intuitively, under the null of no spillover effect, the underlying network structure does not contribute to the total treatment effect. This can also be seen in Hypothesis 4 where the network variable $G$ is unrestricted in the, otherwise, non-sharp null. This leads to the imputability of a test statistic of choice by fixing the direct treatment assignment $Z$ and randomizing over the distribution of network $G$. We formally define the conditional test statistic, $T_{c}$, below and show its imputability under the null Hypothesis 4.
Proposition 1.
Define $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$. $T_{c}(G|Z_{obs})$ is imputable under the null hypothesis of no spillover effects in Hypothesis 4.
Proof.
Note that under $\mathcal{H}_{0}$,
$Y(Z,G)=Y(Z^{\prime},G^{\prime})\qquad Z=Z^{\prime},$
Thus,
$T_{c}(G|Z_{obs})$ $=T(Z_{obs},G,Y(Z_{obs},G)),$
$=T(Z_{obs},G,Y(Z_{obs},G^{\prime}))\qquad P(G^{\prime})>0\text{ under }\mathcal{H}_{0}.$
Since $Z_{obs}$ is arbitrary, we conclude imputability of the $T_{c}(G|Z_{obs})$ under the null $\mathcal{H}_{0}$. ∎
We state our proposed quasi-randomization testing procedure below. The procedure assumes the knowledge of the network-generating process $P(G)$. An elaborate consideration of the distribution of the network will be presented in Section 4.
Procedure 1: Let $(Z_{obs},G_{obs},Y_{obs})$ be the observed joint treatment assignment and its corresponding outcome. Consider a test statistic $T(Z,G,Y_{obs})$.
  1. 1.
    Define a new test statistic $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$.
  2. 2.
    Impute the observed test statistic value ${T_{c}}_{obs}=T_{c}(G_{obs}|Z_{obs})$.
  3. 3.
    Consider the randomization distribution $P(G|Z)$.
  4. 4.
    Compute $pval(Z_{obs},G_{obs},Y_{obs}):=\mathbb{E}_{G|Z_{obs}}[\mathcal{I}(T_{c}(G|Z_{obs})>{T_{c}}_{obs})|Z_{obs}]$.
The following theorem proves the finite sample validity of Procedure 1.
Theorem 1.
Consider the null hypothesis, $\mathcal{H}_{0}$, in Hypothesis 4. Let $Z_{obs}\sim P(Z)$ be the treatment assignment of a randomized experiment. We assume that $G_{obs}\sim P(G)$ where $P(G)$ is known. Let $T(G|Z)$ be an imputable test statistic under the null $\mathcal{H}_{0}$. Then, Procedure 1 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
$\mathbb{E}\bigl{[}\mathcal{I}(\text{p-val}(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1),$
Here, the expectation is with respect to the distribution of $P(G_{obs}|Z_{obs})$.
Proof.
The proof is given in Appendix B.1. ∎
Note that conditional validity implies unconditional validity of the procedure too. A crucial component of the proof is the identification of the conditional distribution $P(G|Z)$ for the validity of Procedure 1. This is contingent upon the experimental design, and different experimental designs will lead to different conditional randomization distributions. We compare this with the state-of-the-art methodology presented in Section 2 where the null randomization distribution comes from $P(Z|C)$. Here, $C$ is a conditioning mechanism consisting of a subset of chosen units. Thus, different experimental designs lead to different $P(Z|C)$. A common problem faced in the literature is in the implementation of complex experimental designs like cluster randomized experiments. In a cluster-randomized experiment, the population is classified into a set of clusters where treatments are assigned at the cluster level. If a cluster is treated, then all units in that cluster receive the treatment or vice-versa. Fixing the treatment assignment of units in the conditioning mechanism might render the null randomization distribution $P(Z|C)$ degenerate, given the restricted span of cluster treatment assignments space. For example, assume there is a focal unit in every cluster. Then, conditioning on the observed treatment of focal units, we obtain $Z|C=\{Z_{obs}\}$ since all the unit treatment assignments can be deterministically imputed by the treatment assignment of the focal unit in that cluster. This reduction in the support of conditional distribution leads to either degeneracy or enormous loss of power. Our methodology captures the null distribution by randomizing over $G|Z$, where the null distribution admits large support. We discuss this in Sections 3.1.1 and 3.1.2, where we identify the null randomization distribution under different experimental designs.
3.1   Experimental design
Here, we discuss how Procedure 1 can be implemented in different experimental designs. A change in experimental design corresponds to a change in (direct) treatment assignment mechanism. Thus, correctly identifying the distribution $P(G|Z)$ is important to ensure the validity of Procedure 1. We find $P(G|Z)$ for two experimental designs, completely randomized experiments and cluster randomized experiments.
3.1.1   Completely randomized experiment
Consider a random sample of units of size $N_{t}$ from $\mathbb{P}$, who are selected in the treatment group. Let $N_{t}$ be fixed. Thus, the remaining units are in the control group. Let $N_{c}=N-N_{t}$ be the size of the units in the control group.
$P(Z=z)=\begin{cases}\frac{1}{\binom{N}{N_{t}}}&\sum_{i=1}^{N}z_{i}=N_{t},\\ 0&\text{otherwise}.\end{cases}$ (1)
We remark that $G\!\perp\!\!\!\perp Z$ in Equation 1. Hence, for completely randomized experiments,
$P(G|Z)=P(G).$
Since $P(G)$ is not identified by design, we will explore random graph models in Section 4.1.
3.1.2   Cluster randomized experiment
Consider the following partition of the population $\mathbb{P}$ into clusters $\{C_{1},C_{2},....,C_{k}\}$. That is,
$C_{i}\cap C_{j}=\emptyset$ $\hskip 14.22636pt\forall i\neq j\in[k],$
$\cup_{i=1}^{k}C_{i}=\mathbb{P}.$
Given such a partition, treatment assignment is done at the cluster level instead of the unit level. Following Ugander et al. (2013), we perform a Bernoulli experiment at the cluster level. Thus, we obtain
$W_{i}\sim Bernoulli(p),$
$Z_{j}=\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i}).$
[…]
$P(G|Z)$ $=\frac{P(Z|G)\cdot P(G)}{P(Z)},$ (2)
$=\frac{\prod_{i=1}^{k}p^{\mathcal{I}(\exists j\in[N]\text{ s.t. }j\in C_{i},z_{j}=1))}\cdot\prod_{i=1}^{k}(1-p)^{\mathcal{I}(\exists j\in[N]\text{ s.t. }j\in C_{i},z_{j}=0))}\cdot P(G)}{P(Z)}.$
Here, we can see that identifying null randomization distribution becomes challenging in the case of graph cluster randomization experiments. We propose a new approach that considers sufficient statistics of the random graph model to build a permutation test.
4   Quasi-permutation test via distributional properties
Identification of null randomization distribution $P(G|Z)$ involves knowledge of the distribution of the network-generating process, which is not identified by the design of the experiment. To this end, we explore random graph models to model $P(G)$. As discussed in the previous section, this identification of null randomization distribution becomes challenging when the complexity of the experimental design increases. This challenge is overcome in our proposed methodology, which is a conditional permutation test. The method does not involve estimating the null distribution. We discuss this in Section 4.1.2.
4.1   Random graph models
In this section, we describe various random graph models based on which $P(G)$ can be modeled. The seminal paper by Erdős and Rényi Erdős et al. (1960) introduced the concept of random graph for the first time. The Erdős-Rényi random graph model has since been studied widely, owing to exact computations of desirable graph properties. We now present the model formally in Section 4.1.1. However, the Erdős-Rényi random graph model does not represent most real-world networks, such as social networks, supply-chain networks, and scientific collaboration networks, among many others. We look at a generalization of Erdős-Rényi random graph that is better adept at modeling such real-world networks, proposed by Newman et al. (2002), in Section 4.1.2.
4.1.1   Erdős-Rényi random graph model
We formally describe the Erdős-Rényi random graph model (Erdős et al. (1960)). The below-described model is the probabilist variant of the (Erdős-Rényi) random graph model. We consider the number of nodes, $N$, and edges, $|E|$, to be fixed, and any two nodes will be linked independently with an edge with a fixed probability $p$. That is,
$P(G=G_{obs}|p)=\binom{\binom{N}{2}}{|E|}\cdot p^{|E|}\cdot(1-p)^{\binom{N}{2}-|E|}.$ (3)
In order to model $P(G)$ using Equation (3), we have to estimate the parameter $p$. We obtain the MLE of the parameter $p$ and estimate it as $\frac{|E|}{\binom{N}{2}}$. We conduct an extensive numerical simulation study and verify the finite sample validity of the proposed randomization test. We also obtain the power of the test under different scenarios, which are presented in Appendix C. An Erdős-Rényi random graph model is the simplest network model one can consider, and it may not be representative of the kind of networks we observe in the real world. For example, in social networks, one observes a high degree of clustering in the graph (Watts and Strogatz (1998)). One drawback of the model is the independence and identical probability of formation of two edges in the graph. One way to characterize this is to evaluate the degree distribution of the random graph. It has been shown that the degree distribution of a large Erdős-Rényi random graph is a Poisson distribution. Most real-life networks violate this property and have highly skewed degree distribution. In the next section, we generalize this and characterize random graphs with arbitrary degree distribution.
4.1.2   Random graphs with arbitrary degree distribution
To represent more realistic degree distributions observed in real life, Newman et al. (2001) considered characterizing the random graph by its degree distribution. We formulate the model below.
Definition 3.
Consider a random graph $G\in\{0,1\}^{N\times N}$ such that $P(G)=f(deg(i):i\in\mathbb{P}|p_{1},p_{2},...,p_{N})$. Here, $\{p_{i}\}_{i=1,...,N}$ represent the degree distribution, that is, $p_{i}$ is the probability that a randomly selected unit has degree $i$ and $f(.)$ represents the probability mass function. We call $G$ a random graph with arbitrary degree distribution.
A key feature of random graphs with arbitrary degree distribution is their ability to capture a large class of degree distribution. This generalizes the Erdős-Rényi random graph model and, at the same time, ensures tractability of important structural graph properties (Newman et al. (2002)). In order to deploy Procedure 1, we have to identify $P(G)$. This involves estimating $(p_{1},p_{2},...,p_{N})$ with a single realization of the graph, and we cannot directly implement Procedure 1 without estimating $P(G)$. In the case of random graphs with arbitrary degree distribution, it is proven that we can possibly estimate the parameters of the random graph with just one realization but with some uncertainty (Chatterjee et al. (2011)). Owing to estimation issues of the null distribution, we propose a new method that does not involve such estimations and reduces to a simple permutation test. In our proposed methodology, we use sufficient statistics of the random graph distribution, degree sequence in this case, and develop a method conditioning on the sufficient statistic. We use the symmetric nature of the $P(G)$ via the equivalence relation and build a conditional permutation test, as described below.
4.2   Permutation test based on equiprobability events
Here, we present our main result, which proposes the procedure to obtain a conditional randomization test without estimating the distribution. Our main idea is to identify a sufficient statistic for the random graph model, which is degree sequence in the case of random graphs with arbitrary degree distribution. We then sample the null distribution over all the graphs with the same degree sequence as the observed graph. Using the equiprobability of the event set of all the graphs with the same degree sequence, we obtain a permutation test. We now formally present the definitions below and proceed to state and prove the finite sample validity of the proposed method.
We define the degree sequence of a graph $G=(V,E)$.
Definition 4.
Consider $D=(d_{1},d_{2},...,d_{N})$ such that $d_{1}\leq d_{2}\leq...\leq d_{N}$. We call $D$ to be the degree sequence of $G=(V,E)$ if there exists a one-to-one mapping from $\{deg(i):i\in V\}$ to $D$. We denote $D$ as $deg(G)$.
Not all $D\in\mathbb{W}^{N}$ can be a valid degree sequence. For example, consider $D=\{1,2\}$ as a candidate for degree sequence for a graph of size $2$. We remark that the maximum degree of a node in the graph is at most $(N-1)$. Since no vertex in a graph of size two can have a degree of more than 1, $D$ is not a valid degree sequence. Note that graphs with the same degree sequence are not necessarily isomorphic graphs. This implies that the vertices of the graph cannot be permuted to obtain the other graph with edges preserved (see Figure 1).
Figure 1: Non-isomorphic graphs with same degree sequence
The converse holds true. Consider a set of distinct graphs with the number of units fixed to $N$ and a degree sequence of the graphs fixed to a valid degree sequence $D$. Then, this set of all graphs with degree sequence $D$ will form an equivalence class, and we can partition the space of all random graphs with this equivalence relation. We define this below.
Definition 5.
Consider a random graph $G\in\{0,1\}^{N}\sim P(G)$ where $P(G)=f(deg(i):i\in V|p_{1},p_{2},...,p_{N})$ where $\{p_{i}\}_{i=1,...,N}$ represent the degree distribution of $G$. Let $\mathcal{G}$ be the support of $P(G)$. We define a relation $d:\mathcal{G}\rightarrow\mathcal{G}$ such that $G\equiv_{d}G^{\prime}$ if $deg(G)=deg(G^{\prime})$.
It is a quick check that $d$ defines an equivalence relation. Hence, we can partition the support of $P(G)$, where each part in the partition represents an equivalence class of $D$.
Proposition 2.
Consider $d:\mathcal{G}\rightarrow\mathcal{G}$ such that $G\equiv_{d}G^{\prime}$ if $deg(G)=deg(G^{\prime})$, as defined in Definition 5. Then, $d$ is an equivalence relation.
Proof.
$G\equiv_{d}G\quad\forall G\in\mathcal{G}$. Hence, $d$ is reflexive. If for any distinct $G$ and $G^{\prime}$ in $\mathcal{G}$, $G\equiv_{d}G^{\prime}\implies deg(G)=deg(G^{\prime})\implies deg(G^{\prime})=deg(G)\implies G^{\prime}\equiv_{d}G$. Hence, $d$ is symmetric. Similarly, for any three distinct $G,G^{\prime}$ and $G^{\prime\prime}$ in $\mathcal{G}$, $G\equiv_{d}G^{\prime},G^{\prime}\equiv_{d}G^{\prime\prime}\implies deg(G)=deg(G^{\prime})$ and $deg(G^{\prime})=deg(G^{\prime\prime})\implies deg(G)=deg(G^{\prime\prime})\implies G\equiv_{d}G^{\prime\prime}$. Hence, $d$ is transitive. ∎
Procedure 2: Let $(Z_{obs},G_{obs},Y_{obs})$ be the joint treatment assignment vector and its corresponding outcome. Consider a test statistic $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$.
  1. 1.
    Impute the observed value of test statistic ${T_{c}}_{obs}=T_{c}(G_{obs}|Z_{obs})$.
  2. 2.
    Define the equivalence class $[G_{obs}]=\{G\in\mathcal{G}|\quad G\equiv_{d}G_{obs}\}$. Here, $\equiv_{d}$ is the relation defined in Definition 5.
  3. 3.
    Compute p-value as $p(Z_{obs},G_{obs},Y_{obs}):=\frac{1}{|[G_{obs}]|}\cdot\sum_{G\in[G_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs}))$.
Theorem 2.
Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$. Let $Z_{obs}\sim P(Z)$ be the treatment assignment of a cluster randomized experiment. We assume that $G_{obs}\sim P(G)$ where $G$ is a random graph with arbitrary degree distribution. Let $T(G|Z)$ be an imputable test statistic under $\mathcal{H}_{0}$. Then, Procedure 2 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
$\mathbb{E}\bigl{[}\mathcal{I}(p(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1),$
Here, the expectation is taken with respect to the distribution of $P(G_{obs}|Z_{obs})$.
Proof.
Refer to Appendix B.2. ∎
As a corollary, one can take all the clusters to be of size one, implying every unit is a cluster. This retrieves the setup for a completely randomized experiment, and thus, Theorem 2 also holds true for a completely randomized experiment.
We emphasize that the generation of graphs with a given degree sequence admits a polynomial time algorithm. Given fixed parameters like the maximum degree of the graph, one can use state-of-the-art algorithms to efficiently sample graphs with a given degree sequence uniformly from the space of all the graphs with the same degree sequence (Arman et al. (2021)). It has been shown that the size of this equivalence class can be exponentially large (Barvinok and Hartigan (2013)). It should be noted that one should use a uniform sampler in order to avoid introducing any bias in the null distribution attributed to sampling. We can also overcome this by taking a large number of samples to approximate the null distribution effectively.
4.3   Test statistic
The choice of the test statistic is not unique, and the validity of the proposed test is upheld for any choice of test statistic. Test statistics capture departures from the null distribution, and better-suited test statistics in responding to deviations from the null will improve the power of the test. Here, we propose a variation of the Has-Treated-Neighbour test statistic considered in Athey et al. (2018). This looks at the difference in average treatment effect amongst control units with at least one treated neighbor and control units with only control units as neighbors. Essentially, we treat focal units as control units and auxiliary units as treated units.
$T_{I_{c}}(G|Z_{obs})$ $=$ (4)
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)},$
Note that the above test statistic considers the spillover on the control units from the treated. Under the null of no spillover effect for all the units regardless of direct treatment status, one can similarly define the above test statistic for the treated units.
$T_{I_{t}}(G|Z_{obs})$ $=$ (5)
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)},$
We now define the test statistic, denoted by $T_{I}(G|Z_{obs})$, for testing the spillover effect for all the units. It will be a weighted average of the two test statistics defined above.
$T_{I}(G|Z_{obs})=\frac{N_{c}}{N}\cdot T_{I_{c}}+\frac{N_{t}}{N}\cdot T_{I_{t}}$ (6)
It should be noted the above test statistics can only be defined if the sub-populations in consideration are non-empty. For example, if there exists no control unit with all neighbors in control or the number of such units is few, we deal with either a definition or power issue. Thus, we propose a generalization of the above-stated test statistic for robustness. We look at the difference in average treatment effect among control units with greater than 0.75 quantile of the proportion of treated units and control units with less than 0.25 quantile of the proportion of treated units.
$T_{quant_{c}}(G|Z_{obs})$ $=$ (7)
$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\leq p_{0.25})}{|P_{0.25}|},$
Here, $p_{0.25}$ and $p_{0.75}$ represent the first and the third quartile of the vector
$P:=\left\{\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j}}{\sum_{j=1}^{N}G_{ij}}:Z_{i}=0\quad\forall i\in[N]\right\}.$
$P_{0.25}$ and $P_{0.75}$ represent the set of elements in $P$ in the first and third quartile, respectively. In a similar fashion to the test statistic $T_{I}(G|Z_{obs})$, we define the above for treated units and take a weighted average for the two test statistic to obtain $T_{quant}(G|Z_{obs})$. We remark that one can improve the test statistic with the knowledge of appropriate covariates or nodal attributes. This can be incorporated by considering regression-adjusted test statistics taking residuals for the potential outcome (Rosenbaum (2002)). One can also build test statistics based on models for interference structures. For example, parametric test statistics can be constructed of the linear-in-means model of interference proposed by Manski (1993). See Athey et al. (2018) for a more elaborate discussion on this.
5   Simulation study
We present Monte Carlo simulations to validate our proposed testing procedure. We consider the test statistic based on the quantile of the proportion of treated neighbors in the simulations, as discussed in Section 4.3. We will describe the data-generating process and the general setup in the next section. We conduct the simulation study for two experimental designs, namely, completely randomized experiments and cluster randomized experiments. As a robustness check, we repeat the study for two network-generating processes.
Table 1: Rejection rates for the null hypothesis of no spillover effects for completely randomized experiment
Network Distribution Direct Effect Spillover Effect Degree Permutation Test Conditional Focal Test
$T_{quant}$ $T_{elc}$
Small world network 0 0 0.049 0.050
4 0 0.046 0.046
0 0.4 0.106 0.112222We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
4 0.4 0.101 0.060
Stochastic block model 0 0 0.062 0.054
4 0 0.055 0.052
0 0.4 0.130 0.065
4 0.4 0.140 0.048
5.1   Data generating mechanism
We now present the data-generating process considered in the simulations. Our potential outcome function is linearly separable in direct treatment effect and spillover effect. The spillover mechanism is a function of the unit’s neighbor. We propose the mechanism to be contingent upon the proportion of treated neighbors.
$Y_{i}(\mathbf{z})=\begin{cases}\tau_{\text{direct}}\cdot z_{i}+\tau_{\text{spill}}\cdot\frac{\sum_{j=1}^{N}G_{ij}\cdot z_{j}}{\sum_{j=1}^{N}G_{ij}}+\epsilon_{i}&\sum_{j=1}^{N}G_{ij}>0,\\ \tau_{\text{direct}}\cdot z_{i}+\epsilon_{i}&\text{otherwise}.\end{cases}$ (8)
We note that $\epsilon_{i}$ can be viewed as the baseline level of outcomes for all the units if they were untreated. Without loss of generality, we take the mean of the normal distribution to be zero, but it can be arbitrarily chosen. We first present the setup for a completely randomized experiment. The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and assigning them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). We state that $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have $599$ units and $300$ treated units in the model in one case. We study the setup for two generating processes, namely, small-world networks and the stochastic block model. Both models exhibit a high degree of clustering, which is characterized by the clustering coefficient. The clustering coefficient is the number of connections in a randomly chosen node and its neighborhood relative to the total number of connections possible in the neighborhood. Consider a regular graph, which is a graph where all the nodes have the same number of connections. Let us denote the number of connections a unit has in a regular graph by $K$. Such graphs have a high clustering coefficient but also distance between two nodes in such graphs is also high. It is observed in real-life networks that the average distance between two nodes is low (Amaral et al. (2000)). To obtain a graph with a high degree of clustering but a low average path length, one can rewire the edges of a $K$-regular graph with some probability, say $p$. This parametrization characterizes a small world network. The higher the probability of rewiring, the lower the average path length gets. Another class of random graph models that achieves the same property of a high clustering coefficient but a low average path length is that of the stochastic block model. Consider a graph of size $N$ with a partition of nodes. The size of the parts are $\{n_{1},n_{2},...,n_{K}\}$. These parts represent clusters in the graph, and we can model a random graph using it by defining the probability of an edge between nodes within a part and between two parts. Hence, we obtain a probability (symmetric) matrix of size $K\times K$. Note that because we want a high degree of clustering in the graph, the diagonal elements of the probability matrix are typically higher than the off-diagonal elements. The probability matrix is also referred to as the preference matrix in the literature. We now define the setup for the simulation study:
  1. 1.
    Network: We first present results taking network distribution as a small-world network with the following parameters: K (initial connection of all units)= 10 and $p_{rw}$ (probability of rewiring) = 0.1. We then consider a stochastic block model with block sizes $\{50,100,40,110,299\}$. The preference matrix which specifies the probability of an edge between two groups is given below:
    0.08 0.01 0.01 0.01 0.01
    0.05 0.01 0.01 0.01
    0.05 0.01 0.01
    0.05 0.01
    0.09
  2. 2.
    Direct treatment effect: We take $\tau_{direct}$ equal to 4 and 0 in our simulations.
  3. 3.
    Spillover effect: We take the spillover effect $\tau_{spill}$ to be 0 and 0.4.
  4. 4.
    Test statistic: We consider the $T_{quant}$ test statistic described in Section 4.3. It takes the difference in the treatment effect of control units in the first quartile of the proportion of treated neighbors to control units in the last quartile of the proportion of treated neighbors. We define the same for treated units and take the weighted average of both.
The null randomization distribution was generated by sampling graphs with the same degree sequence as the observed graph. The null distribution was approximated by generating 1000 graphs with the same degree sequence as the observed graph. The p-value is calculated as the sample proportion of times the imputed test statistic for different graph realizations, with the same degree sequence as the original graph, is more than the observed test statistic. If the p-value obtained is less than the chosen significance level of 0.05, we reject the null hypothesis. We replicated the above-described setup 4000 times and calculated the proportion of times we rejected the null. We present this result, namely rejection rate, in the Table 1. For robustness, we perform the simulations for two network-generating processes. We obtain the rejection rates for the same setup for the procedure proposed by Athey et al. (2018). We implement random focal vertex selection and show our results for the Edge-Level-Contrast test statistic (Athey et al. (2018)).
We then perform the simulations for cluster randomized experimental setup. We present the results in Table 2. The total number of units in the simulation is the same as before, that is, 599. We obtain the clusters in the graph by implementing the epsilon-net clustering algorithm, described in Algorithm 1 (Appendix A). We take epsilon to be equal to 3. This implies all the clusters contain units that are less than 3 hops away from each other, whereas between cluster distance is at least 3 hops for any two units. We take the probability of treating a cluster to be 0.5. We implemented a Bernoulli trial at the cluster level. The null distribution is again approximated by taking 1000 samples of graphs with the same degree sequence as the original graph. We replicated this setup 2000 times and calculated the proportion of times the proposed test rejected the null. We reduce the replication size to 2000 for cluster randomized setup observing appropriate convergence of the p-value estimate. This was performed for two network-generating processes as before.
Table 2: Rejection rates for the null hypothesis of no spillover effects for cluster randomized experiment
Network Distribution Direct Effect Spillover Effect Degree Permutation Test
$T_{quant}$
Small world network 0 0 0.055
4 0 0.055
0 0.4 0.213
4 0.4 0.217
Stochastic block model 0 0 0.053
4 0 0.055
0 0.4 0.168
4 0.4 0.164
5.2   Comparison with related work
We see that our proposed method has type 1 error controlled appropriately. This corresponds to rows with $\tau_{spill}$ equal to zero in Tables 1 and 2. When $\tau_{spill}$ is equal to 0.4, we see that our method performs with competitive power compared to existing methods in the literature (Athey et al. (2018)). We highlight the case of $\tau_{direct}$ equal to 4 and $\tau_{spill}$ equal to 0.4, where our method displays substantial improvement. The significance of the setup lies in its resemblance to many naturally occurring settings, where the direct effect is meaningfully more than the spillover effect (for example, see Cai et al. (2015)). We see that when the clustering of the network-generating process increases, as is the case with the stochastic block model, our proposed method displays a boost in power. This is in contrast with the method proposed by Athey et al. (2018), where the method does not retain power. An interesting result is in the case of cluster-randomized experiments where a higher degree of clustering shows a drop in power. This can be observed when comparing the power values of the small-world network to the stochastic block model. This can be attributed to a reduction of sample space of imputed values of the test statistic. Hence, a higher degree of clustering increases the cluster sizes, which restricts the direct treatment assignment. The observed graph in the case of the small-world network had 16 clusters, as compared to 11 clusters in the case of the stochastic block model. We also perform the simulation study for cluster randomized experiments for the methods proposed in Athey et al. (2018) and Puelz et al. (2022). As discussed in Section 3, we encounter the degenerate null distribution and hence, do not present the corresponding results. This is more prominent in the method proposed by Athey et al. (2018), in comparison with Puelz et al. (2022). Since Puelz et al. (2022) takes into account the design of the experiment when selecting focal units, we see more robustness of the biclique test (Puelz et al. (2022)) where degenerate null distributions are reduced in frequency. Overall, our method stands more robust to degeneracy and performs well in terms of power in our simulations.
6   Application: Weather insurance adoption in rural China
We re-analyze a field experiment conducted in rural China on rice crop farmers studying neighborhood interference effects (Cai et al. (2015)). The research motive behind the experiment was to see if social networks among the farmers have any influence on weather insurance adoption. The experiment involved conducting intensive training sessions to promote crop insurance products to farmers against extreme weather. A primary goal of the experiment was to identify if there is information dissemination in the farmers’ community and, if so, what mechanisms drive these information exchanges. These are important questions to be answered in this setting as they are consequential to government policy action. A deeper understanding of insurance product adoption via social network mechanisms helps devise a more cost-effective government policy campaign to enhance product adoption among the target community. Once interaction mechanisms can be established, better “seeding” or targeting strategies can be rolled out for future campaigns (Ok et al. (2014)). Such strategies identify influential figures in the community to maximize the interference effect and, hence, the program’s overall impact. Another important reason to study interference in such a setup is to avoid biased estimates and accurately evaluate policy performance measures.
We discuss the experimental design setup by the researchers. The hypothesis of interest is the underlying assumption that access to better information about weather insurance products would increase the adoption of the product in the rice farmers’ community. To capture access to information differentially as treatment and control, two types of information sessions were created: simple sessions of 20 minutes, which provided an introduction to the insurance contract, and intensive sessions of 45 minutes that covered the insurance contract in much more detail, including explaining to them what the benefits are of taking up insurance. To capture information dissemination across the social network, these sessions were conducted in two rounds with a gap of 3 days. The idea behind this setup is to allow the farmers some time to discuss the insurance product among themselves. The interference effect is identified by examining whether households in round 2 with more friends who received intensive sessions in round 1 displayed an enhanced insurance take-up rate. Second-round participants were further randomized into three groups. After the round 2 sessions, the first group was given no additional information, the second group was told about the overall take-up rate in the previous round, and the third group was told in detail about who purchased the product and who did not. This additional information randomization of the second group was done to derive the main mechanism via which information diffusion took place. For illustration purposes, we do not analyze this part of the experiment.
There were a total of 4902 households across 47 different villages in the experiment. The social network among a village is known and is assumed to be independent between two distinct villages. Each household is randomized among the 4 different groups: (i) Receiving simple session in Round 1,(ii) Receiving intensive session in Round 1, (iii) Receiving simple session in Round 2, (iv) Receiving intensive session in Round 2. After the session ends, the households make the decision to purchase insurance. This is captured as a binary outcome variable of the model, with 1 indicating a purchase decision made and 0 when no purchase is made. Under the assumption that no information diffusion can take place between participant households of Round 1 because of the immediate purchase decision the household had to make, we can assume there are no edges among the household participants in Round 1.
In our analysis, we take all the households receiving simple sessions under the control group and intensive sessions under the treated group. Since spillover can only occur from the treated group of Round 1 to the control group of Round 2, we only consider the network structure amongst these two groups in our study. We retain the network structure among the households between the two rounds and nullify the remaining network structure. This includes network structure among households in Round 1 only and network structure among households in Round 2 only. Our analysis was run on 433 households of randomly chosen villages in the experiment, of which 222 were in the treated group. We ran our proposed Procedure 2 and obtained a p-value of 0.0002, which indicates the presence of interference.
7   Conclusion
In this paper, we develop a new methodology to test for the non-sharp null hypothesis of no spillover effect in an experimental setup. Our proposed method assumes mild distributional conditions in the network-generating process to build a quasi-randomization test. Whilst promising, there are several open problems that remain. Incorporating covariates or nodal attribute information to better characterize the random graph null model may help enhance the power of the test and is an interesting direction for future work. We also assume fully and correctly specified information of the given network. An analysis of the random graph null models under partial information and network misspecification is another interesting avenue for future work. Finally, our focus in this paper has been on randomized experiments exclusively. Since the sampling of null distribution is based on quasi-randomization of the network-generating process, a natural and important extension is to study random graph null testing procedures in observational studies.
Acknowledgments
We thank Sumit Kunnumkal, Parshuram Hotkar, and ISB seminar participants for their helpful comments and feedback. We gratefully acknowledge Hemanth Kumar and the ISB Institute of Data Science for providing cluster access. Pallavi Basu’s research is partially supported by the SERB MATRICS award MTR/2022/000073. There is no conflicting interest to declare.
Appendix
Appendix A Algorithm: $\epsilon$-net cluster
An $\epsilon$-net in the graph is a set of units of the graph such that any two units in the set are at least $\epsilon$ distance away from each other, and any unit outside the set is within $\epsilon$ distance. Here, distance represents the shortest path length between two units in the graph. To form an $\epsilon$-net cluster, one can pick a vertex and remove all the vertices within $\epsilon$ distance of the vertex. Then, pick another vertex and iteratively perform the same process until you exhaust all the vertices. The resultant set will be an $\epsilon$-net. We present an algorithm below to obtain an $\epsilon$-net cluster.
Input:
  • Given network $G_{obs}$
Procedure:
$S$ is equal to $G_{obs}$
while $S$ is not empty do
  • Step 1: Pick the first vertex in $S$, say $v_{1}$
  • Step 2: Find all units within $\epsilon$ distance of $v_{1}$
  • Step 3: Store the units found in Step 2 as a cluster and remove these units from $S$
end while
Store the resultant cluster outputs in $C$
Output: C
Algorithm 1 $\epsilon$-net cluster
Appendix B Proofs
B.1   Proof of Theorem 1:
Proof.
We define p-value as
$p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs}),$
Given $Z_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution as induced by $P(G|Z_{obs})$. Thus,
$p_{Z_{obs}}=1-F_{U}(G_{obs}|Z_{obs}).$
Here, $F_{U}(.)$ represents the cumulative distribution of $U$. Then,
$T_{c}(G|Z_{obs})=T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0},$
Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z_{obs})$ and we get
$p_{Z_{obs}}=1-F_{U}(U),$
Hence, by using the probability integral transform theorem, we obtain
$P(p_{Z_{obs}}\leq\alpha)\leq\alpha.$
B.2   Proof of Theorem 2:
Proof.
We define p-value as
$p_{Z_{obs},G_{obs}}$ $=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z=Z_{obs},G\in[G_{obs}]),$
$p_{Z_{obs}}$ $=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z=Z_{obs}).$
Given $Z_{obs}$ and $[G_{obs}]$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution as induced by $P(G|Z=Z_{obs},G\in[G_{obs}])$. Thus,
$p_{Z_{obs},G_{obs}}=1-F_{U}(G_{obs}|Z=Z_{obs},G\in[G_{obs}]),$
Here, $F_{U}(.)$ represents the cumulative distribution of $U$. Then,
$T_{c}(G|Z_{obs})=T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0},$
Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z=Z_{obs},G\in[G_{obs}])$ and we get
$p_{Z_{obs},G_{obs}}=1-F_{U}(U),$
Hence, by using the probability integral transform theorem, we obtain
$P(p_{Z_{obs},G_{obs}}\leq\alpha)\leq\alpha.$
Let $P_{1},P_{2},...,P_{d}$ be the partition induced by $\equiv_{d}$. Pick $G_{1},G_{2},...,G_{d}$ such that $G_{i}\in P_{i}$ for all $i=1,2,...,d$. Hence,
$P(p_{Z_{obs},G_{i}}\leq\alpha)$ $\leq\alpha\qquad i\in[d],$
$\sum_{i\in[d]}P(p_{Z_{obs},G_{i}}\leq\alpha)\cdot P(G\in[G_{i}]))$ $\leq\sum_{i\in[d]}\alpha\cdot P(G\in[G_{i}]),$
$P(p_{Z_{obs}}\leq\alpha)$ $\leq\alpha.$
We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}])$ is a uniform distribution under a cluster randomized experiment. Consider $g,g^{\prime}\in[G_{obs}]$.
$P(G=g|Z=Z_{obs},G\in[G_{obs}])=\frac{P(Z=Z_{obs}|G=g,G\in[G_{obs}])\cdot P(G=g|G\in[G_{obs}])}{P(Z=Z_{obs}|G\in[G_{obs}])},$ (9)
$P(G=g|G\in[G_{obs}])$ $=\frac{f_{p_{0},p_{1},...,p_{N-1}}(deg(g))}{\sum_{G\in[G_{obs}]}f_{p_{0},p_{1},...,p_{N-1}}(deg(G))},$ (10)
$=\frac{f_{p_{0},p_{1},...,p_{N-1}}(deg(g^{\prime}))}{\sum_{G\in[G_{obs}]}f_{p_{0},p_{1},...,p_{N-1}}(deg(G))},$
$=P(G=g^{\prime}|G\in[G_{obs}]).$
$P(Z=Z_{obs}|G=g,G\in[G_{obs}])$ $=P(\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\hskip 5.69054ptj\in[N]|G=g,G\in[G_{obs}]),$ (11)
$=P(W=W_{obs}|G=g,G\in[G_{obs}]),$
$=P(W=W_{obs})$
$=P(W=W_{obs})|G=g^{\prime},G\in[G_{obs}]),$
$=P(Z=Z_{obs}|G=g^{\prime},G\in[G_{obs}]).$
Note that we take cluster partition to be fixed as part of the experimental design. Using Equation (10) and (11) in Equation (9), we obtain
$P(G=g|Z=Z_{obs},G\in[G_{obs}])=P(G=g^{\prime}|Z=Z_{obs},G\in[G_{obs}]),$
Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}]\}$ is uniform under cluster randomized experiment. Hence,
$p_{Z_{obs},G_{obs}}=\frac{1}{|[G_{obs}]|}\cdot\sum_{G\in[G_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})).$
Appendix C Additional numerical simulations
We present the data-generating process considered in the simulations. To give an overview, we consider a linearly separable potential outcome model for the outcome function, which is a composition of two effects: a direct treatment effect and a spillover treatment effect. The spillover parameter is weighted by a function of the unit’s neighbor. We consider two such outcome models for the simulations: one that captures spillovers by the presence of a treated neighbor and the other that captures it via the proportion of treated neighbors.
C.1   Data generating mechanism
The following describes the two potential outcome models:
$Y_{i}(\textbf{z})=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}>0)+\epsilon_{i},$ (12)
$\epsilon_{i}\sim\mathcal{N}(0,1)\qquad\forall i\in[N].$
Here, $\epsilon_{i}$ represents the baseline level of outcomes without any treatment assignment. We take the mean of the normal distribution to be zero without loss of generality. We now re-state the second potential outcome model from Section 5.
$Y_{i}(\mathbf{z})=\begin{cases}\tau_{\text{direct}}\cdot z_{i}+\tau_{\text{spill}}\cdot\frac{\sum_{j=1}^{N}G_{ij}\cdot z_{j}}{\sum_{j=1}^{N}G_{ij}}+\epsilon_{i}&\sum_{j=1}^{N}G_{ij}>0,\\ \tau_{\text{direct}}\cdot z_{i}+\epsilon_{i}&\text{otherwise}.\end{cases}$ (13)
The treatment assignment vector is chosen by randomly sampling $N_{t}$ units out of $N$ units and giving them treatment (that is, $Z_{i}=1$ for all $i$ corresponding to units in the random sample). Here, $N_{t}$ is fixed and pre-determined by the experimenter. In the performed simulations, we have 100 units and 5 treated units in the model in one case. We then extended the study to 500 units, and 250 were treated in the second case.
Consider the following setup:
  1. 1.
    Network: We work with the standard Erdős-Rényi random graph model $G(n,p)$ where $n$ is the number of nodes and $p$ is the probability of edge formation between two nodes. In our setup, we take $p=0.2$.
  2. 2.
    Direct treatment effect: We take $\tau_{direct}=4$ in our simulations. For robustness, we also perform the simulations with $\tau_{direct}=1$. It is noted that $\tau_{direct}$ has no impact on the validity and power of the test.
  3. 3.
    Spillover effect: We vary the spillover effect, $\tau_{spill}$ from 0 to 7, keeping $\tau_{direct}$ fixed.
  4. 4.
    Test statistic: We consider the test statistic taking the difference of treatment effect on control units with control neighbors and control units with treated neighbors, $T_{I_{c}}(G|Z_{obs})$. We also consider the test statistic based on the difference in treatment effect of a low proportion of treated neighbors and a high proportion of treated neighbors, $T_{{quant}_{c}}(G|Z_{obs})$. Refer to Section 4.3 for further details.
We generate the randomization distribution of the test statistic in consideration by repeated drawing from the Erdős-Rényi random graph model with the probability of an edge between two nodes being 0.2. The simulations in Table 3 were carried out on 100 units, with 5 being treated and the rest in control. We approximated the randomization distribution of the test statistic by taking 200 samples of the graph. We calculate the corresponding p-value by the proportion of times the test statistic imputed was larger than the test statistic for the observed network. If the p-value is less than 0.05, we reject the null hypothesis of no spillover effects. We replicated the above 480 times and calculated rejection rates as the proportion of times the p-value was significant. For this setup, the outcome model (12) with $\tau_{direct}$ equals to 4 and $\tau_{spill}$ is presented in Table 3.
We then repeat the simulations for the above setup for outcome model (13) and present them in Table 4.
We note that the power of the test decreases substantially for model (13), compared with model (12). This is owing to the simplistic nature of the outcome model (12), and we present its results for verification purposes. Another caveat in the above setting is the nature of test statistic $T_{I_{c}}(G|Z_{obs})$. Depending on the input graph, we might not get any control units with all neighboring ones in control, invalidating the procedure. So, the above setup involves 100 units with 5 units in treatment. We numerically found the procedure failing when the number of treated units was increased to more than 5. We present the results for model (13) with $T_{{quant}_{c}}$ as test statistic in Table 4. Here, the number of units in the population is 500, with the number of treated units being 250. $\tau_{direct}$ is 4 and varying $\tau_{spill}$ is used for power performance measure.
In all the above models, we took the probability of an edge between two nodes as $0.2$. We present the same setup in 4, with the probability of an edge between two nodes estimated as below. $G_{obs}$ is generated from Erdős-Rényi random graph with $p=0.2$.
$\hat{p}=\frac{|E_{G_{obs}}|}{\binom{N}{2}}.$
Here, $|E_{G_{obs}}|$ represents number of edges in the observed graph $G_{obs}$. The results are presented in Table 4.
We highlight how type 1 error is controlled under 0.05 in all the setups (case when $\tau_{spill}$ is zero). When the null hypothesis is not true, we observe good power against the test in line with results from Athey et al. [2018]. This remains true when the network distribution is estimated parametrically.
Table 3: Rejection rates for null hypothesis of no spillover effects; Outcome model (12) with test statistic $T_{I_{c}}$
$\tau_{spill}$ Rejection Rate
0.0 0.035
0.1 0.083
0.2 0.173
0.3 0.279
0.4 0.402
0.5 0.579
0.6 0.745
0.7 0.887
0.8 0.943
0.9 0.987
1.0 0.993
Table 4: Rejection rates for null hypothesis of no spillover effects; Outcome model (13) with test statistic $T_{I_{c}}$, $T_{{quant}_{c}}$ and $T_{{quant}_{c}}$ with estimated Erdős-Rényi random graph
$\tau_{spill}$ Rejection Rate
$T_{I_{c}}$ $T_{{quant}_{c}}$ $T_{{quant}_{c}}$ with $\hat{p}$
0 0.054 0.040 0.054
1 0.058 0.080 0.069
2 0.110 0.238 0.229
3 0.181 0.419 0.419
4 0.256 0.694 0.693
5 0.365 0.881 0.881
6 0.485 0.954 0.954
7 0.612 0.990 0.994

### v2

@@
 Abstract
-Many classical inferential approaches fail to hold when interference exists among the population units.
-This amounts to the treatment status of one unit affecting the potential outcome of other units in the population.
-Testing for such spillover effects in this setting makes the null hypothesis non-sharp.
+Network interference amounts to the treatment status of one unit affecting the potential outcome of other units in the population.
+Testing for spillover effects in this setting makes the null hypothesis non-sharp.
 An interesting approach to tackling the non-sharp nature of the null hypothesis in this setup is constructing conditional randomization tests such that the null is sharp on the restricted population.
-In randomized experiments, conditional randomized tests hold finite sample validity.
 Such approaches can pose computational challenges as finding these appropriate sub-populations based on experimental design can involve solving an NP-hard problem.
@@
 We propose a new approach that builds a conditional quasi-randomization test.
-Our main idea is to build the (non-sharp) null distribution of no spillover effects using random graph null models.
-We show that our method is exactly valid in finite-samples under mild assumptions.
-Our method displays enhanced power over other methods, with substantial improvement in complex experimental designs.
-We highlight that the method reduces to a simple permutation test, making it easy to implement in practice.
-We conduct a simulation study to verify the finite-sample validity of our approach and illustrate our methodology to test for interference in a weather insurance adoption experiment run in rural China.
-Keywords: Interference, Networks, Causal Inference, Quasi-Randomization Tests, Peer Effects, Random Graphs
+We build the (non-sharp) null distribution of no spillover effects using random graph null models.
+We show that our method is exactly valid in finite samples under mild assumptions.
+Our method displays enhanced power over other methods, substantially improving cluster randomized trials.
+We illustrate our methodology to test for interference in a weather insurance adoption experiment run in rural China.
+Keywords: Causal Inference, Network Interference, Peer Effects, Randomization Tests, Random Graphs
 1 Introduction
 A common assumption in causal inference is that of the Stable Unit Treatment Value Assumption (SUTVA).
-This assumes away any dependence of one unit’s treatment status on the outcome of another unit.
-Whilst this allows for unbiased treatment effect estimation and valid inferential strategies, there are many experimental and observational settings where interference between units takes place (Cox (1958)).
+SUTVA assumes no dependence of a unit’s outcome on another unit’s treatment status.
+While this allows for unbiased treatment effect estimation and valid inferential strategies, there are many experimental and observational settings where interference between units takes place (Cox (1958)).
 There are multiple pathways via which interference can take place in the population.
-It can be a result of the spread of treatment amongst the population units, the respective outcomes of treated units affecting the outcomes of others, or a combination of both.
-For example, the spread of a contagious disease is contingent upon the number of infected people in the population (Halloran and Struchiner (1995)).
+It can result from the spread of treatment amongst the population units, the respective outcomes of treated units affecting the outcomes of others, or a combination of both.
+For example, the spread of a contagious disease depends on the number of infected people in the population (Halloran and Struchiner (1995)).
 Thus, the estimation of any preventive measures taken will be challenged because of a likely SUTVA violation.
@@
 (2019); Blake and Coey (2014)).
-The total treatment effect, in such cases, can be decomposed into two effects, one of the direct effect and the other of spillover effects (Sobel (2006)).
+The total treatment effect, in such cases, can be decomposed into two effects, one of the direct effects and the other of spillover effects (Sobel (2006)).
 The classical method for testing treatment effects is the “Fisher Randomization Test” or FRT (Fisher (1936)).
-It is used to test for a sharp null hypothesis of no treatment effect for all units in the population.
+FRT is used to test for a sharp null hypothesis of no treatment effect for all units in the population.
 The test exploits the imputability of counterfactuals under the null.
-This enables the determination of the randomization null distribution of a test statistic of choice.
+Imputation enables the determination of the randomization null distribution of a test statistic of choice.
 If one were to test for spillover effects, it would render the null non-sharp, leading to imputability issues for the test statistic.
 There is a growing body of literature that addresses this issue by considering a carefully determined subset of units in the population and treatment assignment vector, such that on these selected units and treatment vectors, the null is sharp (Aronow and Samii (2017); Athey et al.
-(2018); (Basse et al., 2019b)).
-Such conditional randomization tests also lead to valid exact p-values, as is the case with the Fisher Randomization Test.
+(2018); Basse et al.
+(2019b)).
+Such conditional randomization tests also lead to valid exact p-values, as with the Fisher Randomization Test.
 Basse et al.
-(2019b) build a permutation test under the special case of clustered interference.
-Whilst a permutation test is computationally cheaper, the method lacks generalizability because of the structured nature of interference considered.
+(2019b) build a permutation test under the particular case of clustered interference.
+While a permutation test is computationally cheaper, the method lacks generalizability because of the structured nature of the interference considered.
 Aronow and Samii (2017) and Athey et al.
-(2018) generalize the above for a given network and consider a range of non-sharp null hypotheses for testing.
+(2018) generalize the above for a given network and evaluate a range of non-sharp null hypotheses for testing.
 Puelz et al.
 (2022) presents a graph-theoretic approach to constructing the conditioning mechanism of a sub-population of units and the corresponding treatment assignment vectors for which the null is sharp.
-However, the proposed methodology is NP-hard in implementation and, thus, in practice, does not involve observing the treatment assignment to construct the conditioning mechanism.
-Whilst the above considers the problem of testing for interference in an experimental setting, the literature on testing for spillover effects in an observational setting is still in its infancy.
-In this paper, we consider the network to be coming from a distribution itself.
-This is in contrast with the setup considered in the current methods that assume the network to be fixed.
+However, the proposed methodologies lose power in cluster randomized experiments.
+This paper considers the network to be coming from a distribution itself.
+This idea contrasts with the setup considered in the current methods, assuming the network is fixed.
 We draw a parallel here to estimating the probability of unit treatment assignment via propensity scores in the case of observational studies.
-The propensity of two nodes in a network being connected can be looked at as being a function of their network characteristics, covariates, or both.
+The propensity of two nodes in a network being connected can be considered a function of their network characteristics, covariates, or both.
 This leads us to think of these linkages as coming from a generating process with an associated probability distribution.
 The current methods are based on strict randomization inference with no distributional assumptions for either the network or the potential outcomes, achieving a generalization of the Fisher Randomization Test.
-Whilst this allows for widespread validity of the methods proposed, it comes at a cost of loss of power.
-We emphasize that leveraging appropriate stochastic assumptions can significantly boost power and lead to methodological advancements.
-In this paper, we restrict the network data-generating process to be characterized by any arbitrary degree distribution.
-First proposed in Newman et al.
+This allows for the widespread validity of the methods proposed, albeit it might be at the cost of power.
+This paper explores how leveraging appropriate stochastic assumptions can significantly boost power and lead to methodological advancements.
+We assume that a given vector of sufficient statistics characterizes the network data-generating process and can thus be represented as an exponential random graph model.
+We first take the degree sequence of the graph as a sufficient statistic.
+Proposed in Newman et al.
 (2001), random graphs with arbitrary degree distribution represent real-world social networks closely owing to their ability to capture heterogeneity in degree.
-Given the mild nature of the distributional assumption, we obtain a large class of network data-generating processes.
-This assists us in building a quasi-randomization test based on the network distribution.
-We show that conditional on the observed treatment, the (quasi) randomization distribution of a test statistic leads to valid p-values.
-The challenge of the imputability of a test statistic of choice under the null, being the major methodological contribution of the current methods in the literature, is naturally overcome in our method.
-Model-based approaches with distribution assumptions on outcome model function have been previously considered in the literature for identifying and estimating peer effects (Manski (1993); Manski (2013); Bowers et al.
+This assists us in building a quasi-randomization test based on the network distribution by randomizing over graphs from this generating process.
+We show that conditional on the observed treatment, a test statistic’s (quasi) randomization distribution leads to valid p-values.
+We then generalize the model to a vector of sufficient statistics and present a method that works for any set of such network statistics.
+Model-based approaches with distribution assumptions on outcome model function have been previously considered in the literature for identifying and estimating peer effects (Manski (1993, 2013); Bowers et al.
 (2013); Toulis and Kao (2013); Blume et al.
@@
 More recently, Li and Wager (2022) and Leung (2020) proposed estimation strategies for spillover effects under a distributional assumption on the network formation model.
-The inferential methods for the proposed estimation procedures involve either non-standard procedures or asymptotic approximations (Imbens and Manski (2004)).
+The inferential methods for the proposed estimation procedures involve non-standard or asymptotic approximations (Imbens and Manski (2004)).
 With a careful identification of causal estimates, such large sample normal approximations of confidence interval conform to a Neymanian perspective.
@@
 Building upon this, we develop a randomization-based test for a given random graph model.
-This leads to finite sample validity of the p-value obtained for the oracle method, in line with Fisherian procedures.
-Our method addresses several technical challenges encountered in the current methods.
+This leads to finite sample validity of the p-value obtained, in line with Fisherian procedures.
+As a second contribution, our method addresses several technical challenges encountered in the current methods.
 We show that randomization tests based on network-generating processes lead to imputable test statistics under the null, conditional on the treatment assignment.
-This overcomes a major challenge faced in randomization methods based on random treatment assignment.
-The current methodology involves restricting attention to a carefully chosen subset of focal units and claims null of no spillover effects for only these units.
-In contrast, our method tests for a stricter class of nulls with no spillover effects on all the units in line with the Fisher Randomization Test.
-Another hurdle faced by the current methodologies is the extension of the method to complex experimental designs.
+This overcomes a significant challenge in randomization methods based on random treatment assignment.
+The state-of-the-art techniques restrict attention to a carefully chosen subset of focal units and claim null of no spillover effects for only these units.
+In contrast, our method tests for a stricter class of nulls with no spillover effects on all the units.
+Another hurdle the current methodologies face is the loss of power in cluster randomized trials.
 Determination of conditional randomization distribution given a conditioning mechanism is not straightforward.
@@
 It can lead to an enormous loss of power or render the null degenerate completely.
-There is a burgeoning set of literature focussing on improving the current methods by incorporating design-based techniques (Puelz et al.
-(2022); (Basse et al., 2019a)).
-We emphasize that our method is easily generalized to various experimental designs, and we show this for two experimental designs, namely, completely randomized experiments and cluster randomized experiments.
-Our second major contribution is the computational efficiency of the proposed method.
-We propose that the random graph model be characterized by any arbitrary degree distribution.
-We build a permutation test by generating the null over the set of graphs with the same degree sequence as the observed graph.
+A burgeoning literature set focuses on improving current methods by incorporating design-based techniques (Puelz et al.
+(2022); Basse et al.
+(2019a)).
+Unlike distribution-free methods, we show how incorporating information about network generation into the methodology can boost power.
+Our third significant contribution is the computational efficiency of the proposed method.
+We suggest that sufficient network statistics can characterize the random graph model.
+We first take a single sufficient statistic as the network’s degree sequence and build a permutation test by generating the null over the set of graphs with the same degree sequence as the observed graph.
 Random graph generation of a prescribed degree sequence admits fast and efficient polynomial time algorithms for large graphs (Viger and Latapy (2005); Fosdick et al.
 (2018)).
-Thus, via permutation of edges of the graph in a manner that preserves degree sequence, we obtain valid randomization tests.
-This is closely related to literature in random graph null model used in network analysis (Van Koevering et al.
-(2021)) with applications in protein structure discovery and community detection among many (Milenković et al.
+Thus, we obtain valid randomization tests via permutation of edges of the graph in a manner that preserves degree sequence.
+We then present a method for a general exponential random graph model characterized by an arbitrary set of network statistics and build a permutation test by generating graphs in the same isomorphism class under the null.
+This is closely related to the literature on the random graph null model used in network analysis with applications in protein structure discovery and community detection among many (Milenković et al.
 (2009); Sah et al.
 (2014)).
-To the best of our knowledge, we are the first to formally introduce random graph null models as an extension of the Fisher Randomization Test.
-To illustrate our method, we re-analyze a completely randomized experiment in rural China where rice crop farmers are informed about a weather insurance product (Cai et al.
+To illustrate our method, we re-analyze a randomized experiment in rural China where rice crop farmers are informed about a weather insurance product (Cai et al.
 (2015)).
@@
 We display our methodology on a subset of the original experiment performed and obtain significant information spillover among the farmers.
-The setting displays the importance of such spillover effects in evidence-based policymaking under resource-constrained settings.
 The remaining paper is structured as follows.
 In Section 2, we set up notation and formalize the null hypothesis of interest, incorporating network random variable.
-We also describe the present state-of-the-art methodology of using a conditioning mechanism for testing non-sharp null.
 Section 3 presents a randomization test based on a given network-generating process and its application to different experimental designs.
-Section 4 extends this further to random graph models and presents our main methodology, which can be implemented as a permutation test.
-In Section 5, we present extensive numerical simulations to evaluate and compare the power of the proposed test with the current methods in the literature.
-Section 6 and 7 illustrate our methodology as an empirical application and conclude the paper.
-All theorem proofs, algorithms, and additional numerical simulation results are presented in the Appendix.
+Section 4 extends this further to random graph models and presents our central methodology introducing the permutation test.
+In Section 5, we present extensive numerical simulations to evaluate the performance of the proposed test.
+Sections 6 and 7 illustrate our methodology as an empirical application and conclude the paper.
+Proofs, algorithms, and additional numerical results are presented in the Appendix.
 2 Setup
@@
 Here, $\mathbb{Z}$ represents the set of possible treatments to which the population units can be exposed.
-For the purposes of this article, we take =$\mathbb{Z}=\{0,1\}$.
+We take =$\mathbb{Z}=\{0,1\}$ for this article.
 Hence, any treatment assignment vector is an N-dimensional binary vector.
-Consider the treatment vector to be drawn from a probability distribution $P_{Z}:\{0,1\}^{N}\rightarrow[0,1]\subset\mathbb{R}$.
-Let $Z_{obs}$ be the realized outcome drawn from the treatment assignment distribution $P_{Z}$.
 We consider the population $\mathbb{P}$ to be connected via an undirected graph $G=(\mathbb{P},E)$ where $E\subseteq\mathbb{P}\times\mathbb{P}$.
 For distinct units $i$ and $j$ such that $(i,j)\in E$ represents a connection between the two units.
-With some slight abuse of notation, we also represent the adjacency matrix of the graph by $G$.
+With some slight abuse of notation, we also represent the graph’s adjacency matrix by $G$.
 Thus, $G$ is an $N\times N$ binary matrix.
@@
 Here, $\mathcal{N}_{G}(i)$ represents all the neighbors of unit $i$ in graph $G$.
-We may simply write $\mathcal{N}(i)$ if the graph is clear in context.
-We define the total number of neighbors of unit $i$ as the degree of $i$.
-This is denoted by $deg(i)$ and is equal to $|N(i)|$.
+We may write $\mathcal{N}(i)$ if the graph is clear in context.
+We define the total number of unit neighbors $i$ as the degree of $i$.
+This is denoted by $deg(i)$ and equals $|N(i)|$.
 We denote the distance between two units $i$ and $j$ as $dist(i,j)$.
@@
 By convention, we take $dist(i,i)=0$ and $dist(i,j)=\infty$ if no path exists between the two distinct units.
-Consider the graph $G$ to be coming from the probability distribution $P_{G}:\{0,1\}^{N\times N}\rightarrow[0,1]\subset\mathbb{R}$.
+Consider $\mathbb{P^{\prime}}\subseteq\mathbb{P}$.
+We denote $G[\mathbb{P^{\prime}}]$ as the sub-graph induced by $\mathbb{P^{\prime}}$.
+That is, $G[\mathbb{P^{\prime}}]=(\mathbb{P^{\prime}},E^{\prime})$ where $E^{\prime}=\{(i,j)\in\mathbb{P^{\prime}}\times\mathbb{P^{\prime}}|\>(i,j)\in E\}$.
+Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$ where $X_{i}$ is a m-dimensional vector.
+$X_{i}$ is fixed and influences the outcome via the network’s potential outcome function or formation.
+Let $X$ denote the $N\times M$ matrix where each row corresponds to the respective unit’s covariate vector.
+Consider the treatment vector to be drawn from a probability distribution $P_{X}(Z,G):\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow[0,1]$.
+We define the conditional probability distribution $P(Z|G):\{0,1\}^{N}\rightarrow[0,1]\subset\mathbb{R}$ as $P_{Z|G}$.
+Let $Z_{obs}$ be the realized treatment drawn from the treatment assignment distribution $P_{Z|G}$.
+Consider the graph $G$ to be coming from the marginal probability distribution $P(G):\{0,1\}^{N\times N}\rightarrow[0,1]\subset\mathbb{R}$, denoted as $P_{G}$.
 We denote the observed sample from the distribution $P_{G}$ to be $G_{obs}$.
-Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$ where $X_{i}$ is a m-dimensional vector.
-$X_{i}$ is fixed and influences the outcome via the potential outcome function.
-Let $X$ denote the $N\times M$ matrix where each row corresponds to the respective unit’s covariate vector.
 We define the potential outcome function $Y:\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow\mathbb{R}^{N}$ as a real-valued function that takes treatment assignment and network adjacency matrix as arguments.
@@
 The assignment mechanism defines the joint probability distribution of treatment assignment vectors given covariates and potential outcomes, denoted by $P(Z|X,Y(0),Y(1))$.
-The following are three key assumptions that are considered in the causal inference literature for assignment mechanism: (i) individualistic assignment, which restricts the treatment assignment of a unit to its own covariates and outcomes (ii) probabilistic assignment, which assumes non-integer probabilities of treatment assignment to units (iii) unconfoundedness assumption, which restricts the dependence of the probability assignment of units to only covariates, and not the potential outcome (Imbens and Rubin (2015)).
-Individualistic assignment mechanism assumption, given a well-defined potential outcome function, constitutes Stable Unit Treatment Value Assumption (SUTVA).
+The following are three key assumptions that are considered in the causal inference literature for assignment mechanism: (i) individualistic assignment, which restricts the treatment assignment of a unit to its covariates and outcomes (ii) probabilistic assignment, which assumes non-integer probabilities of treatment assignment to units (iii) unconfoundedness assumption, which restricts the dependence of the probability assignment of units to only covariates, and not the potential outcome (Imbens and Rubin (2015)).
+Given a well-defined potential outcome function, individualistic assignment mechanism assumption constitutes Stable Unit Treatment Value Assumption (SUTVA).
 We argue that the underlying network structure among the population is an implicit form of treatment.
 Hence, we generalize the treatment to a bivariate treatment assignment vector $(Z,G)$.
-Whilst the usual assumptions described above are plausible for the marginal distribution of direct treatment $Z$, the individualistic assumption is not valid for the marginal distribution of network treatment $G$ given the nature of the variable.
-We look at the distribution of joint treatment assignment vector $(Z,G)$ in this setup.
-$P(Z,G|X)=P(G|Z,X)\cdot P(Z|X),$
-assuming joint treatment assignment is unconfounded, that is,
+We assume that the joint treatment assignment is unconfounded; that is,
 $P(Z,G|X,Y(0),Y(1))=P(Z,G|X).$
-We remark that we took the assumption of the unconfoundedness of both treatment assignments into consideration.
 We formally define the standard sharp null hypothesis of no (direct) treatment effect when the population is not interconnected.
 Hypothesis 1.
-Sharp null hypothesis of no treatment effect:
+The sharp null hypothesis of no treatment effect:
 $\forall i\in[N]\quad Y_{i}(Z)=Y_{i}(Z^{\prime})\qquad\forall\quad Z,Z^{\prime}\in\{0,1\}^{N}.$
-The above stated null hypothesis imposes specified forms of restrictions on the potential outcome function. Extending this to our setup, we get the following definition of the sharp null hypothesis.
+The above-stated null hypothesis imposes specified restrictions on the potential outcome function. Extending this to our setup, we get the following definition of the sharp null hypothesis.
 Hypothesis 2.
-Sharp null hypothesis of no (joint) treatment effect:
+The sharp null hypothesis of no (joint) treatment effect:
 $\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}.$
-We now give a definition of a test statistic that will be used for testing such null hypotheses. Following our setup, we will use the joint treatment assignment, $(Z,G)$, for the rest of the article.
+We now define a test statistic that will be used for testing such null hypotheses. Following our setup, we will use the joint treatment assignment, $(Z,G)$, for the rest of the article.
 Definition 1.
 A test statistic $T$ is a real-valued function of the treatment assignment and $Y_{obs}$ given the covariates $X$, $T_{X}(Z,G,Y_{obs})$.
-Under sharp null, given an observed outcome, $Y_{obs}$, we can impute the outcome function $Y$ for all treatment assignments.
+Under the sharp null, given an observed outcome, $Y_{obs}$, we can impute the outcome function $Y$ for all treatment assignments.
 This is not possible when the null hypothesis is non-sharp.
@@
 $\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\textbf{Z}_{i}\times\textbf{G}_{i}\subset\{0,1\}^{N}\times\{0,1\}^{N\times N}.$
-Here, the subset restriction, denoted by $\textbf{Z}\times\textbf{G}:=\textbf{Z}_{i}\times\textbf{G}_{i}$, on treatment assignment is chosen in accordance with the null hypothesis of interest.
-We elaborate further using the example of testing null of no spillover effects below.
+Here, the subset restriction, denoted by $\textbf{Z}\times\textbf{G}:=\textbf{Z}_{i}\times\textbf{G}_{i}$, on treatment assignment is chosen by the null hypothesis of interest.
+We elaborate further using the testing null of no spillover effects example below.
 Consider the null hypothesis of no spillover effects if the distance between two units is at least k.
@@
 (Null Hypothesis: no spillover effects at distance k and beyond) Let the $k-$distance function $D_{G}(i)$ for a unit $i$ in graph $G$ be defined as below:
+$D_{G}^{k}(i)=\begin{cases}1&\text{if dist(i,j)}\leq k,\\ 0&\text{otherwise}.\end{cases}$
+$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
 […]
-$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
-[…]
-Note the non-imputability of potential outcome function under the null considered above for different treatment assignments.
-Specifically, given $(Z_{obs},G_{obs})$, we can compute potential outcomes for treatment assignments in the restriction set defined above.
-Since the restriction set is a strict subset of the treatment assignment space, the null is not sharp.
+Note the non-imputability of the potential outcome function under the null considered above for different treatment assignments.
+Precisely, given $(Z_{obs},G_{obs})$, we can compute potential outcomes for treatment assignments in the restriction set defined above.
+The null is not sharp since the restriction set is a strict subset of the treatment assignment space.
 2.3 Null hypothesis of no spillover and conditional randomization test
 In this paper, we are concerned with testing first-order spillover effects. We take $k=1$ in Example 1 and state the null hypothesis of no spillover effects below.
+{mdframed}
+[leftmargin=0.1in,rightmargin=0.1in]
 Hypothesis 4.
 (Null hypothesis of no spillover effects)
-$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
+$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
 […]
@@
 Using the above assumption of linearly separable joint treatment effect, we get
-$\forall i\in[N]\quad Y^{2}(Z_{\mathcal{N}_{G}(i)})=Y^{2}(Z^{\prime}_{\mathcal{N}_{G^{\prime}}(i)})\qquad$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
+$\forall i\in[N]\quad Y^{2}(Z_{\mathcal{N}_{G}(i)})=Y^{2}(Z^{\prime}_{\mathcal{N}_{G^{\prime}}(i)})$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
 […]
-A similar approach to defining non-sharp null of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al.
+This approach of defining non-sharp null of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al.
 (2019b); Puelz et al.
@@
 The exposure functions are functions of the treatment assignment vector, $Z$, such that restriction sets can be created on $Z$ by partitioning $Z$ on the range of the exposure function.
-In the above Example 1, we can define an exposure function as below.
+We can define an exposure function in the above Example 1:
 $f_{G}(Z)=\frac{Z\cdot D_{G}^{(k-1)}(i)}{|D_{G}^{(k-1)}(i)|}.$
@@
 Since we consider the graph underlying the population to be random, we define restrictions on the joint treatment assignment in Hypothesis 3.
-A spillover mechanism, if existing, will be captured in the potential outcome function, and hence, we do not separately define a joint exposure mapping function by generalizing the exposure mappings.
-Careful identification of a spillover mechanism will be required for the inversion of the hypothesis test to obtain the confidence interval.
+A spillover mechanism, if existing, will be captured in the potential outcome function.
 We describe the currently proposed method in the literature of a conditional-randomization test procedure to obtain valid p-values (Puelz et al.
 (2022)).
-Procedure: Consider $C$ to be a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|C)$ is imputable under the null.
+Procedure 0: Consider $C$ as a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|C)$ is imputable under the null.
   1. 1.
@@
 The above-stated procedure holds finite-sample validity.
-We note that the conditioning mechanism $C$ is drawn from $P(C|Z_{obs})$, which should be determined by the analyst.
+The conditioning mechanism $C$ is drawn from $P(C|Z_{obs})$, which the analyst should determine.
 One must ensure the imputability of the test statistic $T(Z|C)$ when constructing a mechanism for $C$.
@@
 This is determined by the test statistic’s responsiveness to the null hypothesis.
-3   Quasi-randomization test based on network distribution
+3 Quasi-randomization test based on network
 We present the quasi-randomization test for the sharp null of no spillover effects.
 Our main idea is to generate the null sampling distribution of a test statistic by randomizing over the network-generating process.
-Intuitively, under the null of no spillover effect, the underlying network structure does not contribute to the total treatment effect.
+Intuitively, the underlying network structure does not contribute to the total treatment effect under the null of no spillover effect.
 This can also be seen in Hypothesis 4 where the network variable $G$ is unrestricted in the, otherwise, non-sharp null.
-This leads to the imputability of a test statistic of choice by fixing the direct treatment assignment $Z$ and randomizing over the distribution of network $G$.
+This leads to the imputability of a test statistic of choice by fixing the direct treatment assignment $Z$ and randomizing over the network distribution $G$.
 We formally define the conditional test statistic, $T_{c}$, below and show its imputability under the null Hypothesis 4.
@@
 Note that under $\mathcal{H}_{0}$,
-$Y(Z,G)=Y(Z^{\prime},G^{\prime})\qquad Z=Z^{\prime},$
+$Y(Z,G)=Y(Z^{\prime},G^{\prime})\qquad Z=Z^{\prime}.$
 Thus,
 $T_{c}(G|Z_{obs})$ $=T(Z_{obs},G,Y(Z_{obs},G)),$
-$=T(Z_{obs},G,Y(Z_{obs},G^{\prime}))\qquad P(G^{\prime})>0\text{ under }\mathcal{H}_{0}.$
+$=T(Z_{obs},G,Y(Z_{obs},G^{\prime}))\qquad P(G^{\prime}|Z_{obs})>0\text{ under }\mathcal{H}_{0}.$
 Since $Z_{obs}$ is arbitrary, we conclude imputability of the $T_{c}(G|Z_{obs})$ under the null $\mathcal{H}_{0}$. ∎
@@
 The procedure assumes the knowledge of the network-generating process $P(G)$.
-An elaborate consideration of the distribution of the network will be presented in Section 4.
+An elaborate consideration of the network distribution will be presented in Section 4.
 Procedure 1: Let $(Z_{obs},G_{obs},Y_{obs})$ be the observed joint treatment assignment and its corresponding outcome. Consider a test statistic $T(Z,G,Y_{obs})$.
@@
 Consider the null hypothesis, $\mathcal{H}_{0}$, in Hypothesis 4.
-Let $Z_{obs}\sim P(Z)$ be the treatment assignment of a randomized experiment.
-We assume that $G_{obs}\sim P(G)$ where $P(G)$ is known.
-Let $T(G|Z)$ be an imputable test statistic under the null $\mathcal{H}_{0}$.
+Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a randomized experiment.
+We assume that $P(Z,G)$ is known.
+Consider a test statistic $T(Z,G,Y(Z,G))$.
 Then, Procedure 1 described above is conditionally valid at level $\alpha\in(0,1)$.
 That is,
-$\mathbb{E}\bigl{[}\mathcal{I}(\text{p-val}(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1),$
-Here, the expectation is with respect to the distribution of $P(G_{obs}|Z_{obs})$.
+$\mathbb{E}\bigl{[}\mathcal{I}(\text{p-val}(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1).$
+Here, the expectation is for the distribution of $P(G_{obs}|Z_{obs})$.
 Proof.
 The proof is given in Appendix B.1. ∎
-Note that conditional validity implies unconditional validity of the procedure too.
-A crucial component of the proof is the identification of the conditional distribution $P(G|Z)$ for the validity of Procedure 1.
+Note that conditional validity implies unconditional validity of the procedure, too.
+A crucial component of the proof is identifying the conditional distribution $P(G|Z)$ for the validity of Procedure 1.
 This is contingent upon the experimental design, and different experimental designs will lead to different conditional randomization distributions.
@@
 Thus, different experimental designs lead to different $P(Z|C)$.
-A common problem faced in the literature is in the implementation of complex experimental designs like cluster randomized experiments.
-In a cluster-randomized experiment, the population is classified into a set of clusters where treatments are assigned at the cluster level.
-If a cluster is treated, then all units in that cluster receive the treatment or vice-versa.
+A common problem faced in the literature is implementing the methods in cluster randomized experiments.
+In a cluster-randomized experiment, the population is classified into clusters where treatments are assigned at the cluster level.
+If a cluster is treated, all units in that cluster receive the treatment or vice-versa.
 Fixing the treatment assignment of units in the conditioning mechanism might render the null randomization distribution $P(Z|C)$ degenerate, given the restricted span of cluster treatment assignments space.
@@
 This reduction in the support of conditional distribution leads to either degeneracy or enormous loss of power.
-Our methodology captures the null distribution by randomizing over $G|Z$, where the null distribution admits large support.
+Our methodology captures the null distribution by randomizing over $G|Z$, where the null distribution admits ample support.
 We discuss this in Sections 3.1.1 and 3.1.2, where we identify the null randomization distribution under different experimental designs.
@@
 Let $N_{c}=N-N_{t}$ be the size of the units in the control group.
-$P(Z=z)=\begin{cases}\frac{1}{\binom{N}{N_{t}}}&\sum_{i=1}^{N}z_{i}=N_{t},\\ 0&\text{otherwise}.\end{cases}$ (1)
-We remark that $G\!\perp\!\!\!\perp Z$ in Equation 1. Hence, for completely randomized experiments,
+$P(Z=z\;|\;G)=\begin{cases}\frac{1}{\binom{N}{N_{t}}}&\sum_{i=1}^{N}z_{i}=N_{t},\\ 0&\text{otherwise}.\end{cases}$ (1)
+We remark that $G\!\perp\!\!\!\perp Z$ by design. Hence, for completely randomized experiments,
 $P(G|Z)=P(G).$
@@
 We propose a new approach that considers sufficient statistics of the random graph model to build a permutation test.
-4   Quasi-permutation test via distributional properties
+4 Permutation test via distributional properties
 Identification of null randomization distribution $P(G|Z)$ involves knowledge of the distribution of the network-generating process, which is not identified by the design of the experiment.
 To this end, we explore random graph models to model $P(G)$.
-As discussed in the previous section, this identification of null randomization distribution becomes challenging when the complexity of the experimental design increases.
+As discussed in the previous section, identifying null randomization distribution becomes challenging in cluster-randomized experiments.
 This challenge is overcome in our proposed methodology, which is a conditional permutation test.
-The method does not involve estimating the null distribution.
 We discuss this in Section 4.1.1.
 4.1 Random graph models
-In this section, we describe various random graph models based on which $P(G)$ can be modeled.
-The seminal paper by Erdős and Rényi Erdős et al.
-(1960) introduced the concept of random graph for the first time.
-The Erdős-Rényi random graph model has since been studied widely, owing to exact computations of desirable graph properties.
-We now present the model formally in Section 4.1.1.
+This section describes various random graph models based on which $P(G)$ can be modeled.
+The Erdős-Rényi random graph model has been studied widely, owing to exact computations of desirable graph properties.
+We present the model formally in Appendix C.1.
 However, the Erdős-Rényi random graph model does not represent most real-world networks, such as social networks, supply-chain networks, and scientific collaboration networks, among many others.
-We look at a generalization of Erdős-Rényi random graph that is better adept at modeling such real-world networks, proposed by Newman et al.
+We look at a generalization of Erdős-Rényi random graph better at modeling such real-world networks, proposed by Newman et al.
 (2002), in Section 4.1.1.
-4.1.1   Erdős-Rényi random graph model
-We formally describe the Erdős-Rényi random graph model (Erdős et al.
-(1960)).
-The below-described model is the probabilist variant of the (Erdős-Rényi) random graph model.
-We consider the number of nodes, $N$, and edges, $|E|$, to be fixed, and any two nodes will be linked independently with an edge with a fixed probability $p$.
-That is,
-$P(G=G_{obs}|p)=\binom{\binom{N}{2}}{|E|}\cdot p^{|E|}\cdot(1-p)^{\binom{N}{2}-|E|}.$ (3)
-In order to model $P(G)$ using Equation (3), we have to estimate the parameter $p$.
-We obtain the MLE of the parameter $p$ and estimate it as $\frac{|E|}{\binom{N}{2}}$.
-We conduct an extensive numerical simulation study and verify the finite sample validity of the proposed randomization test.
-We also obtain the power of the test under different scenarios, which are presented in Appendix C.
-An Erdős-Rényi random graph model is the simplest network model one can consider, and it may not be representative of the kind of networks we observe in the real world.
-For example, in social networks, one observes a high degree of clustering in the graph (Watts and Strogatz (1998)).
-One drawback of the model is the independence and identical probability of formation of two edges in the graph.
-One way to characterize this is to evaluate the degree distribution of the random graph.
-It has been shown that the degree distribution of a large Erdős-Rényi random graph is a Poisson distribution.
-Most real-life networks violate this property and have highly skewed degree distribution.
-In the next section, we generalize this and characterize random graphs with arbitrary degree distribution.
+For this article, we restrict the models to be covariate-free.
 4.1.1 Random graphs with arbitrary degree distribution
@@
 (2002)).
-In order to deploy Procedure 1, we have to identify $P(G)$.
+To deploy Procedure 1, we must identify $P(G)$.
 This involves estimating $(p_{1},p_{2},...,p_{N})$ with a single realization of the graph, and we cannot directly implement Procedure 1 without estimating $P(G)$.
-In the case of random graphs with arbitrary degree distribution, it is proven that we can possibly estimate the parameters of the random graph with just one realization but with some uncertainty (Chatterjee et al.
+In the case of random graphs with arbitrary degree distribution, it is proven that we can estimate the parameters of the random graph with just one realization but with some uncertainty (Chatterjee et al.
 (2011)).
+However, there might be other relational dependencies characterizing the formation of the network that are not captured by random graphs with arbitrary degree distribution.
+For example, the phenomena of transitivity are not captured in this model, where two nodes have higher chances of being connected if they have a common connecting node.
+4.1.2 Exponential random graph model
+The state-of-the-art method for characterizing random graph formation of social networks is through the exponential random graph model or ERGM (Snijders et al.
+(2006)).
+It is a generalization of the random graphs with arbitrary degree distribution, where the characterization of the distribution of the random graph is through a set of given sufficient statistics.
+This could also contain information about covariates to capture homophily in the network formation process.
+In this article, we consider the random graph formation free of covariates.
+Definition 4.
+Consider a random graph $G\in\{0,1\}^{N\times N}$ such that $P(G=g|\eta)=\frac{1}{\kappa}\cdot{e^{\eta^{T}\cdot s(g)}}$.
+Here, $s(g)$ represents a vector of sufficient statistics, and $\kappa$ is the normalizing constant.
+We call $G$ an exponential random graph.
+Common sufficient statistics used in the literature for the characterization of social networks include the number of edges, the number of triangles, and the number of stars of different sizes, say stars of size two or stars of size three (Robins et al.
+(2007)).
+It is important to carefully choose the model as an ill-posed ERGM can lead to near degeneracy (Chatterjee and Diaconis (2013); Schweinberger et al.
+(2020)).
+As with random graphs with arbitrary degree distribution, estimating ERGM is fraught with challenges given just a single realization.
 Owing to estimation issues of the null distribution, we propose a new method that does not involve such estimations and reduces to a simple permutation test.
-In our proposed methodology, we use sufficient statistics of the random graph distribution, degree sequence in this case, and develop a method conditioning on the sufficient statistic.
-We use the symmetric nature of the $P(G)$ via the equivalence relation and build a conditional permutation test, as described below.
+In our proposed methodology, we first consider the sufficient statistics of the random graph distribution as degree sequence and develop a method conditioning on the sufficient statistic.
+We then generalize the method to a set of arbitrary sufficient statistics.
+We use the symmetric nature of the $P(G)$ via an equivalence relation and build a conditional permutation test, as described below.
 4.2 Permutation test based on equiprobability events
-Here, we present our main result, which proposes the procedure to obtain a conditional randomization test without estimating the distribution.
-Our main idea is to identify a sufficient statistic for the random graph model, which is degree sequence in the case of random graphs with arbitrary degree distribution.
-We then sample the null distribution over all the graphs with the same degree sequence as the observed graph.
-Using the equiprobability of the event set of all the graphs with the same degree sequence, we obtain a permutation test.
+4.2.1 Test for a random graph with arbitrary degree distribution
+Here, we present our main result, which proposes obtaining a conditional randomization test without estimating the distribution.
+Our main idea is to identify a sufficient statistic for the random graph model, which is degree sequence in the case of random graphs with arbitrary degree distribution, and condition on this sufficient statistic.
+We sample the null distribution over all the graphs with the same degree sequence as the observed graph.
+We obtain a permutation test using the equiprobability of the event set of all the graphs with the same degree sequence.
 We now formally present the definitions below and proceed to state and prove the finite sample validity of the proposed method.
@@
 Consider $D=(d_{1},d_{2},...,d_{N})$ such that $d_{1}\leq d_{2}\leq...\leq d_{N}$.
-We call $D$ to be the degree sequence of $G=(V,E)$ if there exists a one-to-one mapping from $\{deg(i):i\in V\}$ to $D$.
+We call $D$ to be the degree sequence of $G=(V=\{v_{1},v_{2},...,v_{N}\},E)$ if there exists a one-to-one mapping $\delta_{G}:V\rightarrow D$ such that $\delta(v_{i})=deg(v_{j})$ for some $v_{j}\in V$ for all $v_{i}\in V$.
 We denote $D$ as $deg(G)$.
@@
 Note that graphs with the same degree sequence are not necessarily isomorphic graphs.
-This implies that the vertices of the graph cannot be permuted to obtain the other graph with edges preserved (see Figure 1).
+This implies that the graph’s vertices cannot be permuted to obtain the other graph with edges preserved (see Figure 1).
 Figure 1: Non-isomorphic graphs with same degree sequence
-The converse holds true.
+The converse holds.
 Consider a set of distinct graphs with the number of units fixed to $N$ and a degree sequence of the graphs fixed to a valid degree sequence $D$.
@@
 We define a relation $d:\mathcal{G}\rightarrow\mathcal{G}$ such that $G\equiv_{d}G^{\prime}$ if $deg(G)=deg(G^{\prime})$.
-It is a quick check that $d$ defines an equivalence relation. Hence, we can partition the support of $P(G)$, where each part in the partition represents an equivalence class of $D$.
+It is a quick check that $d$ defines an equivalence relation. Hence, we can partition the support of $P(G)$, where each part represents an equivalence class of $D$.
 Proposition 2.
@@
 Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$.
-Let $Z_{obs}\sim P(Z)$ be the treatment assignment of a cluster randomized experiment.
-We assume that $G_{obs}\sim P(G)$ where $G$ is a random graph with arbitrary degree distribution.
-Let $T(G|Z)$ be an imputable test statistic under $\mathcal{H}_{0}$.
+Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a completely randomized experiment.
+We assume that $G_{obs}\sim P_{G}$ where $G$ is a random graph with arbitrary degree distribution.
+Consider a test statistic $T(Z,G,Y(Z,G))$.
 Then, Procedure 2 described above is conditionally valid at level $\alpha\in(0,1)$.
 That is,
-$\mathbb{E}\bigl{[}\mathcal{I}(p(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1),$
-Here, the expectation is taken with respect to the distribution of $P(G_{obs}|Z_{obs})$.
+$\mathbb{E}\bigl{[}\mathcal{I}(p(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1).$
+Here, the expectation is taken for the distribution of $P(G_{obs}|Z_{obs})$.
 Proof.
 Refer to Appendix B.2. ∎
-As a corollary, one can take all the clusters to be of size one, implying every unit is a cluster.
-This retrieves the setup for a completely randomized experiment, and thus, Theorem 2 also holds true for a completely randomized experiment.
-We emphasize that the generation of graphs with a given degree sequence admits a polynomial time algorithm.
+As a corollary, we can substitute any sub-class of the degree sequence equivalence class in Procedure 2, and the finite-sample validity of the Procedure is retained.
+For example, we can take the equivalence class to be all the graphs isomorphic to the observed graph in the second step of Procedure 2.
+We present results for both these Procedures in Section 5.
+We emphasize that generating graphs with a given degree sequence admits a polynomial time algorithm.
 Given fixed parameters like the maximum degree of the graph, one can use state-of-the-art algorithms to efficiently sample graphs with a given degree sequence uniformly from the space of all the graphs with the same degree sequence (Arman et al.
@@
 It has been shown that the size of this equivalence class can be exponentially large (Barvinok and Hartigan (2013)).
-It should be noted that one should use a uniform sampler in order to avoid introducing any bias in the null distribution attributed to sampling.
-We can also overcome this by taking a large number of samples to approximate the null distribution effectively.
+It should be noted that one should use a uniform sampler to avoid introducing any bias in the null distribution attributed to sampling.
+We can also overcome this by effectively taking many samples to approximate the null distribution.
+4.2.2 Test for exponential random graph model
+We generalize the above method to a class of exponential random graph models where the sufficient statistics of the model are permutation invariant.
+We achieve this by conditioning on the graph isomorphism class and generating the null cases from this equivalence class.
+We formally present the results below.
+A graph is considered isomorphic to another graph if a one-to-one correspondence between the vertices of the two graphs preserves the edges between the graphs.
+Below, we present an equivalence condition for the same.
+Definition 7.
+Two graphs $G$ and $G^{\prime}$ are isomorphic if and only if a permutation matrix $P$ exists such that $G=P^{T}G^{\prime}P$.
+We denote $G$ and $G^{\prime}$ as isomorphic graphs using $G\cong G^{\prime}$.
+We note that graph isomorphism defines an equivalence relation.
+The permutation matrix comprises row binary vectors indexed by the one-to-one correspondence between the vertices of the graphs.
+We now define a class of exponential random graph models based on the permutation invariance property of its sufficient statistic.
+Definition 8.
+An exponential random graph model $G\in\{0,1\}^{N\times N}$ is called permutation-invariant if $s(\Pi(g))=s(g)$.
+Here, $\Pi$ represents the permutation function of the graph’s vertices, and $s(.)$ represents the vector of sufficient statistics of the exponential random graph model.
+Many sufficient statistics used in real-life social network modeling follow this property.
+As discussed, some graph statistics include the number of triangles, degree sequence, and number of stars of size two, and they are permutation invariant.
+We model the data-generating process as an exponential random graph with permutation invariance property and present a permutation testing procedure.
+Procedure 3: Let $(Z_{obs},G_{obs},Y_{obs})$ be the joint treatment assignment vector and its corresponding outcome. Consider a test statistic $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$.
+  1. 1.
+    Impute the observed value of test statistic ${T_{c}}_{obs}=T_{c}(G_{obs}|Z_{obs})$.
+  2. 2.
+    Consider the sets $Z^{t}:=\{i\in\mathbb{P}:\>Z_{i}=t\}\;\forall\;t\in\{0,1\}$.
+  3. 3.
+    Define the equivalence class
+    $[G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\;|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
+  4. 4.
+    Compute p-value as
+    $p(Z_{obs},G_{obs},Y_{obs}):=\frac{1}{|[G_{obs}|Z_{obs}]|}\cdot\sum_{G\in[G_{obs}|Z_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})).$
+Theorem 3.
+Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$.
+Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a cluster randomized experiment.
+We assume that $G_{obs}\sim P_{G}$ where $G$ is a permutation-invariant exponential random graph model.
+Consider a test statistic $T(Z,G,Y(Z,G))$.
+Then, Procedure 3 described above is conditionally valid at level $\alpha\in(0,1)$.
+That is,
+$\mathbb{E}\bigl{[}\mathcal{I}(p(Z_{obs},G_{obs},Y_{obs})\leq\alpha)|\mathcal{H}_{0}\bigl{]}\leq\alpha\qquad\forall\text{ }\alpha\in(0,1).$
+Here, the expectation is taken for the distribution of $P(G_{obs}|Z_{obs})$.
+Proof.
+Refer to Appendix B.2. ∎
 4.3 Test statistic
@@
 This looks at the difference in average treatment effect amongst control units with at least one treated neighbor and control units with only control units as neighbors.
-Essentially, we treat focal units as control units and auxiliary units as treated units.
-$T_{I_{c}}(G|Z_{obs})$ $=$ (4)
-$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)},$
-Note that the above test statistic considers the spillover on the control units from the treated.
-Under the null of no spillover effect for all the units regardless of direct treatment status, one can similarly define the above test statistic for the treated units.
-$T_{I_{t}}(G|Z_{obs})$ $=$ (5)
-$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)},$
+Let $T_{I_{c}}(G|Z_{obs})=$
+$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=0)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=0)}.$ (3)
+The above test statistic considers the spillover on the control units from the treated.
+Under the null of no spillover effect for all the units regardless of direct treatment status, one can also define the above test statistic for the treated units.
+Similarly, let $T_{I_{t}}(G|Z_{obs})$ equals
+$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}>0,Z_{i_{obs}}=1)}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)}{\sum_{i=1}^{N}\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}=0,Z_{i_{obs}}=1)}.$ (4)
 We now define the test statistic, denoted by $T_{I}(G|Z_{obs})$, for testing the spillover effect for all the units. It will be a weighted average of the two test statistics defined above.
-$T_{I}(G|Z_{obs})=\frac{N_{c}}{N}\cdot T_{I_{c}}+\frac{N_{t}}{N}\cdot T_{I_{t}}$ (6)
+$T_{I}(G|Z_{obs})=\frac{N_{c}}{N}\cdot T_{I_{c}}+\frac{N_{t}}{N}\cdot T_{I_{t}}.$ (5)
 It should be noted the above test statistics can only be defined if the sub-populations in consideration are non-empty.
-For example, if there exists no control unit with all neighbors in control or the number of such units is few, we deal with either a definition or power issue.
+For example, if no control unit exists with all neighbors in control or the number of such units is few, we deal with either a definition or power issue.
 Thus, we propose a generalization of the above-stated test statistic for robustness.
 We look at the difference in average treatment effect among control units with greater than 0.75 quantile of the proportion of treated units and control units with less than 0.25 quantile of the proportion of treated units.
-$T_{quant_{c}}(G|Z_{obs})$ $=$ (7)
-$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\leq p_{0.25})}{|P_{0.25}|},$
+Let $T_{quant_{c}}(G|Z_{obs})=$
+$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}$ (6)
+$-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\leq p_{0.25})}{|P_{0.25}|}.$
 Here, $p_{0.25}$ and $p_{0.75}$ represent the first and the third quartile of the vector
@@
 $P_{0.25}$ and $P_{0.75}$ represent the set of elements in $P$ in the first and third quartile, respectively.
-In a similar fashion to the test statistic $T_{I}(G|Z_{obs})$, we define the above for treated units and take a weighted average for the two test statistic to obtain $T_{quant}(G|Z_{obs})$.
+In a similar fashion to the test statistic $T_{I}(G|Z_{obs})$, we define the above for treated units and take a weighted average for the two test statistics to obtain $T_{quant}(G|Z_{obs})$.
 We remark that one can improve the test statistic with the knowledge of appropriate covariates or nodal attributes.
@@
 For example, parametric test statistics can be constructed of the linear-in-means model of interference proposed by Manski (1993).
-See Athey et al.
-(2018) for a more elaborate discussion on this.
+We also consider more test statistics, given that the choice is not unique and different test statistics would differ in their power.
+We take the test statistic proposed in Bond et al.
+(2012), which takes the difference between the average over all the edges where the neighbor is a treated or control unit.
 5 Simulation study
 We present Monte Carlo simulations to validate our proposed testing procedure.
-We consider the test statistic based on the quantile of the proportion of treated neighbors in the simulations, as discussed in Section 4.3.
-We will describe the data-generating process and the general setup in the next section.
+We consider two test statistics, one that is based on the quantile of the proportion of treated neighbors in the simulations and another that is weighted by treated neighbors (first proposed in Bond et al.
+(2012)), as discussed in Section 4.3.
+The following subsection will describe the data-generating process and the general setup.
 We conduct the simulation study for two experimental designs, namely, completely randomized experiments and cluster randomized experiments.
 As a robustness check, we repeat the study for two network-generating processes.
-Table 1: Rejection rates for the null hypothesis of no spillover effects for completely randomized experiment
-Network Distribution Direct Effect Spillover Effect Degree Permutation Test Conditional Focal Test
-$T_{quant}$ $T_{elc}$
-Small world network 0 0 0.049 0.050
-4 0 0.046 0.046
-0 0.4 0.106 0.112222We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
-4 0.4 0.101 0.060
-Stochastic block model 0 0 0.062 0.054
-4 0 0.055 0.052
-0 0.4 0.130 0.065
-4 0.4 0.140 0.048
 5.1 Data generating mechanism
@@
 We state that $N_{t}$ is fixed and pre-determined by the experimenter.
-In the performed simulations, we have $599$ units and $300$ treated units in the model in one case.
-We study the setup for two generating processes, namely, small-world networks and the stochastic block model.
-Both models exhibit a high degree of clustering, which is characterized by the clustering coefficient.
-The clustering coefficient is the number of connections in a randomly chosen node and its neighborhood relative to the total number of connections possible in the neighborhood.
-Consider a regular graph, which is a graph where all the nodes have the same number of connections.
+In the performed simulations, we have $599$ total units and $300$ treated units in the model in one case.
+We study the setup for two generating processes: small-world networks and the stochastic block model.
+Both models exhibit a high degree of clustering, characterized by the clustering coefficient.
+The clustering coefficient is the number of connections in a randomly chosen node and its neighborhood relative to the total number of connections possible.
+Consider a regular graph where all the nodes have the same cardinality of connections.
 Let us denote the number of connections a unit has in a regular graph by $K$.
-Such graphs have a high clustering coefficient but also distance between two nodes in such graphs is also high.
-It is observed in real-life networks that the average distance between two nodes is low (Amaral et al.
+Such graphs have a high clustering coefficient, but the distance between two nodes in such graphs is also high.
+Table 1: Rejection rates for the null hypothesis of no spillover effects for completely randomized experiment.
+Network Distribution Direct Effect Spillover Effect Degree Permutation Test Isomorphism Permutation Test Conditional Focal Test
+$T_{bond}$ $T_{quant}$ $T_{bond}$ $T_{quant}$ $T_{elc}$
+Small world network 0 0 0.045 0.049 0.049 0.047 0.050
+4 0 0.054 0.046 0.056 0.045 0.046
+0 0.4 0.187 0.148 0.191 0.162 0.112111We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
+4 0.4 0.074 0.151 0.064 0.147 0.060222We obtain the same rejection rate in our replication as reported in Athey et al. (2015).
+Stochastic block model 0 0 0.045 0.062 0.050 0.050 0.054
+4 0 0.050 0.055 0.047 0.052 0.052
+0 0.4 0.191 0.217 0.189 0.216 0.065
+4 0.4 0.064 0.220 0.063 0.217 0.048
+The average distance between two nodes in real-life networks is low (Amaral et al.
 (2000)).
@@
 This parametrization characterizes a small world network.
-The higher the probability of rewiring, the lower the average path length gets.
+The higher the rewiring probability, the lower the average path length.
 Another class of random graph models that achieves the same property of a high clustering coefficient but a low average path length is that of the stochastic block model.
@@
 Note that because we want a high degree of clustering in the graph, the diagonal elements of the probability matrix are typically higher than the off-diagonal elements.
-The probability matrix is also referred to as the preference matrix in the literature.
+The probability matrix is also called the preference matrix in the literature.
 We now define the setup for the simulation study:
@@
   2. 2.
-    Direct treatment effect: We take $\tau_{direct}$ equal to 4 and 0 in our simulations.
+    Direct treatment effect: We take $\tau_{direct}$ equal to 0 and 4 in our simulations.
   3. 3.
@@
     Test statistic: We consider the $T_{quant}$ test statistic described in Section 4.3.
-It takes the difference in the treatment effect of control units in the first quartile of the proportion of treated neighbors to control units in the last quartile of the proportion of treated neighbors.
-We define the same for treated units and take the weighted average of both.
-The null randomization distribution was generated by sampling graphs with the same degree sequence as the observed graph.
-The null distribution was approximated by generating 1000 graphs with the same degree sequence as the observed graph.
-The p-value is calculated as the sample proportion of times the imputed test statistic for different graph realizations, with the same degree sequence as the original graph, is more than the observed test statistic.
-If the p-value obtained is less than the chosen significance level of 0.05, we reject the null hypothesis.
+It takes the difference in the treatment effect of control units in the first quartile of the proportion of treated neighbors to control units in the last quartile.
+We define the same for treated units and take the weighted average.
+We also consider the $T_{bond}$ test statistic, proposed in Bond et al.
+(2012), defined as follows:
+    $T_{bond}(G|Z_{obs})=\frac{\sum_{i,j\in[N]}G_{ij}\cdot Z_{j}\cdot Y_{i}}{\sum_{i,j\in[N]}G_{ij}\cdot Z_{j}}-\frac{\sum_{i,j\in[N]}G_{ij}\cdot(1-Z_{j})\cdot Y_{i}}{\sum_{i,j\in[N]}G_{ij}\cdot(1-Z_{j})}.$
+The null distribution was generated by sampling 1000 graphs with the same degree sequence as the observed graph and isomorphic to the observed graph in a second option.
+The p-value is calculated as the sample proportion of times the imputed test statistic for the sampled graph realizations is more than the observed test statistic.
+We reject the null hypothesis if the p-value obtained is less than the chosen significance level of 0.05.
 We replicated the above-described setup 4000 times and calculated the proportion of times we rejected the null.
@@
 (2018)).
+We label this test as ‘Conditional Focal Test’.
 We then perform the simulations for cluster randomized experimental setup.
@@
 We take epsilon to be equal to 3.
-This implies all the clusters contain units that are less than 3 hops away from each other, whereas between cluster distance is at least 3 hops for any two units.
+This implies all the clusters contain units less than three hops away from each other, whereas the between-cluster distance is at least three hops for any two units.
 We take the probability of treating a cluster to be 0.5.
 We implemented a Bernoulli trial at the cluster level.
-The null distribution is again approximated by taking 1000 samples of graphs with the same degree sequence as the original graph.
-We replicated this setup 2000 times and calculated the proportion of times the proposed test rejected the null.
-We reduce the replication size to 2000 for cluster randomized setup observing appropriate convergence of the p-value estimate.
+Under Procedure 3, the null distribution is approximated by taking 1000 samples of graphs isomorphic to the original graph, such that the corresponding subgraphs on treated units and controlled units remain isomorphic.
+This is obtained by considering permutations where treated units can be mapped only to treated units (and thus, control units can only be mapped to control units).
+We replicated this setup 4000 times and calculated the proportion of times the proposed test rejected the null.
+We label Procedure 3 as ‘Block Isomorphism Permutation Test’ in Table 2.
 This was performed for two network-generating processes as before.
 Table 2: Rejection rates for the null hypothesis of no spillover effects for cluster randomized experiment
-Network Distribution Direct Effect Spillover Effect Degree Permutation Test
-$T_{quant}$
-Small world network 0 0 0.055
-4 0 0.055
-0 0.4 0.213
-4 0.4 0.217
-Stochastic block model 0 0 0.053
-4 0 0.055
-0 0.4 0.168
-4 0.4 0.164
+Network Distribution Direct Effect Spillover Effect Block Isomorphism Permutation Test
+$T_{bond}$ $T_{quant}$
+Small world network 0 0 0.052 0.053
+4 0 0.056 0.051
+0 0.4 0.429 0.352
+4 0.4 0.429 0.348
+Stochastic block model 0 0 0.051 0.057
+4 0 0.048 0.046
+0 0.4 0.225 0.237
+4 0.4 0.227 0.242
 5.2 Comparison with related work
-We see that our proposed method has type 1 error controlled appropriately.
+We see that our proposed methods have type 1 error controlled appropriately.
 This corresponds to rows with $\tau_{spill}$ equal to zero in Tables 1 and 2.
-When $\tau_{spill}$ is equal to 0.4, we see that our method performs with competitive power compared to existing methods in the literature (Athey et al.
+When $\tau_{spill}$ equals 0.4, our method performs with competitive power compared to existing methods in the literature (Athey et al.
 (2018)).
-We highlight the case of $\tau_{direct}$ equal to 4 and $\tau_{spill}$ equal to 0.4, where our method displays substantial improvement.
+We highlight the $\tau_{direct}$ equal to 4 and $\tau_{spill}$ equal to 0.4, where our method significantly improves.
 The significance of the setup lies in its resemblance to many naturally occurring settings, where the direct effect is meaningfully more than the spillover effect (for example, see Cai et al.
 (2015)).
-We see that when the clustering of the network-generating process increases, as is the case with the stochastic block model, our proposed method displays a boost in power.
-This is in contrast with the method proposed by Athey et al.
+We see in Table 1 that our proposed method boosts power when the clustering of the network-generating process increases, as with the stochastic block model.
+This contrasts with the method proposed by Athey et al.
 (2018), where the method does not retain power.
+In Appendix D, we add more insights on the performance of the $T_{bond}$ test statistic
+as observed in Table 1.
 An interesting result is in the case of cluster-randomized experiments where a higher degree of clustering shows a drop in power.
@@
 This can be attributed to a reduction of sample space of imputed values of the test statistic.
-Hence, a higher degree of clustering increases the cluster sizes, which restricts the direct treatment assignment.
-The observed graph in the case of the small-world network had 16 clusters, as compared to 11 clusters in the case of the stochastic block model.
+Hence, a higher degree of clustering increases the cluster sizes, restricting the direct treatment assignment.
+For instance, an observed graph in the case of the small-world network had 16 clusters, compared to 11 clusters in the case of the stochastic block model.
 We also perform the simulation study for cluster randomized experiments for the methods proposed in Athey et al.
@@
 (2022).
-As discussed in Section 3, we encounter the degenerate null distribution and hence, do not present the corresponding results.
+As discussed in Section 3, we encounter the degenerate null distribution and do not present the corresponding results.
 This is more prominent in the method proposed by Athey et al.
-(2018), in comparison with Puelz et al.
+(2018), compared with Puelz et al.
 (2022).
 Since Puelz et al.
-(2022) takes into account the design of the experiment when selecting focal units, we see more robustness of the biclique test (Puelz et al.
+(2022) considers the experiment’s design when selecting focal units, we see more robustness of the biclique test (Puelz et al.
 (2022)) where degenerate null distributions are reduced in frequency.
 Overall, our method stands more robust to degeneracy and performs well in terms of power in our simulations.
-6   Application: Weather insurance adoption in rural China
+We also see that the $T_{bond}$ test statistic performs slightly better than $T_{quant}$ test statistic in some cluster-randomized setups.
+This can be attributed to $T_{bond}$ being weighted by treated numbers in the graph instead of the corresponding quantile-based weights in $T_{quant}$ and the heterogeneity of the clustered assignment setup.
+We emphasize that the $T_{bond}$ in conjunction with our methodology continues to be valid regardless of the size of the direct effect.
+The method proposed by Bond et al.
+(2012) does not work for settings with non-zero direct effects (Refer Athey et al.
+(2018) for an elaborate discussion).
+6 Weather insurance adoption in rural China
 We re-analyze a field experiment conducted in rural China on rice crop farmers studying neighborhood interference effects (Cai et al.
@@
 The research motive behind the experiment was to see if social networks among the farmers have any influence on weather insurance adoption.
-The experiment involved conducting intensive training sessions to promote crop insurance products to farmers against extreme weather.
+The experiment involved intensive training sessions to promote crop insurance products to farmers against extreme weather.
 A primary goal of the experiment was to identify if there is information dissemination in the farmers’ community and, if so, what mechanisms drive these information exchanges.
@@
 A deeper understanding of insurance product adoption via social network mechanisms helps devise a more cost-effective government policy campaign to enhance product adoption among the target community.
-Once interaction mechanisms can be established, better “seeding” or targeting strategies can be rolled out for future campaigns (Ok et al.
-(2014)).
-Such strategies identify influential figures in the community to maximize the interference effect and, hence, the program’s overall impact.
-Another important reason to study interference in such a setup is to avoid biased estimates and accurately evaluate policy performance measures.
+An important reason to study interference in such a setup is to avoid biased estimates and accurately evaluate policy performance measures.
 We discuss the experimental design setup by the researchers.
-The hypothesis of interest is the underlying assumption that access to better information about weather insurance products would increase the adoption of the product in the rice farmers’ community.
+Researchers hypothesize that access to better information about weather insurance products would increase the adoption of the product in the rice farmers’ community.
 To capture access to information differentially as treatment and control, two types of information sessions were created: simple sessions of 20 minutes, which provided an introduction to the insurance contract, and intensive sessions of 45 minutes that covered the insurance contract in much more detail, including explaining to them what the benefits are of taking up insurance.
@@
 After the round 2 sessions, the first group was given no additional information, the second group was told about the overall take-up rate in the previous round, and the third group was told in detail about who purchased the product and who did not.
-This additional information randomization of the second group was done to derive the main mechanism via which information diffusion took place.
+This additional information randomization of the second group was done to derive the primary information diffusion mechanism.
 For illustration purposes, we do not analyze this part of the experiment.
-There were a total of 4902 households across 47 different villages in the experiment.
-The social network among a village is known and is assumed to be independent between two distinct villages.
-Each household is randomized among the 4 different groups: (i) Receiving simple session in Round 1,(ii) Receiving intensive session in Round 1, (iii) Receiving simple session in Round 2, (iv) Receiving intensive session in Round 2.
-After the session ends, the households make the decision to purchase insurance.
+There were 4902 households across 47 different villages in the experiment.
+A village’s social network is known and assumed to be independent between two distinct villages.
+Each household is randomized among the 4 different groups: (i) Receiving a simple session in Round 1, (ii) Receiving an intensive session in Round 1, (iii) Receiving a simple session in Round 2, (iv) Receiving an intensive session in Round 2.
+After the session ends, the households decide to purchase insurance.
 This is captured as a binary outcome variable of the model, with 1 indicating a purchase decision made and 0 when no purchase is made.
-Under the assumption that no information diffusion can take place between participant households of Round 1 because of the immediate purchase decision the household had to make, we can assume there are no edges among the household participants in Round 1.
+Under the assumption that no information diffusion can occur between participant households of Round 1 because of the immediate purchase decision the household had to make, we can assume there are no edges among the household participants in Round 1.
 In our analysis, we take all the households receiving simple sessions under the control group and intensive sessions under the treated group.
 Since spillover can only occur from the treated group of Round 1 to the control group of Round 2, we only consider the network structure amongst these two groups in our study.
-We retain the network structure among the households between the two rounds and nullify the remaining network structure.
-This includes network structure among households in Round 1 only and network structure among households in Round 2 only.
+This includes network structure among households in Round 1 and Round 2 only.
 Our analysis was run on 433 households of randomly chosen villages in the experiment, of which 222 were in the treated group.
-We ran our proposed Procedure 2 and obtained a p-value of 0.0002, which indicates the presence of interference.
+We ran our proposed Procedure 2, the degree permutation test, and approximated the null distribution by taking 10,000 samples of graphs with the same degree sequence as the observed graph.
+We obtained a p-value of 0.0002, which indicates the presence of interference.
 7 Conclusion
-In this paper, we develop a new methodology to test for the non-sharp null hypothesis of no spillover effect in an experimental setup.
-Our proposed method assumes mild distributional conditions in the network-generating process to build a quasi-randomization test.
-Whilst promising, there are several open problems that remain.
-Incorporating covariates or nodal attribute information to better characterize the random graph null model may help enhance the power of the test and is an interesting direction for future work.
-We also assume fully and correctly specified information of the given network.
-An analysis of the random graph null models under partial information and network misspecification is another interesting avenue for future work.
+We develop a new methodology to test for the non-sharp null hypothesis of no spillover effect in an experimental setup.
+Our proposed method assumes natural distributional conditions in the network-generating process to build a quasi-randomization test.
+While promising, several open problems remain.
+Incorporating covariates or nodal attribute information to characterize the random graph null model better may help enhance the power of the test and is an interesting direction for future work.
+We also assume fully and correctly specified information about the given network.
+Analyzing the random graph null models under partial information and network misspecification is another exciting avenue for future work.
 Finally, our focus in this paper has been on randomized experiments exclusively.
-Since the sampling of null distribution is based on quasi-randomization of the network-generating process, a natural and important extension is to study random graph null testing procedures in observational studies.
-Acknowledgments
-We thank Sumit Kunnumkal, Parshuram Hotkar, and ISB seminar participants for their helpful comments and feedback.
+Since the sampling of null distribution is based on quasi-randomization of the network-generating process, a natural and essential extension is to study random graph null testing procedures in observational studies.
+Acknowledgments.
+We thank Avi Feller, Fabrizia Mealli, Panos Toulis, ACIC 2024 conference participants, and the Tom Ten Have Award committee for their encouragement and feedback.
+We also thank Sai Sriramya Gorripati, Parshuram Hotkar, Sumit Kunnumkal, and ISB seminar participants for their enthusiasm and comments.
 We gratefully acknowledge Hemanth Kumar and the ISB Institute of Data Science for providing cluster access.
-Pallavi Basu’s research is partially supported by the SERB MATRICS award MTR/2022/000073.
-There is no conflicting interest to declare.
 Appendix
@@
 Procedure:
+Sort $G_{obs}$ by degree of vertices in ascending order
 $S$ is equal to $G_{obs}$
 while $S$ is not empty do
-  • Step 1: Pick the first vertex in $S$, say $v_{1}$
-  • Step 2: Find all units within $\epsilon$ distance of $v_{1}$
+  • Step 1: Pick the last vertex in $S$, denoted by $v$, and store in $V$
+  • Step 2: Find all units within $\epsilon$ distance of $v$
   • Step 3: Store the units found in Step 2 as a cluster and remove these units from $S$
@@
 Store the resultant cluster outputs in $C$
-Output: C
+Permute the outputs to the original ordering and store them in $C,V$
+Output: C and V
 Algorithm 1 $\epsilon$-net cluster
@@
 We define p-value as
-$p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs}),$
-Given $Z_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution as induced by $P(G|Z_{obs})$. Thus,
+$p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs}).$
+Given $Z_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution induced by $P(G|Z_{obs})$. Thus,
 $p_{Z_{obs}}=1-F_{U}(G_{obs}|Z_{obs}).$
 Here, $F_{U}(.)$ represents the cumulative distribution of $U$. Then,
-$T_{c}(G|Z_{obs})=T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0},$
+$T_{c}(G|Z_{obs})\stackrel{{\scriptstyle d}}{{=}}T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0}.$
 Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z_{obs})$ and we get
-$p_{Z_{obs}}=1-F_{U}(U),$
+$p_{Z_{obs}}=1-F_{U}(U).$
 Hence, by using the probability integral transform theorem, we obtain
 $P(p_{Z_{obs}}\leq\alpha)\leq\alpha.$
-B.2   Proof of Theorem 2:
+B.2 Proof of Theorem 2 and Theorem 3:
 Proof.
@@
 Thus,
-$p_{Z_{obs},G_{obs}}=1-F_{U}(G_{obs}|Z=Z_{obs},G\in[G_{obs}]),$
+$p_{Z_{obs},G_{obs}}=1-F_{U}(G_{obs}|Z=Z_{obs},G\in[G_{obs}]).$
 Here, $F_{U}(.)$ represents the cumulative distribution of $U$. Then,
-$T_{c}(G|Z_{obs})=T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0},$
+$T_{c}(G|Z_{obs})\stackrel{{\scriptstyle d}}{{=}}T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0}.$
 Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z=Z_{obs},G\in[G_{obs}])$ and we get
-$p_{Z_{obs},G_{obs}}=1-F_{U}(U),$
+$p_{Z_{obs},G_{obs}}=1-F_{U}(U).$
 Hence, by using the probability integral transform theorem, we obtain
@@
 $P(p_{Z_{obs}}\leq\alpha)$ $\leq\alpha.$
-We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}])$ is a uniform distribution under a cluster randomized experiment. Consider $g,g^{\prime}\in[G_{obs}]$.
-$P(G=g|Z=Z_{obs},G\in[G_{obs}])=\frac{P(Z=Z_{obs}|G=g,G\in[G_{obs}])\cdot P(G=g|G\in[G_{obs}])}{P(Z=Z_{obs}|G\in[G_{obs}])},$ (9)
+Let $[G_{obs}]=\{G\in\{0,1\}^{N\times N}:\>deg(G)=deg(G_{obs})\}$.
+We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}])$ is a uniform distribution under a completely randomized experiment.
+Consider $g,g^{\prime}\in[G_{obs}]$.
+$P(G=g|Z=Z_{obs},G\in[G_{obs}])=\frac{P(Z=Z_{obs}|G=g,G\in[G_{obs}])\cdot P(G=g|G\in[G_{obs}])}{P(Z=Z_{obs}|G\in[G_{obs}])}.$ (8)
 $P(G=g|G\in[G_{obs}])$ $=\frac{f_{p_{0},p_{1},...,p_{N-1}}(deg(g))}{\sum_{G\in[G_{obs}]}f_{p_{0},p_{1},...,p_{N-1}}(deg(G))},$ (9)
@@
 $=P(G=g^{\prime}|G\in[G_{obs}]).$
-$P(Z=Z_{obs}|G=g,G\in[G_{obs}])$ $=P(\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\hskip 5.69054ptj\in[N]|G=g,G\in[G_{obs}]),$ (11)
-$=P(W=W_{obs}|G=g,G\in[G_{obs}]),$
-$=P(W=W_{obs})$
-$=P(W=W_{obs})|G=g^{\prime},G\in[G_{obs}]),$
+$P(Z=Z_{obs}|G=g,G\in[G_{obs}])$ $=\frac{1}{\binom{N}{\sum Z_{obs}}},$ (10)
 $=P(Z=Z_{obs}|G=g^{\prime},G\in[G_{obs}]).$
-Note that we take cluster partition to be fixed as part of the experimental design. Using Equation (10) and (11) in Equation (9), we obtain
-$P(G=g|Z=Z_{obs},G\in[G_{obs}])=P(G=g^{\prime}|Z=Z_{obs},G\in[G_{obs}]),$
-Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}]\}$ is uniform under cluster randomized experiment. Hence,
+Using Equation (9) and (10) in Equation (8), we obtain
+$P(G=g|Z=Z_{obs},G\in[G_{obs}])=P(G=g^{\prime}|Z=Z_{obs},G\in[G_{obs}]).$
+Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}]\}$ is uniform under completely randomized experiment. Therefore,
 $p_{Z_{obs},G_{obs}}=\frac{1}{|[G_{obs}]|}\cdot\sum_{G\in[G_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})).$
+Hence, we prove that Procedure $2$ is valid. We now prove the validity of Procedure $3$. Let
+$[G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\>|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
+Here, ${Z_{obs}}^{t}=\{i\in\mathbb{P}:\>{Z_{obs}}_{i}=t\}$ for $t\in\{0,1\}$.
+Note that $[G_{obs}|Z_{obs}]\subseteq[G_{obs}]$, where $[G_{obs}]=\{G\in\mathcal{G}|G\cong G_{obs}\}$.
+We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}|Z_{obs}])$ is a uniform distribution under a cluster randomized experiment.
+Let $C$ be a random variable defined as $\mathcal{A}\circ G$.
+Here, $\mathcal{A}$ represents the clustering algorithm in Algorithm 1.
+Consider $g,g^{\prime}\in[G_{obs}|Z_{obs}]$. Then $P(G=g|Z=Z_{obs},G\in[G_{obs}|Z_{obs}])=$
+$\frac{P(Z=Z_{obs}|G=g,G\in[G_{obs}|Z_{obs}])\cdot P(G=g|G\in[G_{obs}|Z_{obs}])}{P(Z=Z_{obs}|G\in[G_{obs}|Z_{obs}])}.$ (11)
+$P(G=g|G\in[G_{obs}|Z_{obs}])$ $=\frac{f_{\eta}(s(g))}{\sum_{G\in[G_{obs}|Z_{obs}]}f_{\eta}(s(G)},$ (12)
+$=\frac{f_{\eta}(s(\pi(G_{obs})))}{\sum_{G\in[G_{obs}|Z_{obs}]}f_{\eta}(s(\pi^{\prime}(G_{obs}))},$
+$=\frac{f_{\eta}(s(G_{obs}))}{\sum_{G\in[G_{obs}|Z_{obs}]}f_{\eta}(s(G_{obs})},$
+$=P(G=g^{\prime}|G\in[G_{obs}]).$
+$P(Z=Z_{obs}\>|\>G=g,$ $G\in[G_{obs}|Z_{obs}])$ (13)
+$=P(\sum_{i=1}^{|C|}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\;j\in[N]\>|\>G=g,G\in[G_{obs}|Z_{obs}])$
+$=P(\sum_{i=1}^{|\pi(C_{obs})|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))={Z_{obs}}_{j}\;j\in[N]\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
+$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
+$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(\pi^{-1}(j)\in{C_{obs}}_{i})=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
+$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in{C_{obs}}_{i})=t\quad\forall\pi(j)\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
+$=P(W=W_{obs})=P(Z=Z_{obs}|G=g^{\prime},G\in[G_{obs}|Z_{obs}]).$
+Using Equation (12) and (13) in Equation (11), we obtain
+$P(G=g|Z=Z_{obs},G\in[G_{obs}|Z_{obs}])=P(G=g^{\prime}|Z=Z_{obs},G\in[G_{obs}|Z_{obs}]).$
+Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}|Z_{obs}]\}$ is uniform under cluster randomized experiment.
+Therefore,
+$p_{Z_{obs},G_{obs}}=\frac{1}{|[G_{obs}|Z_{obs}]|}\cdot\sum_{G\in[G_{obs}|Z_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})).$
+Hence, we prove that Procedure $3$ is valid. ∎
 Appendix C Additional numerical simulations
-We present the data-generating process considered in the simulations.
-To give an overview, we consider a linearly separable potential outcome model for the outcome function, which is a composition of two effects: a direct treatment effect and a spillover treatment effect.
+We present the Erdős-Rényi random graph model and the data-generating process considered in the simulations here.
+To give an overview, we consider a linearly separable potential outcome model for the outcome function, composed of two effects: a direct treatment effect and a spillover treatment effect.
 The spillover parameter is weighted by a function of the unit’s neighbor.
 We consider two such outcome models for the simulations: one that captures spillovers by the presence of a treated neighbor and the other that captures it via the proportion of treated neighbors.
-C.1   Data generating mechanism
+C.1 Erdős-Rényi random graph model
+We formally describe the Erdős-Rényi random graph model (Erdős et al.
+[1960]).
+The below-described model is the probabilist variant of the (Erdős-Rényi) random graph model.
+We consider the number of nodes, $N$, and edges, $|E|$, to be fixed, and any two nodes will be linked independently with an edge with a fixed probability $p$.
+That is,
+$P(G=G_{obs}|p)=\binom{\binom{N}{2}}{|E|}\cdot p^{|E|}\cdot(1-p)^{\binom{N}{2}-|E|}.$ (14)
+To model $P(G)$ using Equation (14), we have to estimate the parameter $p$.
+We obtain the MLE of the parameter $p$ and estimate it as $\frac{|E|}{\binom{N}{2}}$.
+We conduct an extensive numerical simulation study and verify the finite sample validity of the proposed randomization test.
+We also obtain the power of the test under different scenarios.
+An Erdős-Rényi random graph model is the simplest network model one can consider, and it may not be representative of the kind of networks we observe in the real world.
+For example, one observes a high degree of clustering in social networks in the graph (Watts and Strogatz [1998]).
+One drawback of the model is the independence and identical probability of formation of two edges in the graph.
+One way to characterize this is to evaluate the degree distribution of the random graph.
+It has been shown that the degree distribution of a large Erdős-Rényi random graph is a Poisson distribution.
+Most real-life networks violate this property and have highly skewed degree distribution.
+In the main text, we generalize this and characterize random graphs with arbitrary degree distribution.
+C.2 Data generating mechanism
 The following describes the two potential outcome models:
-$Y_{i}(\textbf{z})=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}>0)+\epsilon_{i},$ (12)
+$Y_{i}(\textbf{z})=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot z_{j}>0)+\epsilon_{i},$ (15)
 $\epsilon_{i}\sim\mathcal{N}(0,1)\qquad\forall i\in[N].$
@@
 Here, $N_{t}$ is fixed and pre-determined by the experimenter.
-In the performed simulations, we have 100 units and 5 treated units in the model in one case.
+In the performed simulations, we have 100 total units and 5 treated units in the model in one case.
 We then extended the study to 500 units, and 250 were treated in the second case.
@@
     Direct treatment effect: We take $\tau_{direct}=4$ in our simulations.
-For robustness, we also perform the simulations with $\tau_{direct}=1$.
+We also perform the simulations with $\tau_{direct}=1$ for robustness.
 It is noted that $\tau_{direct}$ has no impact on the validity and power of the test.
   3. 3.
-    Spillover effect: We vary the spillover effect, $\tau_{spill}$ from 0 to 7, keeping $\tau_{direct}$ fixed.
+    Spillover effect: We vary the spillover effect, $\tau_{spill}$ from 0 to 5, keeping $\tau_{direct}$ fixed.
   4. 4.
     Test statistic: We consider the test statistic taking the difference of treatment effect on control units with control neighbors and control units with treated neighbors, $T_{I_{c}}(G|Z_{obs})$.
-We also consider the test statistic based on the difference in treatment effect of a low proportion of treated neighbors and a high proportion of treated neighbors, $T_{{quant}_{c}}(G|Z_{obs})$.
+We also consider the test statistic based on the difference in treatment effect of a low proportion of treated neighbors and a high proportion of treated neighbors, $T_{quant}(G|Z_{obs})$.
 Refer to Section 4.3 for further details.
@@
 If the p-value is less than 0.05, we reject the null hypothesis of no spillover effects.
-We replicated the above 480 times and calculated rejection rates as the proportion of times the p-value was significant.
-For this setup, the outcome model (12) with $\tau_{direct}$ equals to 4 and $\tau_{spill}$ is presented in Table 3.
+We replicated the above 450 times and calculated rejection rates as the proportion of times the p-value was significant.
+For this setup, the outcome model (15) with $\tau_{direct}$ equals to 4 and $\tau_{spill}$ is presented in Table 3, varying from $0.0$ to $0.7$.
 We then repeat the simulations for the above setup for outcome model (16) and present them in Table 4.
@@
 Another caveat in the above setting is the nature of test statistic $T_{I_{c}}(G|Z_{obs})$.
-Depending on the input graph, we might not get any control units with all neighboring ones in control, invalidating the procedure.
+We might not get any control units with all neighboring ones in control depending on the input graph, invalidating the procedure.
 So, the above setup involves 100 units with 5 units in treatment.
 We numerically found the procedure failing when the number of treated units was increased to more than 5.
-We present the results for model (13) with $T_{{quant}_{c}}$ as test statistic in Table 4.
+We present the results for model (16) with $T_{quant}$ as test statistic in Table 4.
 Here, the number of units in the population is 500, with the number of treated units being 250.
@@
 In all the above models, we took the probability of an edge between two nodes as $0.2$.
-We present the same setup in 4, with the probability of an edge between two nodes estimated as below.
+We present the same setup in 4, with the probability of an edge between two nodes estimated below.
 $G_{obs}$ is generated from Erdős-Rényi random graph with $p=0.2$.
 $\hat{p}=\frac{|E_{G_{obs}}|}{\binom{N}{2}}.$
-Here, $|E_{G_{obs}}|$ represents number of edges in the observed graph $G_{obs}$. The results are presented in Table 4.
-We highlight how type 1 error is controlled under 0.05 in all the setups (case when $\tau_{spill}$ is zero).
-When the null hypothesis is not true, we observe good power against the test in line with results from Athey et al.
-[2018].
-This remains true when the network distribution is estimated parametrically.
-Table 3: Rejection rates for null hypothesis of no spillover effects; Outcome model (12) with test statistic $T_{I_{c}}$
+Here, $|E_{G_{obs}}|$ represents number of edges in the observed graph $G_{obs}$.
+The results are presented in Table 4.
+We highlight how type 1 error is controlled under 0.05 in all the setups (when $\tau_{spill}$ is zero).
+When the null hypothesis is false, we observe moderate power.
+This holds when the network distribution is estimated parametrically.
+Table 3: Rejection rates for the null hypothesis of no spillover effects. Outcome model (15) with test statistic $T_{I_{c}}$.
 $\tau_{spill}$ Rejection Rate
-0.0 0.035
-0.1 0.083
-0.2 0.173
-0.3 0.279
-0.4 0.402
-0.5 0.579
-0.6 0.745
-0.7 0.887
-0.8 0.943
-0.9 0.987
-1.0 0.993
-Table 4: Rejection rates for null hypothesis of no spillover effects; Outcome model (13) with test statistic $T_{I_{c}}$, $T_{{quant}_{c}}$ and $T_{{quant}_{c}}$ with estimated Erdős-Rényi random graph
+0.0 0.038
+0.1 0.077
+0.2 0.167
+0.3 0.293
+0.4 0.444
+0.5 0.620
+0.6 0.749
+0.7 0.858
+Table 4: Rejection rates for the null hypothesis of no spillover effects. Outcome model (16) with test statistic $T_{quant}$, and $T_{quant}$ with estimated Erdős-Rényi random graph.
 $\tau_{spill}$ Rejection Rate
-$T_{I_{c}}$ $T_{{quant}_{c}}$ $T_{{quant}_{c}}$ with $\hat{p}$
-0 0.054 0.040 0.054
-1 0.058 0.080 0.069
-2 0.110 0.238 0.229
-3 0.181 0.419 0.419
-4 0.256 0.694 0.693
-5 0.365 0.881 0.881
-6 0.485 0.954 0.954
-7 0.612 0.990 0.994
+$T_{quant}$ $T_{quant}$ with $\hat{p}$
+0 0.057 0.044
+1 0.124 0.138
+2 0.402 0.469
+3 0.747 0.729
+4 0.947 0.927
+5 0.984 0.989
+Appendix D Performance of test statistic $T_{bond}$ on Procedure 2
+Figure 2: Distribution of p-values for a completely randomized experiment under Procedure 2 (isomorphism permutation variant) over 2,000 replications; the plots on the top have direct effect 0 with test statistics $T_{bond}$ on the left panel and $T_{quant}$ on the right.
+The plots on the bottom have a direct effect 4.
+All plots have a spillover effect of 0.4.
+The (blue) line indicates the threshold of 0.05.
+[Lower panel, left plot] The closer-to-uniform distribution of the p-values for $T_{bond}$ contributes to the drop in power as observed in Table 1.

### v3

@@
 An interesting approach to tackling the non-sharp nature of the null hypothesis in this setup is constructing conditional randomization tests such that the null is sharp on the restricted population.
-Such approaches can pose computational challenges as finding these appropriate sub-populations based on experimental design can involve solving an NP-hard problem.
-In this paper, we view the network amongst the population as a random variable instead of being fixed.
+In randomized experiments, conditional randomized tests hold finite sample validity and are assumption-lean.
+In this paper, we incorporate the network amongst the population as a random variable instead of being fixed.
 We propose a new approach that builds a conditional quasi-randomization test.
-We build the (non-sharp) null distribution of no spillover effects using random graph null models.
+To build the (non-sharp) null distribution of no spillover effects, we use random graph null models.
 We show that our method is exactly valid in finite samples under mild assumptions.
-Our method displays enhanced power over other methods, substantially improving cluster randomized trials.
+Our method displays enhanced power over state-of-the-art methods, with a substantial improvement in cluster randomized trials.
 We illustrate our methodology to test for interference in a weather insurance adoption experiment run in rural China.
-Keywords: Causal Inference, Network Interference, Peer Effects, Randomization Tests, Random Graphs
+Keywords: Causal Inference, Network Interference, Peer Effects, Randomization Tests, Random Graphs.
 1 Introduction
@@
 (2015)).
-Due to social network ties in the population, information dissemination is a plausible interference mechanism that can occur and affect evaluated estimates.
+Due to social network ties in the population, information dissemination is a plausible interference mechanism that can occur and affect the evaluated estimates.
 Similar phenomena can be seen amongst experiments run on social media networks, like LinkedIn, or online marketplaces like eBay, owing to the interconnected nature of these platforms (Pouget-Abadie et al.
@@
 If one were to test for spillover effects, it would render the null non-sharp, leading to imputability issues for the test statistic.
-There is a growing body of literature that addresses this issue by considering a carefully determined subset of units in the population and treatment assignment vector, such that on these selected units and treatment vectors, the null is sharp (Aronow and Samii (2017); Athey et al.
+There is a growing body of literature that addresses this issue by considering a fixed subset of units in the population and treatment assignment vector, called ‘focal units’, such that on these focal units the null is sharp (Aronow and Samii (2017); Athey et al.
 (2018); Basse et al.
@@
 Basse et al.
-(2019b) build a permutation test under the particular case of clustered interference.
-While a permutation test is computationally cheaper, the method lacks generalizability because of the structured nature of the interference considered.
+(2019b) builds a permutation test under clustered interference.
 Aronow and Samii (2017) and Athey et al.
@@
 Puelz et al.
-(2022) presents a graph-theoretic approach to constructing the conditioning mechanism of a sub-population of units and the corresponding treatment assignment vectors for which the null is sharp.
-However, the proposed methodologies lose power in cluster randomized experiments.
+(2022) presents a graph-theoretic approach to constructing the conditioning mechanism of selecting the focal units and the corresponding treatment assignment vectors for which the null is sharp.
+While the above considers the problem of testing for interference in an experimental setting, the literature on testing for spillover effects in an observational setting is still in its infancy.
 This paper considers the network to be coming from a distribution itself.
-This idea contrasts with the setup considered in the current methods, assuming the network is fixed.
+This idea expands upon the setup considered in the current methods, where the network is taken to be fixed.
 We draw a parallel here to estimating the probability of unit treatment assignment via propensity scores in the case of observational studies.
@@
 The current methods are based on strict randomization inference with no distributional assumptions for either the network or the potential outcomes, achieving a generalization of the Fisher Randomization Test.
-This allows for the widespread validity of the methods proposed, albeit it might be at the cost of power.
-This paper explores how leveraging appropriate stochastic assumptions can significantly boost power and lead to methodological advancements.
+This allows for the widespread validity of the methods proposed.
+Since the experimenter cannot randomize this network, we explore how to incorporate appropriate stochastic assumptions on the network in the context of a conditional randomization test.
 We assume that a given vector of sufficient statistics characterizes the network data-generating process and can thus be represented as an exponential random graph model.
@@
 This assists us in building a quasi-randomization test based on the network distribution by randomizing over graphs from this generating process.
-We show that conditional on the observed treatment, a test statistic’s (quasi) randomization distribution leads to valid p-values.
+We show that, conditional on the observed treatment, a test statistic’s (quasi) randomization distribution leads to valid p-values.
 We then generalize the model to a vector of sufficient statistics and present a method that works for any set of such network statistics.
@@
 More recently, Li and Wager (2022) and Leung (2020) proposed estimation strategies for spillover effects under a distributional assumption on the network formation model.
-The inferential methods for the proposed estimation procedures involve non-standard or asymptotic approximations (Imbens and Manski (2004)).
-With a careful identification of causal estimates, such large sample normal approximations of confidence interval conform to a Neymanian perspective.
-Our first contribution is to extend the random graph interference model into the potential outcome framework.
+The inferential methods for the proposed estimation procedures involve asymptotic approximations (Imbens and Manski (2004)).
+Such large sample normal approximations of confidence interval conform to a Neymanian perspective.
+In line with Fisherian procedures, our first contribution is to extend the random graph interference model into the potential outcome framework.
 We generalize and define sharp/non-sharp null hypotheses for random graph models.
 Building upon this, we develop a randomization-based test for a given random graph model.
-This leads to finite sample validity of the p-value obtained, in line with Fisherian procedures.
-As a second contribution, our method addresses several technical challenges encountered in the current methods.
-We show that randomization tests based on network-generating processes lead to imputable test statistics under the null, conditional on the treatment assignment.
-This overcomes a significant challenge in randomization methods based on random treatment assignment.
-The state-of-the-art techniques restrict attention to a carefully chosen subset of focal units and claim null of no spillover effects for only these units.
-In contrast, our method tests for a stricter class of nulls with no spillover effects on all the units.
-Another hurdle the current methodologies face is the loss of power in cluster randomized trials.
-Determination of conditional randomization distribution given a conditioning mechanism is not straightforward.
+This leads to finite sample validity of the p-value obtained.
+As a second contribution, we show that randomization tests based on network-generating processes lead to imputable test statistics under the null, conditional on the treatment assignment and network covariates.
+This overcomes a technical challenge in randomization methods based on random treatment assignment, where the null of no spillover effects is tested only upon the focal units (Athey et al.
+(2018)).
+In contrast, our method tests for a stricter class of nulls with no spillover effects on the given population.
+Another hurdle the current methodologies face is low power.
+Determination of the conditional randomization distribution given a conditioning mechanism is not straightforward.
 The restriction of fixed focal units in the randomization null reduces the sample space of possible treatment assignment vectors in the null.
 It can lead to an enormous loss of power or render the null degenerate completely.
-A burgeoning literature set focuses on improving current methods by incorporating design-based techniques (Puelz et al.
+A burgeoning literature set focuses on improving current methods by incorporating design-based techniques for selecting conditioning mechanisms (Puelz et al.
 (2022); Basse et al.
 (2019a)).
-Unlike distribution-free methods, we show how incorporating information about network generation into the methodology can boost power.
-Our third significant contribution is the computational efficiency of the proposed method.
+We show how incorporating information about network generation into the methodology can aid in power.
+Our methods are also computationally feasible.
 We suggest that sufficient network statistics can characterize the random graph model.
@@
 (2018)).
-Thus, we obtain valid randomization tests via permutation of edges of the graph in a manner that preserves degree sequence.
-We then present a method for a general exponential random graph model characterized by an arbitrary set of network statistics and build a permutation test by generating graphs in the same isomorphism class under the null.
-This is closely related to the literature on the random graph null model used in network analysis with applications in protein structure discovery and community detection among many (Milenković et al.
+Thus, we obtain valid randomization tests via the permutation of edges of the graph in a manner that preserves the degree sequence.
+We then present a method for a general exponential random graph model characterized by an arbitrary set of network statistics and build a permutation test by generating degree-preserving graphs in the same isomorphism class under the null.
+This is closely related to the literature on the random graph null model used in network analysis with applications in protein structure discovery and community detection among many (Milenkovic et al.
 (2009); Sah et al.
@@
 We display our methodology on a subset of the original experiment performed and obtain significant information spillover among the farmers.
-The remaining paper is structured as follows.
-In Section 2, we set up notation and formalize the null hypothesis of interest, incorporating network random variable.
-Section 3 presents a randomization test based on a given network-generating process and its application to different experimental designs.
-Section 4 extends this further to random graph models and presents our central methodology introducing the permutation test.
-In Section 5, we present extensive numerical simulations to evaluate the performance of the proposed test.
-Sections 6 and 7 illustrate our methodology as an empirical application and conclude the paper.
-Proofs, algorithms, and additional numerical results are presented in the Appendix.
 2 Setup
 2.1 Notation
-Consider a population =$\mathbb{P}=\{1,2,...,N\}$ where $N$ is the total number of units in the population.
+Consider a population $\mathbb{P}=\{1,2,...,N\}$ where $N$ is the total number of units in the population.
 Consider a vector $Z\in{\mathbb{Z}}^{N}$ where each element in the vector indicates treatment assigned to the corresponding indexed unit in the population.
 Here, $\mathbb{Z}$ represents the set of possible treatments to which the population units can be exposed.
-We take =$\mathbb{Z}=\{0,1\}$ for this article.
+We take $\mathbb{Z}=\{0,1\}$ for this article.
 Hence, any treatment assignment vector is an N-dimensional binary vector.
@@
 That is, $G[\mathbb{P^{\prime}}]=(\mathbb{P^{\prime}},E^{\prime})$ where $E^{\prime}=\{(i,j)\in\mathbb{P^{\prime}}\times\mathbb{P^{\prime}}|\>(i,j)\in E\}$.
-Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$ where $X_{i}$ is a m-dimensional vector.
+Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$, where $X_{i}$ is an m-dimensional vector.
 $X_{i}$ is fixed and influences the outcome via the network’s potential outcome function or formation.
@@
 We define the potential outcome function $Y:\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow\mathbb{R}^{N}$ as a real-valued function that takes treatment assignment and network adjacency matrix as arguments.
-We will consider $Y$ to be fixed, and each element $i$ of an image in $Y(Z,G)$, $Y_{i}(Z,G)$ denotes the outcome of $i^{th}$ unit in the population under the treatment assignment $Z$ and network assignment $G$.
+We will consider $Y$ to be fixed, and each element $i$ of an image in $Y(Z,G)$, $Y_{i}(Z,G)$, denotes the outcome of the $i^{th}$ unit in the population under the treatment assignment $Z$ and network assignment $G$.
 We denote the corresponding potential outcome as $Y_{obs}$ that is, $Y_{obs}=Y(Z_{obs},G_{obs})$.
 We note that the potential outcome function implicitly contains information in $X$ and can be written as $Y_{X}(Z,G)$.
-For the remainder of the article, we will use the notation $Y(Z,G)$ for readability.
+For the remainder of the article, we will use the notation $Y(Z,G)$ for readability unless deemed necessary.
 2.2 Definitions
@@
 The following are three key assumptions that are considered in the causal inference literature for assignment mechanism: (i) individualistic assignment, which restricts the treatment assignment of a unit to its covariates and outcomes (ii) probabilistic assignment, which assumes non-integer probabilities of treatment assignment to units (iii) unconfoundedness assumption, which restricts the dependence of the probability assignment of units to only covariates, and not the potential outcome (Imbens and Rubin (2015)).
-Given a well-defined potential outcome function, individualistic assignment mechanism assumption constitutes Stable Unit Treatment Value Assumption (SUTVA).
+Given a well-defined potential outcome function, the individualistic assignment mechanism assumption constitutes Stable Unit Treatment Value Assumption (SUTVA).
 We argue that the underlying network structure among the population is an implicit form of treatment.
@@
 We assume that the joint treatment assignment is unconfounded; that is,
-$P(Z,G|X,Y(0),Y(1))=P(Z,G|X).$
+$P(Z,G|X,Y(0),Y(1))=P(Z,G|X).$ (1)
+Remark 1.
+It should be noted that the graph $G$ only maps the interference structure in the population.
+Any other dependence of the potential outcome function on the network characteristics and other confounding factors comprises the covariates in the study.
+We denote the covariates corresponding to network statistics as $G_{X}$, which may be dependent on G.
+Apart from $G_{X}$, in this article, we restrict the potential outcome function to be covariate-free for the sake of brevity.
+Thus, Equation 1 can be written as
+$P(Z,G|X,Y(0),Y(1))$ $=P(Z,G|G_{X},Y(0),Y(1)),$
+$=P(Z,G|G_{X}).$
+For further details on covariate balancing in randomized experiments, refer to Liu and Yang (2020).
 We formally define the standard sharp null hypothesis of no (direct) treatment effect when the population is not interconnected.
@@
 $\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
-[…]
+$\text{s.t.}\quad Z\cdot D_{G}^{(k-1)}(i)=Z^{\prime}\cdot D_{G^{\prime}}^{(k-1)}(i).$
 Note the non-imputability of the potential outcome function under the null considered above for different treatment assignments.
@@
 In this paper, we are concerned with testing first-order spillover effects. We take $k=1$ in Example 1 and state the null hypothesis of no spillover effects below.
-{mdframed}
-[leftmargin=0.1in,rightmargin=0.1in]
 Hypothesis 4.
-(Null hypothesis of no spillover effects)
-$\forall i\in[N]\quad Y_{i}(Z,G)=Y_{i}(Z^{\prime},G^{\prime})$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
-[…]
-We highlight that the treatment $G$ is not restricted in the above hypothesis.
+(Null hypothesis of no spillover effects) $\forall i\in[N]\quad{Y_{i}}_{G_{X_{i}}}(Z,G)={Y_{i}}_{G^{\prime}_{X_{i}}}(Z^{\prime},G^{\prime})$ $\forall(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$ $\text{s.t.}\quad Z_{i}=Z_{i}^{\prime},{G_{X}}_{i}={G^{\prime}_{X}}_{i}.$
+The treatment $G$ is unrestricted here, fixing the graph covariates.
+We claim that the null hypothesis, in the above, encodes no spillover effects (see Remark 1).
 We now describe the spillover mechanism as a component in the potential outcome function $Y(Z,G)$.
-In the following example, $Y(Z,G)$ is linearly separable in two components, one of direct effect and another of spillover effect, as stated below.
-We re-state the null hypothesis of no spillover effects, as described in Hypothesis 4.
+In the following example, $Y(Z,G)$ is linearly separable into two components, one of direct effect and another of spillover effect, as stated below.
+We restate the null hypothesis of no spillover effects, as described in Hypothesis 4.
 Example 2.
@@
 $\forall i\in[N]\quad Y^{2}(Z_{\mathcal{N}_{G}(i)})=Y^{2}(Z^{\prime}_{\mathcal{N}_{G^{\prime}}(i)})$ $\forall\quad(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$
-[…]
-This approach of defining non-sharp null of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al.
+$\text{s.t.}\quad Z_{i}=Z_{i}^{\prime}.$
+This approach of defining non-sharp nulls of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al.
 (2019b); Puelz et al.
@@
 $f_{G}(Z)=\frac{Z\cdot D_{G}^{(k-1)}(i)}{|D_{G}^{(k-1)}(i)|}.$
-We obtain the setup described in the literature, given a fixed $G$ using the above-described exposure function as a restriction set.
+We obtain the setup described in the literature, given a fixed $G$, using the above-described exposure function as a restriction set.
 The setup assumes a spillover mechanism via an exposure function.
 Since we consider the graph underlying the population to be random, we define restrictions on the joint treatment assignment in Hypothesis 3.
-A spillover mechanism, if existing, will be captured in the potential outcome function.
-We describe the currently proposed method in the literature of a conditional-randomization test procedure to obtain valid p-values (Puelz et al.
+A spillover mechanism, if it exists, will be captured in the potential outcome function.
+We describe the currently proposed method in the literature of a conditional-randomization test procedure to obtain valid p-values using ‘focal units’ (Puelz et al.
 (2022)).
-Procedure 0: Consider $C$ as a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|C)$ is imputable under the null.
+Procedure 0: Consider $F$ as a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|F)$ is imputable under the null.
   1. 1.
@@
   2. 2.
-    Draw $C\sim P(C|Z_{obs})$ and compute $T(Z_{obs}|C)$.
+    Draw $F\sim P(F|Z_{obs})$ and compute $T(Z_{obs}|F)$.
   3. 3.
-    Compute p-value as $P_{Z|C}\left(T(Z|C)>T(Z_{obs}|C)\right)$.
+    Compute p-value as $P_{Z|F}\left(T(Z|F)>T(Z_{obs}|F)\right)$.
 The above-stated procedure holds finite-sample validity.
-The conditioning mechanism $C$ is drawn from $P(C|Z_{obs})$, which the analyst should determine.
-One must ensure the imputability of the test statistic $T(Z|C)$ when constructing a mechanism for $C$.
+The conditioning mechanism $F$, termed as focal units, is drawn from $P(F|Z_{obs})$, which the analyst should determine.
+We consider the random selection method as the conditioning mechanism to obtain the focal units.
+Procedure 0 tests for deviation from the null hypothesis of no spillover effect on focal units.
+This ensures the imputability of the test statistic $T(Z|F)$.
 We require the test statistic to be imputable to obtain its sampling distribution under the null, even though we observe only one treatment assignment vector.
@@
 $T(Z,G,Y(Z,G))$ $=T(Z,G,Y(Z^{\prime},G^{\prime}))$
-$\quad P(Z,G|(Z,G)\in\textbf{Z}\times\textbf{G}\text{ },\mathcal{H}_{0}),P(Z^{\prime},G^{\prime}|(Z^{\prime},G^{\prime})\in\textbf{Z}\times\textbf{G}\text{ },\mathcal{H}_{0})>0.$
+[…]
 The choice of a (imputable) test statistic is not unique, though, and different test statistics can lead to different powers of the test.
@@
 Our main idea is to generate the null sampling distribution of a test statistic by randomizing over the network-generating process.
-Intuitively, the underlying network structure does not contribute to the total treatment effect under the null of no spillover effect.
-This can also be seen in Hypothesis 4 where the network variable $G$ is unrestricted in the, otherwise, non-sharp null.
-This leads to the imputability of a test statistic of choice by fixing the direct treatment assignment $Z$ and randomizing over the network distribution $G$.
+Intuitively, the underlying network structure also contains information about the spillover effect.
+This can also be seen in Hypothesis 4 where the network variable $G$ is unrestricted in the, otherwise, non-sharp null keeping the network covariates fixed.
+Fixing direct treatment assignment and network covariates leads to the imputability of a test statistic of choice.
+To strictly generalize the conditional randomization test over focal units (Athey et al.
+(2018)), one can randomize over the network distribution of $G_{F^{c}}$, that is, over the non-focal units in the population.
+One feature of this generalization is that it ensures the network characteristics of the focal units remain fixed.
+One can, then, also randomize over the joint treatment assignment vector $(Z,G)$.
+This expands the state-of-the-art methods over the bivariate treatment assignment vector.
+The following method focuses on the conditional treatment assignment distribution of $G|Z$.
 We formally define the conditional test statistic, $T_{c}$, below and show its imputability under the null Hypothesis 4.
 Proposition 1.
-Define $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$. $T_{c}(G|Z_{obs})$ is imputable under the null hypothesis of no spillover effects in Hypothesis 4.
+Define $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{{G_{X}}_{obs}}(Z_{obs},G))$.
+$T_{c}(G|Z_{obs})$ is imputable under the null hypothesis of no spillover effects in Hypothesis 4 and is equal to $T(Z_{obs},G,Y_{obs})$.
+Here, $Y_{obs}=Y_{{G_{X}}_{obs}}(Z_{obs},G_{obs})$.
 Proof.
-Note that under $\mathcal{H}_{0}$,
-$Y(Z,G)=Y(Z^{\prime},G^{\prime})\qquad Z=Z^{\prime}.$
-Thus,
-$T_{c}(G|Z_{obs})$ $=T(Z_{obs},G,Y(Z_{obs},G)),$
-$=T(Z_{obs},G,Y(Z_{obs},G^{\prime}))\qquad P(G^{\prime}|Z_{obs})>0\text{ under }\mathcal{H}_{0}.$
-Since $Z_{obs}$ is arbitrary, we conclude imputability of the $T_{c}(G|Z_{obs})$ under the null $\mathcal{H}_{0}$. ∎
+The proof is given in Supplementary material B.1
 We state our proposed quasi-randomization testing procedure below.
-The procedure assumes the knowledge of the network-generating process $P(G)$.
+The procedure assumes the knowledge of the network-generating process $P(G|G_{X})$.
 An elaborate consideration of the network distribution will be presented in Section 5.
@@
   3. 3.
-    Consider the randomization distribution $P(G|Z)$.
+    Consider the randomization distribution $P(G|Z,G_{X})$.
   4. 4.
@@
 Proof.
-The proof is given in Appendix B.1. ∎
+The proof is given in Supplementary material B.3. ∎
+Remark 2.
 Note that conditional validity implies unconditional validity of the procedure, too.
-A crucial component of the proof is identifying the conditional distribution $P(G|Z)$ for the validity of Procedure 1.
+That is, Procedure 1 holds finite sample validity with respect to the bivariate joint distribution of the treatment assignment vector $(Z,G)$, following the law of iterated expectations.
+A crucial component of the proof is identifying the conditional distribution $P(G|Z,G_{X})$ for the validity of Procedure 1.
 This is contingent upon the experimental design, and different experimental designs will lead to different conditional randomization distributions.
-We compare this with the state-of-the-art methodology presented in Section 2 where the null randomization distribution comes from $P(Z|C)$.
-Here, $C$ is a conditioning mechanism consisting of a subset of chosen units.
-Thus, different experimental designs lead to different $P(Z|C)$.
-A common problem faced in the literature is implementing the methods in cluster randomized experiments.
+We compare this with the state-of-the-art methodology presented in Section 2 where the null randomization distribution comes from $P(Z|F)$.
+Here, $F$ represents focal units.
+A common problem faced in the literature is the loss of power when implementing the methods in cluster randomized experiments.
 In a cluster-randomized experiment, the population is classified into clusters where treatments are assigned at the cluster level.
-If a cluster is treated, all units in that cluster receive the treatment or vice-versa.
-Fixing the treatment assignment of units in the conditioning mechanism might render the null randomization distribution $P(Z|C)$ degenerate, given the restricted span of cluster treatment assignments space.
+If a cluster is treated, all units in that cluster receive the treatment or vice versa.
+Fixing the treatment assignment of units in the conditioning mechanism might render the null randomization distribution $P(Z|F)$ degenerate, given the restricted span of cluster treatment assignment space.
 For example, assume there is a focal unit in every cluster.
-Then, conditioning on the observed treatment of focal units, we obtain $Z|C=\{Z_{obs}\}$ since all the unit treatment assignments can be deterministically imputed by the treatment assignment of the focal unit in that cluster.
-This reduction in the support of conditional distribution leads to either degeneracy or enormous loss of power.
-Our methodology captures the null distribution by randomizing over $G|Z$, where the null distribution admits ample support.
-We discuss this in Sections 3.1.1 and 3.1.2, where we identify the null randomization distribution under different experimental designs.
-3.1 Experimental design
+Then, conditioning on the observed treatment of focal units, we obtain $Z|F=\{Z_{obs}\}$ since all the unit treatment assignments can be deterministically imputed by the treatment assignment of the focal unit in that cluster.
+This reduction in the support of the conditional distribution leads to either degeneracy or enormous loss of power.
+To get over this, our methodology captures the null distribution by randomizing over the conditional distribution $G|(Z,G_{X})$, where the null distribution admits ample support.
+We discuss this in Sections 4.1 and 4.2, where we identify the null randomization distribution under different experimental designs.
+4 Experimental design
 Here, we discuss how Procedure 1 can be implemented in different experimental designs.
 A change in experimental design corresponds to a change in (direct) treatment assignment mechanism.
-Thus, correctly identifying the distribution $P(G|Z)$ is important to ensure the validity of Procedure 1.
+Thus, correctly identifying the distribution $P(G|Z)$ (and $P(G|Z,G_{X})$) is important to ensure the validity of Procedure 1.
 We find $P(G|Z)$ for two experimental designs, completely randomized experiments and cluster randomized experiments.
-3.1.1 Completely randomized experiment
-Consider a random sample of units of size $N_{t}$ from $\mathbb{P}$, who are selected in the treatment group.
+4.1 Completely randomized experiment
+Consider a random sample of units of size $N_{t}$ from $\mathbb{P}$, which are selected in the treatment group.
 Let $N_{t}$ be fixed.
@@
 $P(Z=z\;|\;G)=\begin{cases}\frac{1}{\binom{N}{N_{t}}}&\sum_{i=1}^{N}z_{i}=N_{t},\\ 0&\text{otherwise}.\end{cases}$ (2)
-We remark that $G\!\perp\!\!\!\perp Z$ by design. Hence, for completely randomized experiments,
-$P(G|Z)=P(G).$
-Since $P(G)$ is not identified by design, we will explore random graph models in Section 4.1.
-3.1.2 Cluster randomized experiment
-Consider the following partition of the population $\mathbb{P}$ into clusters $\{C_{1},C_{2},....,C_{k}\}$. That is,
-$C_{i}\cap C_{j}=\emptyset$ $\hskip 14.22636pt\forall i\neq j\in[k],$
-$\cup_{i=1}^{k}C_{i}=\mathbb{P}.$
-Given such a partition, treatment assignment is done at the cluster level instead of the unit level.
+We remark that $G\!\perp\!\!\!\perp Z$ by design.
+Hence, for completely randomized experiments, $P(G|Z)=P(G).$ Since $P(G)$ is not identified by design, we will explore random graph models in Section 5.1.
+4.2 Cluster randomized experiment
+Consider the following partition of the population $\mathbb{P}$ into clusters $\{C_{1},C_{2},....,C_{k}\}$.
+That is, $C_{i}\cap C_{j}=\emptyset\forall i\neq j\in[k],\ \ \cup_{i=1}^{k}C_{i}=\mathbb{P}.$ Given such a partition, treatment assignment is done at the cluster level instead of the unit level.
 Following Ugander et al.
 (2013), we perform a Bernoulli experiment at the cluster level.
-Thus, we obtain
-$W_{i}\sim Bernoulli(p),$
-$Z_{j}=\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i}).$
-[…]
+Thus, we obtain $W_{i}\sim Bernoulli(p),\ \ Z_{j}=\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i}).$ Here, $p$ is the probability that a cluster receives treatment, and the experiment is performed on a given cluster partition of the population.
+The cluster formation is part of the design of the experiment and is fixed.
+The experimenter can form these clusters based on the structure of the underlying network.
+Benchmark graph clustering algorithms can be used to obtain the cluster partition, which depends on graph characteristics such as modularity, cuts, and other related quantities (Newman (2006)).
+We use $\epsilon$-net clustering to obtain the cluster partition of the network, previously considered in Eckles et al.
+(2017) and Ugander et al.
+(2013).
+We describe the $\epsilon$-net clustering in detail in Supplementary material A.
+While we present a deterministic version of the clustering procedure, it should be noted that any other algorithm, such as a different selection rule and the resultant $\epsilon$-net cluster output, corresponds to a change in experimental design.
+Unlike completely randomized experiments, in cluster randomized experiments, $G\not\!\perp\!\!\!\perp Z$.
+So,
 $P(G|Z)$ $=\frac{P(Z|G)\cdot P(G)}{P(Z)},$ (3)
 $=\frac{\prod_{i=1}^{k}p^{\mathcal{I}(\exists j\in[N]\text{ s.t. }j\in C_{i},z_{j}=1))}\cdot\prod_{i=1}^{k}(1-p)^{\mathcal{I}(\exists j\in[N]\text{ s.t. }j\in C_{i},z_{j}=0))}\cdot P(G)}{P(Z)}.$
-Here, we can see that identifying null randomization distribution becomes challenging in the case of graph cluster randomization experiments.
+Here, we can see that identifying a null randomization distribution becomes challenging in the case of graph cluster randomization experiments.
 We propose a new approach that considers sufficient statistics of the random graph model to build a permutation test.
 5 Permutation test via distributional properties
-Identification of null randomization distribution $P(G|Z)$ involves knowledge of the distribution of the network-generating process, which is not identified by the design of the experiment.
+Identification of null randomization distribution $P(G|Z,G_{X})$ involves knowledge of the distribution of the network-generating process, which is not identified by the design of the experiment.
 To this end, we explore random graph models to model $P(G)$.
-As discussed in the previous section, identifying null randomization distribution becomes challenging in cluster-randomized experiments.
+As discussed in the previous section, identifying the null randomization distribution becomes challenging in cluster-randomized experiments.
 This challenge is overcome in our proposed methodology, which is a conditional permutation test.
-We discuss this in Section 4.1.1.
+We discuss this in Section 5.2.
 5.1 Random graph models
@@
 The Erdős-Rényi random graph model has been studied widely, owing to exact computations of desirable graph properties.
-We present the model formally in Appendix C.1.
+We present the model formally in Supplementary material C.1.
 However, the Erdős-Rényi random graph model does not represent most real-world networks, such as social networks, supply-chain networks, and scientific collaboration networks, among many others.
 We look at a generalization of Erdős-Rényi random graph better at modeling such real-world networks, proposed by Newman et al.
-(2002), in Section 4.1.1.
-For this article, we restrict the models to be covariate-free.
-4.1.1 Random graphs with arbitrary degree distribution
+(2002), in Section 5.2.
+5.2 Random graphs with arbitrary degree distribution
 To represent more realistic degree distributions observed in real life, Newman et al. (2001) considered characterizing the random graph by its degree distribution. We formulate the model below.
@@
 Here, $\{p_{i}\}_{i=1,...,N}$ represent the degree distribution, that is, $p_{i}$ is the probability that a randomly selected unit has degree $i$ and $f(.)$ represents the probability mass function.
-We call $G$ a random graph with arbitrary degree distribution.
-A key feature of random graphs with arbitrary degree distribution is their ability to capture a large class of degree distribution.
+We call $G$ a random graph with an arbitrary degree distribution.
+A key feature of random graphs with arbitrary degree distribution is their ability to capture a large class of degree distributions.
 This generalizes the Erdős-Rényi random graph model and, at the same time, ensures tractability of important structural graph properties (Newman et al.
@@
 This involves estimating $(p_{1},p_{2},...,p_{N})$ with a single realization of the graph, and we cannot directly implement Procedure 1 without estimating $P(G)$.
-In the case of random graphs with arbitrary degree distribution, it is proven that we can estimate the parameters of the random graph with just one realization but with some uncertainty (Chatterjee et al.
+In the case of random graphs with arbitrary degree distribution, it is proven that we can estimate the parameters of the random graph with just one realization, but with some uncertainty (Chatterjee et al.
 (2011)).
@@
 For example, the phenomena of transitivity are not captured in this model, where two nodes have higher chances of being connected if they have a common connecting node.
-4.1.2 Exponential random graph model
+5.3 Exponential random graph model
 The state-of-the-art method for characterizing random graph formation of social networks is through the exponential random graph model or ERGM (Snijders et al.
@@
 (2020)).
-As with random graphs with arbitrary degree distribution, estimating ERGM is fraught with challenges given just a single realization.
+As with random graphs with arbitrary degree distribution, estimating ERGM is fraught with challenges, given just a single realization.
 Owing to estimation issues of the null distribution, we propose a new method that does not involve such estimations and reduces to a simple permutation test.
-In our proposed methodology, we first consider the sufficient statistics of the random graph distribution as degree sequence and develop a method conditioning on the sufficient statistic.
+In our proposed methodology, we first consider the sufficient statistics of the random graph distribution as the degree sequence and develop a method conditioning on the sufficient statistic.
 We then generalize the method to a set of arbitrary sufficient statistics.
 We use the symmetric nature of the $P(G)$ via an equivalence relation and build a conditional permutation test, as described below.
-4.2 Permutation test based on equiprobability events
-4.2.1 Test for a random graph with arbitrary degree distribution
+6 Permutation test based on equiprobability events
+6.1 Test for a random graph with arbitrary degree distribution
 Here, we present our main result, which proposes obtaining a conditional randomization test without estimating the distribution.
-Our main idea is to identify a sufficient statistic for the random graph model, which is degree sequence in the case of random graphs with arbitrary degree distribution, and condition on this sufficient statistic.
+Our main idea is to identify a sufficient statistic for the random graph model, which is the degree sequence in the case of random graphs with arbitrary degree distribution, and condition on this sufficient statistic.
 We sample the null distribution over all the graphs with the same degree sequence as the observed graph.
@@
 Not all $D\in\mathbb{W}^{N}$ can be a valid degree sequence.
-For example, consider $D=\{1,2\}$ as a candidate for degree sequence for a graph of size $2$.
-We remark that the maximum degree of a node in the graph is at most $(N-1)$.
+For example, consider $D=\{1,2\}$ as a candidate for a degree sequence for a graph of size $2$.
+We remark that the maximum degree of a node in the graph is at most (N1)(N-1).
 Since no vertex in a graph of size two can have a degree of more than 1, $D$ is not a valid degree sequence.
@@
 This implies that the graph’s vertices cannot be permuted to obtain the other graph with edges preserved (see Figure 1).
-Figure 1: Non-isomorphic graphs with same degree sequence
+Figure 1: Non-isomorphic graphs with the same degree sequence
 The converse holds.
@@
 Let $\mathcal{G}$ be the support of $P(G)$.
-We define a relation $d:\mathcal{G}\rightarrow\mathcal{G}$ such that $G\equiv_{d}G^{\prime}$ if $deg(G)=deg(G^{\prime})$.
+We define a relation $d:\mathcal{G}\rightarrow\mathcal{G}$ such that $G\equiv_{d}G^{\prime}$ if $deg(G)=deg(G^{\prime})$ such that $deg_{i}(G)=deg_{i}(G^{\prime})\forall i\in[N]$.
 It is a quick check that $d$ defines an equivalence relation. Hence, we can partition the support of $P(G)$, where each part represents an equivalence class of $D$.
@@
 Proof.
-$G\equiv_{d}G\quad\forall G\in\mathcal{G}$.
-Hence, $d$ is reflexive.
-If for any distinct $G$ and $G^{\prime}$ in $\mathcal{G}$, $G\equiv_{d}G^{\prime}\implies deg(G)=deg(G^{\prime})\implies deg(G^{\prime})=deg(G)\implies G^{\prime}\equiv_{d}G$.
-Hence, $d$ is symmetric.
-Similarly, for any three distinct $G,G^{\prime}$ and $G^{\prime\prime}$ in $\mathcal{G}$, $G\equiv_{d}G^{\prime},G^{\prime}\equiv_{d}G^{\prime\prime}\implies deg(G)=deg(G^{\prime})$ and $deg(G^{\prime})=deg(G^{\prime\prime})\implies deg(G)=deg(G^{\prime\prime})\implies G\equiv_{d}G^{\prime\prime}$.
-Hence, $d$ is transitive.
-∎
+Refer to Supplementary material B.2. ∎
 Procedure 2: Let $(Z_{obs},G_{obs},Y_{obs})$ be the joint treatment assignment vector and its corresponding outcome. Consider a test statistic $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$.
@@
   2. 2.
-    Define the equivalence class $[G_{obs}]=\{G\in\mathcal{G}|\quad G\equiv_{d}G_{obs}\}$. Here, $\equiv_{d}$ is the relation defined in Definition 6.
+    Define the equivalence class $G_{obs}]=\{G\in\mathcal{G}|\quad G\equiv_{d}G_{obs}\}$. Here, $\equiv_{d}$ is the relation defined in Definition 6.
   3. 3.
@@
 We assume that $G_{obs}\sim P_{G}$ where $G$ is a random graph with arbitrary degree distribution.
+Let $G_{X_{i}}=deg(i)$ where $G_{X_{i}}$ denote the graph covariate(s) of unit $i$.
 Consider a test statistic $T(Z,G,Y(Z,G))$.
@@
 Proof.
-Refer to Appendix B.2. ∎
-As a corollary, we can substitute any sub-class of the degree sequence equivalence class in Procedure 2, and the finite-sample validity of the Procedure is retained.
-For example, we can take the equivalence class to be all the graphs isomorphic to the observed graph in the second step of Procedure 2.
+Refer to Supplementary material B.4. ∎
+As a corollary, we can substitute any subclass of the degree sequence equivalence class in Procedure 2, and the finite-sample validity of the Procedure is retained.
+For example, we can take the equivalence class to be all the graphs isomorphic to the observed graph, or automorphic to the observed graph in the second step of Procedure 2.
 We present results for both these Procedures in Section 7.
-We emphasize that generating graphs with a given degree sequence admits a polynomial time algorithm.
+We emphasize that generating graphs with a given degree sequence admits a polynomial-time algorithm.
 Given fixed parameters like the maximum degree of the graph, one can use state-of-the-art algorithms to efficiently sample graphs with a given degree sequence uniformly from the space of all the graphs with the same degree sequence (Arman et al.
@@
 We can also overcome this by effectively taking many samples to approximate the null distribution.
-4.2.2 Test for exponential random graph model
+6.2 Test for exponential random graph model
 We generalize the above method to a class of exponential random graph models where the sufficient statistics of the model are permutation invariant.
-We achieve this by conditioning on the graph isomorphism class and generating the null cases from this equivalence class.
+We can take the graph covariate vector, $G_{X}$, to be the same set of statistics.
+To sample graphs from $G|G_{X}$, we can condition on the graph automorphism class and generate the null cases from this class.
+One caveat is that the size of an automorphic graph class may not be large depending on the symmetry of the graph observed (Erdos and Rényi (1963)).
+To overcome this, we explore sampling approximate automorphic graphs.
+We fix the graph covariate to be the degree of a unit, and sample isomorphic graphs that preserve the degree sequence for the null cases.
+To achieve more generality, one can specify the graph covariate vector of interest and explore advanced graph sampling or rejection sampling techniques.
 We formally present the results below.
@@
 As discussed, some graph statistics include the number of triangles, degree sequence, and number of stars of size two, and they are permutation invariant.
-We model the data-generating process as an exponential random graph with permutation invariance property and present a permutation testing procedure.
+We model the data-generating process as an exponential random graph with a permutation invariance property and present a permutation testing procedure.
 Procedure 3: Let $(Z_{obs},G_{obs},Y_{obs})$ be the joint treatment assignment vector and its corresponding outcome. Consider a test statistic $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$.
@@
     Define the equivalence class
-    $[G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\;|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
+    $G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\;|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}],G\equiv_{d}G_{obs}\}.$
   4. 4.
@@
 We assume that $G_{obs}\sim P_{G}$ where $G$ is a permutation-invariant exponential random graph model.
+Let $G_{X_{i}}=deg(i)$ where $G_{X_{i}}$ denote the graph covariate(s) of unit $i$.
 Consider a test statistic $T(Z,G,Y(Z,G))$.
@@
 Proof.
-Refer to Appendix B.2. ∎
+Refer to Supplementary material B.4. ∎
 6.3 Test statistic
@@
 $T_{I}(G|Z_{obs})=\frac{N_{c}}{N}\cdot T_{I_{c}}+\frac{N_{t}}{N}\cdot T_{I_{t}}.$ (6)
-It should be noted the above test statistics can only be defined if the sub-populations in consideration are non-empty.
-For example, if no control unit exists with all neighbors in control or the number of such units is few, we deal with either a definition or power issue.
+It should be noted that the above test statistics can only be defined if the sub-populations under consideration are non-empty.
+For example, if no control unit exists with all neighbors in control or the number of such units is few, we deal with either a definition or a power issue.
 Thus, we propose a generalization of the above-stated test statistic for robustness.
-We look at the difference in average treatment effect among control units with greater than 0.75 quantile of the proportion of treated units and control units with less than 0.25 quantile of the proportion of treated units.
+We look at the difference in average treatment effect among control units with a greater than 0.75 quantile of the proportion of treated units and control units with a less than 0.25 quantile of the proportion of treated units.
 Let $T_{quant_{c}}(G|Z_{obs})=$
-$\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}$ (6)
+$\hskip-42.67912pt\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\geq p_{0.75})}{|P_{0.75}|}$ (7)
 $-\frac{\sum_{i=1}^{N}Y_{i}\cdot\mathcal{I}(Z_{i_{obs}}=0)\cdot\mathcal{I}(\frac{\sum_{j=1}^{N}G_{ij}\cdot Z_{j_{obs}}}{\sum_{j=1}^{N}G_{ij}}\leq p_{0.25})}{|P_{0.25}|}.$
@@
 In a similar fashion to the test statistic $T_{I}(G|Z_{obs})$, we define the above for treated units and take a weighted average for the two test statistics to obtain $T_{quant}(G|Z_{obs})$.
-We remark that one can improve the test statistic with the knowledge of appropriate covariates or nodal attributes.
+We remark that one can potentially improve the test statistic with the knowledge of appropriate covariates or nodal attributes.
 This can be incorporated by considering regression-adjusted test statistics taking residuals for the potential outcome (Rosenbaum (2002)).
@@
 For example, parametric test statistics can be constructed of the linear-in-means model of interference proposed by Manski (1993).
-We also consider more test statistics, given that the choice is not unique and different test statistics would differ in their power.
+We also consider more test statistics, given that the choice is not unique, and different test statistics would differ in their power.
 We take the test statistic proposed in Bond et al.
@@
 We now present the data-generating process considered in the simulations.
-Our potential outcome function is linearly separable in direct treatment effect and spillover effect.
-The spillover mechanism is a function of the unit’s neighbor.
+Our potential outcome function is linearly separable into direct treatment effect and spillover effect.
+The spillover mechanism is a function of the unit’s neighbors’ direct treatment assignment.
+The potential outcome function also has an additive heterogeneous effect coming from a network-based property, free of a direct treatment assignment vector.
 We propose the mechanism to be contingent upon the proportion of treated neighbors.
-$Y_{i}(\mathbf{z})=\begin{cases}\tau_{\text{direct}}\cdot z_{i}+\tau_{\text{spill}}\cdot\frac{\sum_{j=1}^{N}G_{ij}\cdot z_{j}}{\sum_{j=1}^{N}G_{ij}}+\epsilon_{i}&\sum_{j=1}^{N}G_{ij}>0,\\ \tau_{\text{direct}}\cdot z_{i}+\epsilon_{i}&\text{otherwise}.\end{cases}$ (7)
+$Y_{i}(z)=\begin{cases}\tau_{\text{direct}}\cdot z_{i}+\tau_{\text{spill}}\cdot\frac{\sum_{j=1}^{N}G_{ij}\cdot z_{j}}{\sum_{j=1}^{N}G_{ij}}+\beta_{\text{deg}}\cdot\frac{\sum_{j=1}^{N}G_{ij}}{\text{max}_{i}{\sum_{j=1}^{N}G_{ij}}}+\epsilon_{i}&\sum_{j=1}^{N}G_{ij}>0,\\ \tau_{\text{direct}}\cdot z_{i}+\epsilon_{i}&\text{otherwise}.\end{cases}$ (8)
 We note that $\epsilon_{i}$ can be viewed as the baseline level of outcomes for all the units if they were untreated.
@@
 Such graphs have a high clustering coefficient, but the distance between two nodes in such graphs is also high.
-Table 1: Rejection rates for the null hypothesis of no spillover effects for completely randomized experiment.
-Network Distribution Direct Effect Spillover Effect Degree Permutation Test Isomorphism Permutation Test Conditional Focal Test
-$T_{bond}$ $T_{quant}$ $T_{bond}$ $T_{quant}$ $T_{elc}$
-Small world network 0 0 0.045 0.049 0.049 0.047 0.050
-4 0 0.054 0.046 0.056 0.045 0.046
-0 0.4 0.187 0.148 0.191 0.162 0.112111We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
-4 0.4 0.074 0.151 0.064 0.147 0.060222We obtain the same rejection rate in our replication as reported in Athey et al. (2015).
-Stochastic block model 0 0 0.045 0.062 0.050 0.050 0.054
-4 0 0.050 0.055 0.047 0.052 0.052
-0 0.4 0.191 0.217 0.189 0.216 0.065
-4 0.4 0.064 0.220 0.063 0.217 0.048
 The average distance between two nodes in real-life networks is low (Amaral et al.
@@
 To obtain a graph with a high degree of clustering but a low average path length, one can rewire the edges of a $K$-regular graph with some probability, say $p$.
-This parametrization characterizes a small world network.
+This parametrization characterizes a small-world network.
 The higher the rewiring probability, the lower the average path length.
@@
 The size of the parts are $\{n_{1},n_{2},...,n_{K}\}$.
-These parts represent clusters in the graph, and we can model a random graph using it by defining the probability of an edge between nodes within a part and between two parts.
+These parts represent clusters in the graph, and we can model a random graph using them by defining the probability of an edge between nodes within a part and between two parts.
 Hence, we obtain a probability (symmetric) matrix of size $K\times K$.
-Note that because we want a high degree of clustering in the graph, the diagonal elements of the probability matrix are typically higher than the off-diagonal elements.
+Note that because we also want a high degree of community structure in the graph, the diagonal elements of the probability matrix are typically higher than the off-diagonal elements.
 The probability matrix is also called the preference matrix in the literature.
+We also have a coefficient of network dependence in the potential outcome model, denoted by $\beta_{\text{deg}}$.
+We take $\beta_{\text{deg}}=0$ in Table 2 for a comparative study with the state-of-the-art literature (Athey et al.
+(2018)).
 We now define the setup for the simulation study:
@@
 We then consider a stochastic block model with block sizes $\{50,100,40,110,299\}$.
-The preference matrix which specifies the probability of an edge between two groups is given below:
+The preference matrix, which specifies the probability of an edge between two groups, is given below:
     0.08 0.01 0.01 0.01 0.01
@@
     $T_{bond}(G|Z_{obs})=\frac{\sum_{i,j\in[N]}G_{ij}\cdot Z_{j}\cdot Y_{i}}{\sum_{i,j\in[N]}G_{ij}\cdot Z_{j}}-\frac{\sum_{i,j\in[N]}G_{ij}\cdot(1-Z_{j})\cdot Y_{i}}{\sum_{i,j\in[N]}G_{ij}\cdot(1-Z_{j})}.$
-The null distribution was generated by sampling 1000 graphs with the same degree sequence as the observed graph and isomorphic to the observed graph in a second option.
+The null distribution was generated by sampling 1000 graphs with the same (labelled) degree sequence as the observed graph, and degree-preserving isomorphic to the observed graph in a second option.
+Table 1: Rejection rates for the null hypothesis of no spillover effects for a completely randomized experiment
+Network Distribution Direct Effect Spillover Effect Degree Permutation Test Isomorphism Permutation Test Conditional Focal Test
+$T_{bond}$ $T_{quant}$ $T_{bond}$ $T_{quant}$ $T_{elc}$
+Small world network 0 0 0.045 0.049 0.051 0.045 0.050
+4 0 0.054 0.046 0.046 0.050 0.046
+0 0.4 0.187 0.148 0.193 0.163 0.112111We reported the number presented in Athey et al. (2018). We obtained a rejection rate of 0.073 in our replication.
+4 0.4 0.074 0.151 0.071 0.156 0.060222We obtain the same rejection rate in our replication as reported in Athey et al. (2015).
+Stochastic block model 0 0 0.045 0.062 0.047 0.046 0.054
+4 0 0.050 0.055 0.053 0.056 0.052
+0 0.4 0.191 0.217 0.200 0.206 0.065
+4 0.4 0.064 0.220 0.055 0.214 0.048
 The p-value is calculated as the sample proportion of times the imputed test statistic for the sampled graph realizations is more than the observed test statistic.
@@
 We label this test as ‘Conditional Focal Test’.
-We then perform the simulations for cluster randomized experimental setup.
+We then perform the simulations for a cluster-randomized experimental setup.
 We present the results in Table 2.
 The total number of units in the simulation is the same as before, that is, 599.
-We obtain the clusters in the graph by implementing the epsilon-net clustering algorithm, described in Algorithm 1 (Appendix A).
+We obtain the clusters in the graph by implementing the epsilon-net clustering algorithm, described in Algorithm 1 (Supplementary material A).
 We take epsilon to be equal to 3.
@@
 We implemented a Bernoulli trial at the cluster level.
-Under Procedure 3, the null distribution is approximated by taking 1000 samples of graphs isomorphic to the original graph, such that the corresponding subgraphs on treated units and controlled units remain isomorphic.
-This is obtained by considering permutations where treated units can be mapped only to treated units (and thus, control units can only be mapped to control units).
+Under Procedure 3, the null distribution is approximated by taking 1000 samples of graphs isomorphic to the original graph, such that the corresponding subgraphs on treated units and controlled units remain isomorphic and the labelled degree sequence remains preserved.
+This is obtained by considering permutations where treated units can be mapped only to treated units (thus, control units can only be mapped to control units), and mapping only occurs amongst two units with the same degree.
 We replicated this setup 4000 times and calculated the proportion of times the proposed test rejected the null.
 We label Procedure 3 as ‘Block Isomorphism Permutation Test’ in Table 2.
-This was performed for two network-generating processes as before.
+This was performed for the two network-generating processes as before.
 Table 2: Rejection rates for the null hypothesis of no spillover effects for cluster randomized experiment
@@
 $T_{bond}$ $T_{quant}$
-Small world network 0 0 0.052 0.053
-4 0 0.056 0.051
-0 0.4 0.429 0.352
-4 0.4 0.429 0.348
-Stochastic block model 0 0 0.051 0.057
-4 0 0.048 0.046
-0 0.4 0.225 0.237
-4 0.4 0.227 0.242
+Small world network 0 0 0.053 0.046
+4 0 0.051 0.054
+0 0.4 0.420 0.360
+4 0.4 0.435 0.360
+Stochastic block model 0 0 0.046 0.050
+4 0 0.046 0.052
+0 0.4 0.197 0.206
+4 0.4 0.211 0.214
+For robustness, we also perform simulations for the setup where there is a dependence of the potential outcome function on the graph, devoid of direct treatment assignment.
+For this, we set $\beta_{\text{deg}}=0.4$, installing a positive correlation with the degree of the unit and its corresponding potential outcome.
+We present the results in Table 3.
+The simulations are replicated for both the completely randomized experiment and the cluster randomized experiment.
+The procedures deployed are the degree-preserving isomorphism permutation test and the degree-preserving block isomorphism permutation test.
+The rest of the setup remains the same as before.
+Table 3: Rejection rates for the null hypothesis of no spillover effects for the outcome model including graph covariates.
+Here, $\beta_{deg}=0.4$ and the Procedures deployed are degree-preserving isomorphism/block isomorphism permutation test for a completely/cluster randomized experiment.
+Network Distribution Direct Effect Spillover Effect Completely Randomized Experiment Cluster Randomized Experiment
+$T_{bond}$ $T_{quant}$ $T_{bond}$ $T_{quant}$
+Small world network 0 0 0.052 0.045 0.048 0.046
+4 0 0.050 0.056 0.054 0.053
+0 0.4 0.184 0.162 0.421 0.370
+4 0.4 0.065 0.168 0.415 0.330
+Stochastic block model 0 0 0.050 0.048 0.050 0.048
+4 0 0.051 0.050 0.046 0.043
+0 0.4 0.197 0.216 0.210 0.205
+4 0.4 0.071 0.208 0.186 0.202
 7.2 Comparison with related work
 We see that our proposed methods have type 1 error controlled appropriately.
-This corresponds to rows with $\tau_{spill}$ equal to zero in Tables 1 and 2.
+This corresponds to rows with $\tau_{spill}$ equal to zero in Tables 1, 2, and 3.
 When $\tau_{spill}$ equals 0.4, our method performs with competitive power compared to existing methods in the literature (Athey et al.
@@
 (2018), where the method does not retain power.
-In Appendix D, we add more insights on the performance of the $T_{bond}$ test statistic
-as observed in Table 1.
-An interesting result is in the case of cluster-randomized experiments where a higher degree of clustering shows a drop in power.
+In Supplementary material D, we add more insights on the performance of the $T_{bond}$ test statistic as observed in Table 1.
+An interesting result is in the case of cluster-randomized experiments, where a higher degree of clustering shows a drop in power.
 This can be observed when comparing the power values of the small-world network to the stochastic block model.
-This can be attributed to a reduction of sample space of imputed values of the test statistic.
+This may be attributed to a reduction of sample space of imputed values of the test statistic.
 Hence, a higher degree of clustering increases the cluster sizes, restricting the direct treatment assignment.
@@
 (2022)) where degenerate null distributions are reduced in frequency.
-Overall, our method stands more robust to degeneracy and performs well in terms of power in our simulations.
-We also see that the $T_{bond}$ test statistic performs slightly better than $T_{quant}$ test statistic in some cluster-randomized setups.
-This can be attributed to $T_{bond}$ being weighted by treated numbers in the graph instead of the corresponding quantile-based weights in $T_{quant}$ and the heterogeneity of the clustered assignment setup.
-We emphasize that the $T_{bond}$ in conjunction with our methodology continues to be valid regardless of the size of the direct effect.
+Overall, our method is more robust to degeneracy and performs well in terms of power in our simulations.
+We also see that the $T_{bond}$ test statistic performs slightly better than the $T_{quant}$ test statistic in some cluster-randomized setups.
+This may be attributed to $T_{bond}$ being weighted by treated numbers in the graph instead of the corresponding quantile-based weights in $T_{quant}$ and the heterogeneity of the clustered assignment setup.
+We emphasize that $T_{bond}$ in conjunction with our methodology continues to be valid regardless of the size of the direct effect.
 The method proposed by Bond et al.
@@
 (2018) for an elaborate discussion).
+Our method also retains the same power with graphical confoundedness of the potential outcome function, as shown in Table 3 where $\beta_{\text{deg}}$ = 0.4.
 8 Weather insurance adoption in rural China
@@
 A primary goal of the experiment was to identify if there is information dissemination in the farmers’ community and, if so, what mechanisms drive these information exchanges.
-These are important questions to be answered in this setting as they are consequential to government policy action.
+These are important questions to be answered in this setting, as they are consequential to government policy action.
 A deeper understanding of insurance product adoption via social network mechanisms helps devise a more cost-effective government policy campaign to enhance product adoption among the target community.
@@
 After the round 2 sessions, the first group was given no additional information, the second group was told about the overall take-up rate in the previous round, and the third group was told in detail about who purchased the product and who did not.
-This additional information randomization of the second group was done to derive the primary information diffusion mechanism.
+The randomization of the second group was done to derive the primary information diffusion mechanism.
 For illustration purposes, we do not analyze this part of the experiment.
@@
 We obtained a p-value of 0.0002, which indicates the presence of interference.
-7 Conclusion
+9 Discussion
 We develop a new methodology to test for the non-sharp null hypothesis of no spillover effect in an experimental setup.
@@
 Incorporating covariates or nodal attribute information to characterize the random graph null model better may help enhance the power of the test and is an interesting direction for future work.
+The possibilities of incorporating advanced graph theoretic sampling techniques, and their effect on power, are another interesting line of pursuit.
 We also assume fully and correctly specified information about the given network.
@@
 Since the sampling of null distribution is based on quasi-randomization of the network-generating process, a natural and essential extension is to study random graph null testing procedures in observational studies.
-Acknowledgments.
+Acknowledgement
 We thank Avi Feller, Fabrizia Mealli, Panos Toulis, ACIC 2024 conference participants, and the Tom Ten Have Award committee for their encouragement and feedback.
@@
 We gratefully acknowledge Hemanth Kumar and the ISB Institute of Data Science for providing cluster access.
+The IMS 2024 ICSDS Student Travel Award partially supported Supriya Tiwari’s research.
+Pallavi Basu’s research is partially supported by the SERB MATRICS award MTR/2022/000073.
+Supplementary material
+The Supplementary Material contains proofs, additional simulation results, and algorithmic details.
 Appendix
@@
   • $S$ is equal to $G_{obs}$
-while $S$ is not empty do
+  • While {$S$ is not empty}
   • Step 1: Pick the last vertex in $S$, denoted by $v$, and store in $V$
@@
   • Step 3: Store the units found in Step 2 as a cluster and remove these units from $S$
-end while
-Store the resultant cluster outputs in $C$
-Permute the outputs to the original ordering and store them in $C,V$
-Output: C and V
+  • Step 4: Store the resultant cluster outputs in $C$
+  • Step 5: Permute the outputs to the original ordering and store them in $C,V$
+  • Output C and V
   • Algorithm 1 $\epsilon$-net cluster
     Appendix B Proofs
-B.1 Proof of Theorem 1:
+    B.1 Proof of Proposition 1
+    Proof.
+    Note that under $\mathcal{H}_{0}$,
+    $Y_{{G_{X}}_{i}}(Z,G)=Y_{{G^{\prime}_{X}}_{i}}(Z^{\prime},G^{\prime})\qquad Z_{i}=Z^{\prime}_{i},{G_{X}}_{i}={G^{\prime}_{X}}_{i}.$
+    Thus,
+    $T_{c}(G|Z_{obs})$ $=T(Z_{obs},G,Y_{{G_{X}}_{obs}}(Z_{obs},G)),$
+    $=T(Z_{obs},G,Y_{{G_{X}}_{obs}}(Z_{obs},G^{\prime}))\qquad P(G^{\prime}|Z_{obs},{G_{X}}_{obs})>0\text{ under }\mathcal{H}_{0}.$
+    Since $Z_{obs}$ is arbitrary, we conclude imputability of the $T_{c}(G|Z_{obs})$ under the null $\mathcal{H}_{0}$. Take $G^{\prime}=G_{obs}$, we get $T_{c}(G|Z_{obs})=T(Z_{obs},G,Y_{obs})$. ∎
+    B.2 Proof of Proposition 2
+    Proof.
+    $G\equiv_{d}G\quad\forall G\in\mathcal{G}$.
+Hence, $d$ is reflexive.
+If for any distinct $G$ and $G^{\prime}$ in $\mathcal{G}$, $G\equiv_{d}G^{\prime}\implies deg(G)=deg(G^{\prime})\implies deg(G^{\prime})=deg(G)\implies G^{\prime}\equiv_{d}G$.
+Hence, $d$ is symmetric.
+Similarly, for any three distinct $G,G^{\prime}$ and $G^{\prime\prime}$ in $\mathcal{G}$, $G\equiv_{d}G^{\prime},G^{\prime}\equiv_{d}G^{\prime\prime}\implies deg(G)=deg(G^{\prime})$ and $deg(G^{\prime})=deg(G^{\prime\prime})\implies deg(G)=deg(G^{\prime\prime})\implies G\equiv_{d}G^{\prime\prime}$.
+Hence, $d$ is transitive.
+∎
+    B.3 Proof of Theorem 1
     Proof.
     We define p-value as
-$p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs}).$
-Given $Z_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$. We emphasize that $U$ is a univariate distribution induced by $P(G|Z_{obs})$. Thus,
-$p_{Z_{obs}}=1-F_{U}(G_{obs}|Z_{obs}).$
+    $p_{Z_{obs}}=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z_{obs},{G_{X}}_{obs}).$
+    Given $Z_{obs}$ and ${G_{X}}_{obs}$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$.
+We emphasize that $U$ is a univariate distribution induced by $P(G|Z_{obs},{G_{X}}_{obs})$.
+Thus,
+    $p_{Z_{obs}}=1-F_{U}(G_{obs}|Z_{obs},{G_{X}}_{obs}).$
     Here, $F_{U}(.)$ represents the cumulative distribution of $U$. Then,
     $T_{c}(G|Z_{obs})\stackrel{{\scriptstyle d}}{{=}}T_{c}(G_{obs}|Z_{obs})\qquad\text{given }\mathcal{H}_{0}.$
-Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z_{obs})$ and we get
+    Thus, under $\mathcal{H}_{0}$, the distribution of $U$ is $P(G_{obs}|Z_{obs},{G_{X}}_{obs})$ where the randomness is induced by $G_{obs}\sim P(G|Z_{obs},{G_{X}}_{obs})$ and we get
     $p_{Z_{obs}}=1-F_{U}(U).$
@@
     $P(p_{Z_{obs}}\leq\alpha)\leq\alpha.$
-B.2 Proof of Theorem 2 and Theorem 3:
+    B.4 Proofs of Theorem 2 and Theorem 3
     Proof.
@@
     $p_{Z_{obs}}$ $=P(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})|Z=Z_{obs}).$
-Given $Z_{obs}$ and $[G_{obs}]$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$.
+    Given $Z_{obs}$ and $G_{obs}]$, we define a variable $U$ with the same distribution as $T_{c}(G|Z_{obs})$.
 We emphasize that $U$ is a univariate distribution as induced by $P(G|Z=Z_{obs},G\in[G_{obs}])$.
@@
     $P(p_{Z_{obs},G_{obs}}\leq\alpha)\leq\alpha.$
-Let $P_{1},P_{2},...,P_{d}$ be the partition induced by $\equiv_{d}$. Pick $G_{1},G_{2},...,G_{d}$ such that $G_{i}\in P_{i}$ for all $i=1,2,...,d$. Hence,
+    Let $P_{1},P_{2},...,P_{d}$ be the partition induced by $\equiv_{d}$. Pick $G_{1},G_{2},...,G_{d}$ such that $G_{i}\in P_{i}$ for all i=1,2,,di=1,2,...,d. Hence,
     $P(p_{Z_{obs},G_{i}}\leq\alpha)$ $\leq\alpha\qquad i\in[d],$
@@
     $P(p_{Z_{obs}}\leq\alpha)$ $\leq\alpha.$
-Let $[G_{obs}]=\{G\in\{0,1\}^{N\times N}:\>deg(G)=deg(G_{obs})\}$.
+    Let $G_{obs}]=\{G\in\{0,1\}^{N\times N}:\>deg(G)=deg(G_{obs})\>s.t.\>deg_{i}(G)=deg_{i}(G_{obs})\;\forall i\in[N]\}$.
 We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}])$ is a uniform distribution under a completely randomized experiment.
@@
     $P(G=g|Z=Z_{obs},G\in[G_{obs}])=P(G=g^{\prime}|Z=Z_{obs},G\in[G_{obs}]).$
-Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}]\}$ is uniform under completely randomized experiment. Therefore,
+    Since $g$ and $g^{\prime}$ are arbitrary, we conclude that the distribution of $G$ conditional on the event $\{Z=Z_{obs},G\in[G_{obs}]\}$ is uniform under completely randomized experiment.
+Therefore,
     $p_{Z_{obs},G_{obs}}=\frac{1}{|[G_{obs}]|}\cdot\sum_{G\in[G_{obs}]}\mathcal{I}(T_{c}(G|Z_{obs})>T_{c}(G_{obs}|Z_{obs})).$
     Hence, we prove that Procedure $2$ is valid. We now prove the validity of Procedure $3$. Let
-$[G_{obs}|Z_{obs}]=\{G\in\mathcal{G}\>|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
+    $G_{obs}|Z_{obs}]=\{G\in[G_{obs}]\>|\>G[{Z_{obs}}^{0}]\cong G_{obs}[{Z_{obs}}^{0}],\>G[{Z_{obs}}^{1}]\cong G_{obs}[{Z_{obs}}^{1}]\}.$
     Here, ${Z_{obs}}^{t}=\{i\in\mathbb{P}:\>{Z_{obs}}_{i}=t\}$ for $t\in\{0,1\}$.
-Note that $[G_{obs}|Z_{obs}]\subseteq[G_{obs}]$, where $[G_{obs}]=\{G\in\mathcal{G}|G\cong G_{obs}\}$.
+Note that $G_{obs}|Z_{obs}]\subseteq[G_{obs}]$, where $G_{obs}]=\{G\in\mathcal{G}|G\cong G_{obs}\}$.
 We argue that the distribution of $P(G|Z=Z_{obs},G\in[G_{obs}|Z_{obs}])$ is a uniform distribution under a cluster randomized experiment.
@@
     $P(Z=Z_{obs}\>|\>G=g,$ $G\in[G_{obs}|Z_{obs}])$ (14)
-$=P(\sum_{i=1}^{|C|}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\;j\in[N]\>|\>G=g,G\in[G_{obs}|Z_{obs}])$
-$=P(\sum_{i=1}^{|\pi(C_{obs})|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))={Z_{obs}}_{j}\;j\in[N]\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
-$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
-$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(\pi^{-1}(j)\in{C_{obs}}_{i})=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
-$=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in{C_{obs}}_{i})=t\quad\forall\pi(j)\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
-$=P(W=W_{obs})=P(Z=Z_{obs}|G=g^{\prime},G\in[G_{obs}|Z_{obs}]).$
+    $\hskip-85.35826pt=P(\sum_{i=1}^{|C|}W_{i}\cdot\mathcal{I}(j\in C_{i})={Z_{obs}}_{j}\;j\in[N]\>|\>G=g,G\in[G_{obs}|Z_{obs}])$
+    $\hskip-85.35826pt=P(\sum_{i=1}^{|\pi(C_{obs})|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))={Z_{obs}}_{j}\;j\in[N]\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
+    $\hskip-85.35826pt=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in\pi({C_{obs}}_{i}))=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
+    $\hskip-85.35826pt=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(\pi^{-1}(j)\in{C_{obs}}_{i})=t\quad\forall j\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
+    $\hskip-85.35826pt=P(\sum_{i=1}^{|C_{obs}|}W_{i}\cdot\mathcal{I}(j\in{C_{obs}}_{i})=t\quad\forall\pi(j)\in{Z_{obs}}^{t},t\in\{0,1\}\>|\>G=\pi(G_{obs}),G\in[G_{obs}|Z_{obs}])$
+    $\hskip-85.35826pt=P(W=W_{obs})=P(Z=Z_{obs}|G=g^{\prime},G\in[G_{obs}|Z_{obs}]).$
     Using Equation (13) and (14) in Equation (12), we obtain
@@
 [1960]).
-The below-described model is the probabilist variant of the (Erdős-Rényi) random graph model.
+The model described below is the probabilist variant of the (Erdős-Rényi) random graph model.
 We consider the number of nodes, $N$, and edges, $|E|$, to be fixed, and any two nodes will be linked independently with an edge with a fixed probability $p$.
@@
 It has been shown that the degree distribution of a large Erdős-Rényi random graph is a Poisson distribution.
-Most real-life networks violate this property and have highly skewed degree distribution.
-In the main text, we generalize this and characterize random graphs with arbitrary degree distribution.
+Most real-life networks violate this property and have highly skewed degree distributions.
+In the main text, we generalize this and characterize random graphs with an arbitrary degree distribution.
     C.2 Data generating mechanism
     The following describes the two potential outcome models:
-$Y_{i}(\textbf{z})=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot z_{j}>0)+\epsilon_{i},$ (15)
+    $Y_{i}(z)=\tau_{direct}\cdot z_{i}+\tau_{spill}\cdot\mathcal{I}(\sum_{j=1}^{N}G_{ij}\cdot z_{j}>0)+\epsilon_{i},$ (16)
     $\epsilon_{i}\sim\mathcal{N}(0,1)\qquad\forall i\in[N].$
@@
 Refer to Section 6.3 for further details.
-We generate the randomization distribution of the test statistic in consideration by repeated drawing from the Erdős-Rényi random graph model with the probability of an edge between two nodes being 0.2.
+    We generate the randomization distribution of the test statistic under consideration by repeated drawing from the Erdős-Rényi random graph model with the probability of an edge between two nodes being 0.2.
 The simulations in Table 4 were carried out on 100 units, with 5 being treated and the rest in control.
@@
 This is owing to the simplistic nature of the outcome model (16), and we present its results for verification purposes.
-Another caveat in the above setting is the nature of test statistic $T_{I_{c}}(G|Z_{obs})$.
-We might not get any control units with all neighboring ones in control depending on the input graph, invalidating the procedure.
+Another caveat in the above setting is the nature of the test statistic $T_{I_{c}}(G|Z_{obs})$.
+We might not get any control units with all neighboring ones in control, depending on the input graph, invalidating the procedure.
 So, the above setup involves 100 units with 5 units in treatment.
@@
     In all the above models, we took the probability of an edge between two nodes as $0.2$.
-We present the same setup in 4, with the probability of an edge between two nodes estimated below.
+We present the same setup in 5, with the probability of an edge between two nodes estimated below.
 $G_{obs}$ is generated from Erdős-Rényi random graph with $p=0.2$.